A Complete Bibliography of the ACM Journal on Emerging Technologies in Computing Systems (JETC)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

07 December 2024 Version 1.66

Title word cross-reference

2 [CJ14a, CV12, HYWA09, RSD23b, SMT^{+}17]. 20 [NPS^{+}20]. 3 [ARLB18, AGD^{+}20, BAT^{+}22, VRBS16, BLKM23, BCT^{+}13, CKC^{+}18, CMK^{+}21, CJ14a, CBK^{+}22, CH14, CRKP22, FRB08, KSB^{+}08, KYEB15, KKC17, KWWI17, LKC15, LDK^{+}18, LKK^{+}22, LDP^{+}20, MRG22, SCL17, SKRX13, TJ13b, TTS22, TZS14, VOB19, WL22a, XLBB06, XCF08, XDX14, XPD12, ZJS10, ZMC15, ZCB^{+}22]. 4 [LCSP14]. K [VOB19]. κ [MP10]. μ [RFDT15]. ΣΔ [GGTG^{+}20]. T [YYC07, MPZ21]. Θ(\sqrt{n}) [CV12]. V_{th} [MP10].

-Analog [CFK22]. -Bit [LCSP14]. -D [HYWA09]. -depth [CV12]. -means [VOB19]. -Phase [SMT^{+}17]. -Qubit [NPS^{+}20]. -SNE [MPZ21]. -tree [YYC07].

/high [MP10]. /high- [MP10].

1 [LZBW20]. 100GBd [XLL^{+}18]. 10nm [GVRR17]. 14nm [TGCJ16]. 19 [KKNM22]. 1S1R [BSL^{+}18].

2 [ZLWB20]. 2.0 [BLKM23]. 2006 [CS07]. 2007 [LC08]. 2008 [Bah09]. 2011 [AD14, SN10]. 22nm [TGCJ16].

a-Si [HCTK08]. ABC [PPM^{+}13]. Abnormality [TW22]. Abstraction [DRL^{+}19]. Accelerate [CRSSBMR21, IGGR19]. Accelerated [JCK23, KKNM22, ZJT^{+}14]. Accelerating [AAO21, CRKP22, DEW^{+}23, ERGK21, LWY22]. Acceleration [ASP^{+}18, BRZ21, KMC^{+}22, LZXL22, ORC^{+}24]. Accelerator [DCP^{+}21, DK21, DBS^{+}21, GLL^{+}21b, HM21, KPM22, KPPB17, LLX^{+}18a, NAY24, PJSM17, PAP^{+}22, SBZT20, YLR^{+}23, ZBF^{+}22, ZK18]. Accelerators [BBB^{+}22, CMZR23, GGHP^{+}24, HFLZ22, KAS^{+}22, KKK22, LTO22, LYL19, MEHT21, MMAAK23, WL19, WL22a, XCS^{+}19]. Accelerators-based [MEHT21]. Access [AIK21, GLMG^{+}15, LBJ^{+}16, RSD23b, AKW^{+}13, CSKM13]. Access-Transistor-Free [GLMG^{+}15]. accessibility [GN08]. AccHashtag [JCK23]. Accumulation [BSS16]. Accumulation-Mode [BSS16]. Accuracy [ACH^{+}17, GKT^{+}18, NPH18, TKBM12, WWZ^{+}22]. Accurate [CMJ14, CQZ^{+}21, CLZ^{+}22, ZF15, KTW08]. Achievable [KAKSP14]. achieving [WWJ09]. ACM [BC08, SLCJ22, Shu09, ZLWB20]. across [AMVG12, JRLR15]. Activation [HLH^{+}12, LQYL19, YMWH21]. active [ABS^{+}12, PDLS06]. Activity [MKG^{+}23]. Actuators [ZF15]. Adaptation [LJL18]. Adapting [AIK21]. Adaptive [GLMG^{+}15, MGK18, KMD12, LCT12]. ADC [MSB^{+}22]. Adder [BSL^{+}18, ISI^{+}18, MGMU22, MGZ^{+}17, NV14, CV12, TR13]. Adders [AHHS21]. Addition [CV11]. Address [IASK20]. Addressable [GG17]. addressing [SRD^{+}06]. Adiabatic [KT20]. Advanced [GHHW19, YWF18]. Advances [TSB15]. Adversarial [HMP^{+}22, HYPW22, LDZ22, XZL^{+}21]. AER [KWG^{+}20]. AES [YWF18]. Against [DCSA22, PHS^{+}23, GCC^{+}23, GSC17, LDZ22, RSBA23, ZCX^{+}17]. aggregated [HMP^{+}22]. aging [AMA^{+}14, CNPR14, KCC^{+}14]. Agnostic [CTP14]. AI [JHPG22, PYSJ22].

AI-enhanced [PYSJ22]. Algorithm

[CRSSBMR21, CEW^{+}13, DBS^{+}21, LCSP14, LFDS22, MPZ21, MCT18, PP21, RDH14, SKB13, VOB19, YL14, BBB^{+}22]. Algorithms [CLKG17, CLSD18, CQL21a, CQL21b, MS19, PT14a, SCLW19b, SCLW19a, GMM12, PSM^{+}06]. alive [ABS^{+}12]. All-Optical [GB18, EDCL^{+}22]. All-Spin [MS17, VSRR15, XZR^{+}22]. Alleviate [WXW^{+}17]. Allocation [KXY16, LKL^{+}18, WWJ09]. Allocator [WL22b, YXD^{+}17]. Allow [KMS^{+}20]. alternate [LBGR08]. Alternative [RKM15, TV17]. ALU [SDSS14]. Amplifier [SFD17]. Analog [ASP^{+}18, BY18, BSS16, CHA20, KKK22, MSC^{+}21, RSD^{+}23a, WWZ^{+}22, YLR^{+}23, ZSPC19, CFK22]. Analog-Trojan-resilient [RSD^{+}23a]. Analysis [AHPC21, ARLB18, AHS22, AHHS21, BBB^{+}16, BWL^{+}21, GRPT13, GG17, GFZ13, GPW^{+}15, HON21, KBT24, KYEB15, KCWL^{+}16, LTKP16, MLW^{+}23, NLK^{+}13, PYSJ22, RPBA21, RBHG21, WDG^{+}20, WYZ^{+}20, CCTP08, CSKM13, CWT14, DWL10, HCTK08, KSG14, PHS^{+}23, PFOL07, RBGC14, ZFT13]. Analytic [VOB19]. Analytical [KYEB15, REL^{+}22]. Analytics [CRKP22]. Analyzing [PAC^{+}22]. Annotation [PPM^{+}13]. Anomaly [KPFM16]. Anonymity [WM24]. Anonymous [PHS^{+}23]. ANT [CLZ^{+}22]. ANT-UNet [CLZ^{+}22]. Anti [CCW18]. Anti-Reverse [CCW18]. API [RBG^{+}22]. Application [AIK21, DKK^{+}15, DJH^{+}19, FC18, LZX^{+}21, Tah06, WDG^{+}20, AMVG12, XS14]. Application-driven [WDG^{+}20]. Application-independent [Tah06]. Application-Level [LZX^{+}21]. Application-Specific [DKK^{+}15, FC18]. Applications [ASP^{+}18, BOAC^{+}20, FNO^{+}19, GYM^{+}17, GLL^{+}21b, HVB22, IGGR19, JWJ^{+}17, KPPB17, KHC^{+}22, SJKS20, SLCJ22, SHAC19, TTS22, MFA^{+}13, PFOL07]. Approach [ARD24, BM15, BBB^{+}22, BS15, DRSR14, DJ16, JRLR15, LDK^{+}18, SMR23, SMR^{+}21, ZY18, ZGSA15, CQZK14, RT08, SZSS10]. Approaches [EDZ^{+}23, SMZ^{+}19]. Approximate [AHHS21, GP17, JLL^{+}17, LQYL19, MEHT21, SJKS20, SVA^{+}18, SBT20, TT20]. Approximation [HYPW22]. Arbitrary [Mog14]. Architecting [KBM21, Mit17]. Architectural [ABM21, Mit16, VO06, WKL16]. Architecture [AMH^{+}24, AM18, BYHT18, CRKP22, CNH12, CTT^{+}20, DPB11, GLL^{+}21a, GLL^{+}21b, JOF^{+}15, JMKM21, JRJ22, KDSD24, LZXL22, LLX^{+}18a, LZBW20, LDP^{+}20, LPM^{+}19, MAC^{+}21, NLL^{+}17, SGR^{+}12, VK18, VDB^{+}16, VOB19, WX15, YYBK19, ZLB^{+}22, ZLWB20, ZK19, dLBHC22, CQZK14, CV12, CA11, MTC^{+}08, Moh12, PDLS06, PDL07, SCI^{+}09, TWL09, TCSV09, ZJS09a, ZJS10, ZJS09b]. architecture-level [Moh12]. Architectures [ACM^{+}20, AMF^{+}15, CMM^{+}18, CDP17, CCWCC15, DJ16, ERGK21, GCO^{+}11, GMGA23, GNY^{+}22, KK23, LGL15, MSB^{+}22, MPZ21, Rez23, RMG15, Shu09, WZL16, YJ18, YP17, BPH^{+}11, CDG^{+}12, Deh05, FGZ14, KWFH12, WVGP13, XLBB06, ZMT13]. Area [CKC^{+}18, CCH16, KCWL^{+}16, PFRR17, XZR^{+}22, RT07]. Area-efficient [XZR^{+}22]. Arithmetic [JLL^{+}17, PAP^{+}22, VMNI08, Gla14, WL22a]. AroMa [CMZR23]. Array [KKY^{+}20, LZXL22, LYWW13, MZZ23, MTC^{+}08]. Arrays [BSL^{+}18, CEW^{+}13, CCH16, HHD^{+}23, WHL^{+}21, ZJK22, CCTP08, CSKM13]. Artifact [KRP^{+}21]. Artificial [Dea14, KBM21, WLJC21, XYJ^{+}21]. ASBUS [YWF18]. ASIE [KWG^{+}20]. ASL [ALY^{+}21]. Assays [GCB14]. assembled [GRS05]. Assertions [WJWM23]. Assessment [RNN^{+}22, VOB19]. Assignment [YJ18, SLS^{+}14, ZS08]. Assisted [MPR^{+}22]. Associated [GCO^{+}11]. Associative [CCWCC15, JRJ22, SBR19]. Assurance [BWL^{+}21, GCC^{+}23, AML^{+}23, JDPH^{+}23]. assured [MEHT21]. Asymmetric [CJ16, GVRR17, GJ17, GZZ^{+}21, LPW18]. Asynchronous [GRPT13, GGTG^{+}20, KWG^{+}20, SM11, VGZ11, VSM19, ZSXY11, CB09]. Asynchrony [SN11]. Attack [DDG^{+}22, GCTF20, HMP^{+}22, LZX^{+}21, WHB^{+}21, ZLGL21]. Attacks [ABM21, ASK23, BGX^{+}21, CMZR23, DCSA22, GCC^{+}23, JCK23, JWJ^{+}17, LYEK22, LTO22, LDZ22, PHS^{+}23, RSBA23, WM24]. Audio [HYPW22]. Audit [SPS^{+}24]. Authentic [TFR23]. Authentic/Counterfeit [TFR23]. Authentication [Bis17, IGR^{+}16, MKMS22, RDS24]. Authorization [ASMK22]. Autoencoders [HYPW22]. Automata [DPB11, TNWD20, dLBHC22, DWL10]. Automated [DMR06, GCJ17, WJWM23, XHSC07, ZS08]. Automation [WD22, ZCB^{+}22, CZ05]. Autonomous [LPB^{+}15]. Auxiliary [LDZ22]. Aware [GLMG^{+}15, GUP11, MPM13, MKW^{+}14, NPS^{+}20, PFRR17, PRG^{+}15, SLCJ22, WX15, DJH^{+}19, DLL^{+}19, LHHZ19, MSCS19, RHB^{+}20, RMBC12, SAAR20, STA^{+}12, TMG^{+}21, YWH^{+}13].

B [FCR23]. B-open [FCR23]. Back

[NLW^{+}20]. Backpropagation [DCP^{+}21]. Backside [DDR^{+}16a]. ballistic [HYWA09, PFOL07]. bar [KKY^{+}20]. Barely [ABS^{+}12]. Barnes [MPZ21]. Based [AIK21, BBB^{+}16, BSY^{+}16, BYHT18, CZW^{+}19, DCP^{+}21, DDP20, DKK^{+}15, DJH^{+}19, FYJ^{+}17, GCO^{+}11, GRPT13, GG17, HFLZ22, HC15, HYPW22, KXY16, KBC^{+}23, KHC^{+}22, KT20, LTKP16, LPB^{+}15, LQYL19, LLX^{+}18b, LWY22, MZZ23, MPM13, MSB^{+}22, MKSW17, MAC^{+}21, NLL^{+}17, NPA^{+}12, OBLD14, QGW20, RAC^{+}24, RNN^{+}22, STSG17, SGR^{+}12, SSF^{+}15, TZS14, TSÁ^{+}24, VAK18, VDB^{+}16, VSRR15, WX15, WL19, WLJC21, XLL^{+}18, YXW^{+}12, YLF^{+}17, YYPK17, ZJ11, ZCX^{+}17, ZY18, ZF15, AMH^{+}24, ABR^{+}21, ABM21, ASK23, AHS22, BBB^{+}22, BMTP24, Bis21, CKA23, CZ05, CHN09, DMR06, Deh05, DBS^{+}21, FSD22, GD12, Gla14, GCTF20, HON21, HRR19, HMS^{+}05, IBO22, KPM22, KKY^{+}20, KKNM22, LYEK22, LJ10, LDK^{+}18, LDP^{+}20, LTO22, MDS21, MKMS22, MSCS19, MEHT21, MN06, MS20, NLW^{+}20, RSBA23, Rez23, RDS24, RCYB22, SZSS10, SAAR20, SPR18, SC06, TCSV09, TR13, VOB19]. based [WWG^{+}19, WYZ^{+}20, WL22a, XYJ^{+}21, XZR^{+}22, YYBK19, ZLB^{+}22, ZCDD19, ZH20, WZSC09]. Batching [LML^{+}19]. BCD [TR13]. BD [HYA^{+}20]. BD-Net [HYA^{+}20]. Be [ABM21]. beamed [KK12]. Beat [LTKP16]. Behavior [BOAC^{+}20, GSC17, SMT21]. Behaviors [AIK21]. Benchmark [XCS^{+}19]. Benchmarking [BHLD19]. Benefit [CKC^{+}18]. Best [Bah09]. between [BMTP24, LDPPB21, TGCJ16]. Beyond [KAB^{+}21, KZW^{+}15]. bias [KCC^{+}14]. Biased [WL19]. Big [MLP^{+}20]. BigBus [BPS19]. Binarized [HYA^{+}20, SBR19]. Binary [CCWCC15, GCJ17, HM21, LLX^{+}18a, NHL^{+}17, SBR19, TR13]. biochemical [RBGC14]. Biochips [BMB18, GCB14, HRR19, LBB^{+}18, MPM13, OGB18, CZ05, DJRM09, DDM^{+}06, RMBC12, RBGC14, SC06, SC08, XHSC07, XC08, YYC07]. BioFoundries [DHK^{+}23]. Bioinformatics [Gui13]. Biological [BBR^{+}23]. Biology [Dea14, FHFK14, HD14, MHW14, MSW14, OBLD14, VMV13]. Biomedical [LZCX22, MCH22]. Biosequence [GFZ13]. Bit [GZZ^{+}21, LCSP14, LDZ22]. Bit-plane [LDZ22]. Bit-Widths [GZZ^{+}21]. Bitlet [REL^{+}22]. Bitline [LYWW13]. Black [HYPW22]. Black-box [HYPW22]. Bladder [MGS^{+}12]. Blind [CVK15]. Block [CZQK15, ON15]. Bluetooth [EFRB22]. Bluetooth-enabled [EFRB22]. Board [PB21]. Boards [MGST22]. Boltzmann [YP17]. bond [XS14]. Bonding [KKC17]. Boolean [GKT^{+}18, SWK^{+}16]. Bottom [NLK^{+}13]. Bottom-up [NLK^{+}13]. bound [KHC^{+}22]. bounded [PP21]. box [HYPW22]. BPLight [DCP^{+}21]. BPLight-CNN [DCP^{+}21]. Brain [BY18, KWWI17, WL22a]. Brain-inspired [BY18]. Breaking [WM24]. Bridging [LDPPB21]. broad [GN08]. Brownian [PLC^{+}13]. Budget [TFZ^{+}21]. build [KSB^{+}08]. buildable [LRN05]. Building [BBR^{+}23]. Built [CBK^{+}22, IBO22, MRG22, CWL^{+}13]. Built-In [MRG22, CBK^{+}22, IBO22]. Bulk [BSS16]. Bundled [SMT^{+}17]. Bundled-Data [SMT^{+}17]. Bundling [SBR19]. Bypass [OYLL24, OYLL24].

Cache [ARD24, GB18, HAV^{+}22, NLW^{+}20, VSRR15, ZMC15, TJ13a]. Caches [ON15, PAB^{+}17, SMT21, CWL^{+}13]. CAD [FSDT23]. Calculus [DD14]. Call [SN10]. Can [ABM21, WGY21]. Canonization [GUP11]. Capacity [KXY16, TT21]. Carbon [DLWW08, GRPT13, HC15, LKK^{+}22, HZY^{+}12, MN06, SXL^{+}12]. Carry [MGZ^{+}17]. Case [AGD^{+}20, PPM^{+}13, SCT^{+}22, HCTK08]. Cash [FMTP22]. Causality [MJ11]. CBRAM [MKSW17]. CBRAM-Based [MKSW17]. Cell [BCT^{+}13, GG17, MGMU22, SFD17, XZR^{+}22, SWJ07]. Cell-based [XZR^{+}22]. Cells [CMJ14, DSB16, KCWL^{+}16, MS17, CWL^{+}13]. Cellular [DPB11, LLX^{+}18b, TNWD20, XLW^{+}18, CCTP08, DWL10]. center [BPB^{+}12, KMD12, SMR^{+}12]. centers [AMVG12]. Centric [KPPB17, NYL^{+}20, GNY^{+}22]. CGRAs [JOF^{+}15, TSÁ^{+}24]. Chain [MLP^{+}20, GD12, WFCX09]. chain-based [GD12]. Challenges [BVM^{+}19, FMP^{+}21, Ko12, Nar05, Rez23, MN06, YW13]. Challenging [SMR^{+}21]. Change [WZL16, JRC^{+}13]. Channel [ABM21, VDK^{+}21, CDP17, PYSJ22, PRV^{+}20, ASK23, DDG^{+}22, DCSA22, MLW^{+}23, RPBA21, RBHG21]. Characteristics [KKKK18]. Characterization [ASK23, AMA^{+}14, LCY19, NLW^{+}20, TJ13b, KWFH12]. Charge [RFDT15]. Check [LBJ^{+}16]. Check-Pointing [LBJ^{+}16]. Checking [MZR^{+}14]. Checks [ABR^{+}21]. chemically [CNHL08]. Chip [ASK23, BS21, VRBS16, BKJ19, Bis17, Bis21, CLKG17, CMM^{+}18, CZQK15, CKC^{+}18, CCW18, CTP14, DJ16, FNO^{+}19, GMGA23, KST^{+}22, KDMT22, KK23, LDK^{+}18, LWM^{+}14, LWY22, MSCS19, OYLL24, PDL15, PRG^{+}15, QCF^{+}16, Rez23, SCLW19a, TZS14, VK18, VSM19, WXW^{+}17, WM24, XYM18, ACM^{+}20, BPH^{+}11, CLSD18, CDG^{+}12, CWL^{+}13, CJ14b, CA11, DLL^{+}19, GMM12, KMC^{+}22, LMC^{+}11, LWX^{+}14, SCLW19b, WVGP13, ZFT13, ZXC10, PCD^{+}11, YXW^{+}12]. Chip-Level [CCW18]. Chips [CMK^{+}21, TFR23, PHS^{+}23]. choices [Nar05]. Chronic [MGS^{+}12]. CIM [HVB22]. Circuit [ACH^{+}17, CJ16, DRSR14, GRPT13, GCJ17, GB18, HSZM17, KHR^{+}15, LZBW20, MCT18, NPH18, PB21, SAAR20, TT20, TGCJ16, TW22, TSMCB17, VAK18, ZLWB20, BJ10, DLWW08, KCC^{+}14, MRH12, MMJ09, Moh12, SZSS10, XDX14, YWH^{+}13]. Circuits [AUDS22, AMF^{+}15, BM15, BS15, Che15, CV11, DDP20, DD14, DWK^{+}16, DNHL11, HM14, HLS14, HN12, JLL^{+}17, KZW^{+}15, KDSD24, KKC17, LKK^{+}22, LCSP14, MZR^{+}14, MJ11, PP21, PLC^{+}13, RHB^{+}20, SMZ^{+}19, SM11, SMT^{+}17, TJ13b, TSB15, VGZ11, WWG^{+}19, XZR^{+}22, ZM22, BCT^{+}13, HZY^{+}12, KT14, LRN05, LWH14, LJ14, MHL08, MN06, PSM^{+}06, Sek07, TR10, TR13, WFCX09, XCF08]. Circular [Bis21]. Classical [DD14, MMD^{+}20]. Classification [BYHT18, JLL^{+}17, KPPB17, NYL^{+}20]. Classifier [MSC^{+}21]. Classifiers [LQYL19, LDZ22]. Clock [CH14, Che15, ANR^{+}14, MRH12, XPD12]. Clock-Controlled [Che15]. Clock-Tree [CH14]. Clocking [PP21, SSF^{+}15]. Clockless [MJ11]. Closed [CTT^{+}20, SSN12]. Closed-Loop [SSN12]. Cloud [DHK^{+}23, KK23, PHS^{+}15, AMA^{+}14]. CLU [DK21]. Cluster [DBS^{+}21, YYBK19]. Cluster-based [DBS^{+}21, YYBK19]. Clustering [DRSR14]. Clusters [PPM^{+}13, RT07]. CMOS [ASP^{+}18, CB09, Che15, HN12, HLH^{+}12, KHR^{+}15, MP10, Nar05, NAY24, RT07, RYT^{+}07, SCI^{+}09, SXL^{+}12, ZJS09a, ZJS09c, ZJS09b, ZJS10, ZC07, MRR12]. CMOS-Memristive [NAY24]. CMOS-nano [CB09]. CMOS/ [MRR12]. CMOS/molecular [RYT^{+}07]. CMPs [GB18, SKRX13]. CNFET [MGMU22, PFOL07]. CNN [DCP^{+}21, KPM22]. Co [GKT^{+}18, IGGR19, KKY^{+}20]. Co-design [IGGR19]. Co-optimization [KKY^{+}20]. Co-Processor [GKT^{+}18]. Code [CKA23, HH11, LPW18]. Code-based [CKA23]. Codesign [WXW^{+}17]. Cofactor [SSP14]. Cognitive [KZL15]. Coherent [GB18]. Color [LM13]. Combating [LPW18]. Combination [VMV13]. Combinational [SBR19]. combined [ZFT13]. Communication [KDMT22, KK23, LGL15, NVW^{+}22, WM24, LMC^{+}11, SX11]. Community [HIH18]. Compact [JMKM21, SCT^{+}22, DLWW08]. Comparative [DDR^{+}16a, JLL^{+}17, KWC^{+}20, KST^{+}22, KCWL^{+}16]. Compare [REL^{+}22]. Comparison [LLSO17, PYSJ22]. Compartmentalization [RD22]. Compatible [KCD15, KCC^{+}14]. compensation [MRH12]. Competitive [KSA^{+}22, KHC^{+}22]. Compilation [HFLZ22, NPS^{+}20]. Compiler [LKK^{+}22, RBG^{+}22]. Compilers [SPS^{+}24]. complete [WWG^{+}19]. completion [MNT14]. Complex [PSY^{+}18, TT21, WVGP13]. Complexity [MCH22, SGR^{+}12]. Component [AHPC21]. Components [Mit17]. Composable [MHW14, ZCDD19]. Comprehensive [VOB19]. Compressing [TFZ^{+}21]. Compression [MCH22]. Compressive [QGW20]. Compressors [GP17]. Computation [AL17, CVK15, HVB22, HSZM17, MSB^{+}22, MMD^{+}20, TNWD20, YLF^{+}17, ZLB^{+}22, YWH^{+}13, WDT14]. Computation-In-Memory [HVB22, MSB^{+}22, ZLB^{+}22]. Computational [KHC^{+}22, MSW14, TT21]. Computations [MDS21]. Compute [WWZ^{+}22]. Compute-in-Memory [WWZ^{+}22]. Computed [XYJ^{+}21, XZL^{+}21]. Computer [AVK16, EDZ^{+}23, GNY^{+}22, WZL16, CV12]. Computers [JRLR15, PSY^{+}18, SCT^{+}22]. Computing [Ano18, ASP^{+}18, BY18, BOAC^{+}20, BH17, CQZ^{+}21, DMYT15, EDCL^{+}22, FYJ^{+}17, FSD22, GIS^{+}22, HN15, JRLR15, KZL15, KKK22, KWWI17, LGL15, LP17, Mit16, MS20, NLL^{+}17, NV14, NYL^{+}20, NHL^{+}17, PAF18, RCYB22, SJKS20, SCL17, SMR23, SVA^{+}18, SBR19, SDSS14, SK16, SHB^{+}21, SPR18, TTS22, TSMCB17, TV17, VAK18, WYZ^{+}20, WD22, XNK18, ZSPC19, ZK18, AMA^{+}14, KMD12, KT14, MHL08, McK07, PG12, VO06, WZSC09, WDH^{+}09, YW13].

Computing-In-Memory

[KKK22, WYZ^{+}20]. CONCEALING [RDM^{+}21]. CONCEALING-Gate [RDM^{+}21]. Conditional [PSR17]. Configuration [Bis17]. configuring [RT08]. Congenital [XYJ^{+}21]. Congestion [MKW^{+}14, RMBC12]. Congestion-Aware [MKW^{+}14, RMBC12]. Connected [AGD^{+}20]. Connections [LKC15]. Connectivity [OYLL24]. Conservative [PPM^{+}13]. consideration [LWH14]. Considerations [MRR12, UMB^{+}18, ZM22, BJ10, WOW^{+}10]. considering [RYT^{+}07, SLS^{+}14]. Consistent [XZL^{+}21]. Constants [Mog14]. Constrained [OGB18, SCLW19b, SCLW19a, TSÁ^{+}24, TMG^{+}21, XHSC07]. Constraint [WXW^{+}17]. Constraints [CCH16, CNHL08, XHSC07]. Constructions [CKA23]. Consumption [FC18, LBJ^{+}16, LYWW13]. Contactless [RDM^{+}21]. Content [GG17]. Contradictory [DWK^{+}16]. Control [GYM^{+}17, GCB14, QGW20, RSD23b, ZXC10]. control-path [ZXC10]. Controllable [MGZ^{+}17]. Controllable-Polarity [MGZ^{+}17]. Controlled [Che15, DNHL11, HZSA14, MSB^{+}22, SXL^{+}12]. Controllers [ARD24]. Converter [TZS14]. Convolution [HYA^{+}20, KPM22, SPR18]. Convolution-over-time [SPR18]. Convolutional [AHS17, DK21, KRP^{+}21, LLX^{+}18a, LPM^{+}19, ORC^{+}24, PJSM17, PSL^{+}19, TZR20, ZH20]. Coprocessor [ASP^{+}18]. Core [BKJ19, DMYT15, KDMT22, KPFM16, Rez23, YWF18, DJH^{+}19]. correlated [MLW^{+}23]. Correlation [AAFM13, WM24]. Cortically [DBS^{+}21]. Cosine [DBG^{+}14, KSA^{+}22]. COSMO [GIS^{+}22]. Cost [GLL^{+}21a, GCJ17, HSZM17, KXY16, KHC^{+}22, LCSP14, LNL19, MRG22, LBGR08, TR10]. Cost-competitive [KHC^{+}22]. Cost-Effective [MRG22]. Cost-Efficient [GLL^{+}21a]. count [MCT18]. Counter [HON21]. Counterfeit [TFR23]. Countering [LYEK22]. Countermeasure [BBB^{+}16, GCTF20]. Countermeasures [DLTSNA20]. Counting [VAK18]. Coupled [GKT^{+}18, YLF^{+}17, WWG^{+}19]. Coupling [AGR^{+}23, KKC17]. COVID [KKNM22]. COVID-19 [KKNM22]. CPDI [XDX14]. CPU [REL^{+}22]. Critical [CWT14, SPS^{+}24, SHAC19]. Critical-reliability [CWT14]. Criticality [DRL^{+}19, WDG^{+}20, YWH^{+}13]. Cross [BS15, DDG^{+}22, DKK^{+}15, KWC^{+}20, KST^{+}22, KKY^{+}20, LYWW13, PRG^{+}15, SS15, ZGSA15, XDX14]. Cross-bar [KKY^{+}20]. Cross-device [DDG^{+}22]. Cross-Layer [BS15, DKK^{+}15, KST^{+}22, PRG^{+}15, SS15, ZGSA15, KWC^{+}20]. Cross-Point [LYWW13]. cross-power [XDX14]. Crossbar [BSL^{+}18, KZL15, LLSO17, MZZ23, NHL^{+}17, UMB^{+}18, WDW13, WRWW17, YL14, ZJK22, ZH20, ZK18, CQZK14, Tah09, ZMT13]. Crossbar-based [ZH20]. Crossbars [MS19, PDL15]. Crossover [RHB^{+}20]. Crossover-aware [RHB^{+}20]. Crosstalk [SMR23]. Crypsis [GSC17]. Crypto [PHS^{+}23]. Crypto-analysis [PHS^{+}23]. Cryptographic [ABR^{+}21, RSBA23]. Cryptography [CHA20]. Cryptosystems [CKA23]. Crystal [EDCL^{+}22]. CSMO [WDG^{+}20]. CT [KBC^{+}23]. Cubes [DRSR14]. Current [AMH^{+}24, KKKK18, MGK18, RFDT15]. Current-Voltage [KKKK18]. Cyber [LBB^{+}18]. Cyber-Physical [LBB^{+}18]. Cycle [FSDT23, XZL^{+}21, ZF15, SZSS10]. Cycle-Accurate [ZF15]. cycle-based [SZSS10]. Cycles [JRLR15]. cycling [GD12].

D [CV12, ARLB18, AGD^{+}20, BAT^{+}22, VRBS16, BLKM23, BCT^{+}13, CKC^{+}18, CMK^{+}21, CJ14a, CBK^{+}22, CH14, CRKP22, FRB08, HYWA09, KSB^{+}08, KYEB15, KKC17, KWWI17, LKC15, LDK^{+}18, LKK^{+}22, LDP^{+}20, MRG22, SCL17, SKRX13, TJ13b, TTS22, TZS14, VOB19, WL22a, XLBB06, XCF08, XDX14, XPD12, ZJS10, ZMC15, ZCB^{+}22]. D-IC [BLKM23, CH14]. D-NoC [VOB19]. D-ReG [LDP^{+}20]. DAC [CS07, LC08]. DAHM [AMVG12]. Dark [KDMT22]. Data [ACM^{+}20, FHFK14, NLW^{+}20, PSM^{+}06, SMT^{+}17, SCZ^{+}12, VTKT22, ZPL^{+}20, AMVG12, ABS^{+}12, BPB^{+}12, KMD12, SMR^{+}12, ZJS09a, VTKT22]. Dataflow [GLL^{+}21b]. Dataset [JDPH^{+}23]. DC [TZS14]. DCT [BBB^{+}22]. DDRx [HTMH18]. De-obfuscation [GCTF20]. Deadline [SN10]. Deadlock [LKC15]. Deadlock-Free [LKC15]. decimal [Gla14]. Decipher [JMKM21]. decision [LJ14]. Decomposition [HZSA14]. Deep [ASMK22, AHS17, ATW^{+}22, AAO21, CKC^{+}18, CQZ^{+}21, CMZR23, DCP^{+}21, DDG^{+}22, DEW^{+}23, HHD^{+}23, KKKK18, KMC^{+}22, LDP^{+}20, LDZ22, PSR17, PAP^{+}22, PSY^{+}18, RBHG21, RBG^{+}22, SM19, SMR^{+}21, STNP21, TWLL19, WL22a, YLR^{+}23]. DeepPeep [JMKM21]. Defect [FCR23, GUP11, WDW13, WHL^{+}21, YL14, DWL10, PDL07, SCI^{+}09, Tah06, Tah09, TWL09, XC08, YYC07]. Defect-Aware [GUP11]. Defect-Free [YL14]. defect-tolerant [YYC07, ZMT13]. defect/error [TWL09]. defect/error-tolerant [TWL09]. Defects [CHN09, FMW^{+}22, KKC17, MRG22]. Defending [LDZ22]. Defenses [BGX^{+}21]. degradation [Edi14, SLS^{+}14]. Delay [BY18, CMJ14, CKWK18, Gla14, GCTF20, KKC17, LMM18, SMT^{+}17, TGCJ16, CWT14, TR10]. Delay-based [Gla14, GCTF20]. Delay/Power [TGCJ16]. Delivery [HLH^{+}12, WXW^{+}17, ZSXY11, ZS08].

Demand [HLH^{+}12]. Denoising [XZL^{+}21]. Dense [SBR19]. Denser [RMW^{+}17]. dependability [PUBV07, TG07]. Dependence [NPA^{+}12]. Dependent [AMF^{+}15, ZWL^{+}15, LYL19]. Deployment [PJSM17]. Deposited [LLSO17, BPH^{+}11]. Depth [HYA^{+}20, PP21, CV12]. Depth-bounded [PP21]. Depth-wise [HYA^{+}20]. Depthwise [KRP^{+}21]. Depthwise-Separable [KRP^{+}21]. Descent [HHD^{+}23]. Describing [KAB^{+}21]. Design [ASB^{+}21, AMH^{+}24, ACH^{+}17, AHS22, AGR^{+}23, BBB^{+}22, BKJ19, BSY^{+}16, CZ05, CJ16, CTP14, CTT^{+}20, DKK^{+}15, DRL^{+}19, FSDT23, GRPT13, GP17, GG17, GJ17, HM14, JMKM21, JWJ^{+}17, KBT24, KHC^{+}22, KT20, LPB^{+}15, LDK^{+}18, LPW18, LML^{+}19, LZBW20, LZCX22, MRR12, MSB^{+}22, MS20, MCT18, NPH18, NV14, OBLD14, OYLL24, PCD^{+}11, PUBV07, RDM^{+}21, Rez23, SFD17, SAAR20, SDSS14, SS15, SN11, TJ13a, TR10, TR13, TTS22, TSB15, TSMCB17, UMB^{+}18, WOW^{+}10, WD22, XLBB06, XZR^{+}22, ZJK22, ZJS09a, ZLWB20, ZSPC19, ZCB^{+}22, ZGSA15, dLBHC22, BJ10, BCT^{+}13, CB09, CDG^{+}12, CJ14b, CNHL08, DMR06, DLWW08, Gla14, GRS05, HMS^{+}05, HZY^{+}12, IGGR19, KP10, LBGR08, LMC^{+}11, MLK^{+}08, MRH12, MN06, Nar05, OSLT06, RMBC12, SXL^{+}12, WFCX09, XDX14, XHSC07, ZC07, ZXC10, ZJS09c]. design-considerations [BJ10]. Design-time [AMH^{+}24]. Designing [AVK16, ARD24, DBG^{+}14, RYT^{+}07, TKBM12]. Designs [ACJ17, FNO^{+}19, SM19, TZS14, ANR^{+}14, ZS08]. Detect [JWJ^{+}17, TFR23]. Detecting [JCK23]. Detection [ABM21, ATW^{+}22, BYHT18, CKWK18, DRG21, GLL^{+}21a, HIH18, KBC^{+}23, KPFM16, MGMU22, MLP^{+}20, MKG^{+}23, SGR^{+}12, TW22, ZFT13]. Detection-Based [KBC^{+}23]. Detector [LTKP16]. Detector-Based [LTKP16]. Development [FMW^{+}22]. Device [BJ10, CJ14a, EFRB22, HVB22, HD14, LZBW20, SAAR20, YJ18, ZLWB20, ZK19, DDG^{+}22, Edi14, RYT^{+}07]. Device-aware [SAAR20]. Device-Type [YJ18]. Devices [ALY^{+}21, BOAC^{+}20, CKB20, GBLD15, KT20, LZCX22, MZZ23, NPA^{+}12, PDL15, STSG17, TFZ^{+}21, TMG^{+}21, JRC^{+}13, MHL08, RT08, YW13]. DFR [BY18]. Diagnosis [WHL^{+}21, DJRM09, DDM^{+}06]. Diagonal [HZSA14]. diagram [LJ14]. Diamond [ATW^{+}22]. Digit [KSA^{+}22]. Digital [BMTP24, BMB18, Che15, GLL^{+}21a, GCB14, HM14, HLS14, IBO22, KZL15, LBB^{+}18, MPM13, OGB18, DJRM09, DDM^{+}06, KT14, LZCX22, RMBC12, RBGC14, SC08, XHSC07, XC08, YYC07, ZXC10]. Digital-Microfluidic [LBB^{+}18]. dilution [RBGC14]. Dimension [ST20]. Dimensional [GUP11, RAC^{+}24, MLK^{+}08, WFCX09, XS14]. DINOS [VTKT22]. diode [BJ10, DMR06]. diode-based [DMR06]. diodes [LM13]. Directed [CKB20]. Direction [HMP^{+}22]. Direction-aggregated [HMP^{+}22]. Directional [NVW^{+}22]. Disassembly [PRV^{+}20]. Disease [XYJ^{+}21]. displacements [SWJ07]. dissipation [MHL08]. Distance [CV11, TT20]. Distributed [AAO21, LGYC21, NHL^{+}17, AMVG12, STA^{+}12, VMNI08]. distributed-memory [VMNI08]. Distribution [MGST22, RSBA23, XPD12]. Disturbance [ZCX^{+}17]. Diverse [OK22]. Division [XLL^{+}18]. DL [DDG^{+}22]. DMAC [RSD23b]. DNA [MT14, SKB13, VTKT22]. DNN [HYA^{+}20, LTO22, MMAAK23, SPS^{+}24]. DNNs [JMKM21]. Domain [KWC^{+}20, Mit17, SBZT20, XDX14]. Domain-Wall [Mit17, SBZT20]. Dormant [DRG21]. Dot [DPB11, TNWD20, DWL10, WDH^{+}09]. DPA [DLTSNA20, ZJ11]. DRAM [ARD24, YYBK19]. DRAM-Flash [YYBK19]. DRAMs [BJ10]. Driven [LHW^{+}17, BGX^{+}21, GMM12, WDG^{+}20, XS14]. driver [HCTK08]. Droplet [GHHW19, HRR19, MPM13, XHSC07, XC08]. Droplet-Aware [MPM13]. droplet-interference [XHSC07]. Dropout [IBO22]. Drug [HLH^{+}12]. DSE [WDG^{+}20]. DSP [TWL09]. DTA [TMG^{+}21]. DTA-PUF [TMG^{+}21]. Dual [DLTSNA20, MFA^{+}13, PAF18, PP21, YJ18, MP10]. dual- [MP10]. Dual-Mode [PAF18]. Dual-Rail [DLTSNA20]. Durable [BGX^{+}21]. duty [GD12]. DVFS [MKW^{+}14, ZF15]. DWT [SGR^{+}12]. Dynamic [CThG15, GB18, JOF^{+}15, MRH12, MEHT21, QGW20, SMT21, SHB^{+}21, TMG^{+}21, YMWH21, ZMC15, AMVG12, DLL^{+}19, WWJ09]. Dynamically [AIK21, ZJS09a, ZJS09c, ZJS09b, ZJS10].

Early [Ko12, XLW^{+}18, ZJK22, ZGSA15]. Early-Stage [ZGSA15]. Easy [DDR^{+}16b]. ECC [PFRR17]. ECDSA [BBB^{+}16]. ECG [SCZ^{+}12, ZBF^{+}22]. Ecosystem [OK22]. Edge [CQZ^{+}21, KK23, MZZ23, MGMU22, SPS^{+}24, XZL^{+}21]. Editor [CLKG17, SCLW19a, Ano18]. Editorial [CS07, Cha10, CFK22, HVB22, IN05, JHPG22, Kar20, McK07, Nar08, SLCJ22, SK16, TSB15, TSMCB17, TV17, XCF08, ZLWB20, Shu09]. Editors [CLSD18, FNO^{+}19, GMGA23, SCLW19b]. EEG [KRP^{+}21]. Effect [CV11, LYWW13, XPD12, HZY^{+}12]. Effective [MRG22, MKG^{+}23, ZBF^{+}22]. effects [MLK^{+}08]. Efficiency [CMM^{+}18, LLSO17, TKBM12, YMWH21, ZS08]. Efficient [BYHT18, CQL21a, CQL21b, CRKP22, DDG^{+}22, DJRM09, GLL^{+}21a, GLL^{+}21b, HN12, HHD^{+}23, KDMT22, KT20, LKC15, LPW18, LMZZ24, MGMU22, MMAAK23, NV14, NAY24, PJSM17, PSR17, PAF18, RSD23b, SVA^{+}18, SMT21, SDSS14, SMT^{+}17, SPR18, TZR20, VSRR15, WGY21, XLW^{+}18, ANR^{+}14, BY18, BS21, GD12, KSB^{+}08, LML^{+}19, LYL19, MS20, MSC^{+}21, PT12, TR13, XZR^{+}22, SM11]. Elastic [PHS^{+}15, ZM22]. Elasticity [GOGCK11]. Electric [RFDT15]. Electrical [YLR^{+}23]. Electro [ACJ17, DEW^{+}23]. Electro-Photonic [ACJ17, DEW^{+}23]. Electroencephalography [TKBM12]. Electrograms [ZBF^{+}22]. Electromagnetics [CFK22]. Electron [CEW^{+}13, CCH16, WHL^{+}21, HYWA09]. Electronic [MGST22, YXW^{+}12, JRC^{+}13]. Electronics [BY12, CFK22, Ko12, HCTK08, WZSC09]. Electrostatic [GPW^{+}15, KKC17]. electrostatics [KTW08]. Electrothermal [CSKM13, HLH^{+}12]. Element [SMT^{+}17]. elements [CW08]. EM-X-DL [DDG^{+}22]. Embedded [FKM22, JCK23, JWJ^{+}17, LBJ^{+}16, SMR23, TZR20, WHB^{+}21, MCT10]. Embedding [HIH18, SWK^{+}16]. Embryonics [TMM^{+}07]. Emerging [BSY^{+}16, DMYT15, DLTSNA20, FMP^{+}21, FNO^{+}19, GBLD15, KZW^{+}15, KBM21, MSB^{+}22, MAC^{+}21, MPR^{+}22, TSB15, WZSC09, BC08, Edi14, PUBV07]. Employing [KST^{+}22]. Emulation [RCYB22]. Enabled [CDP17, Rez23, EFRB22, NVW^{+}22, WVGP13]. Encoder [QGW20]. Encoding [ZWL^{+}15]. End [FKM22]. End-to-End [FKM22]. Energy [ACH^{+}17, BY18, CMM^{+}18, CRKP22, DNHL11, GD12, KT20, LPB^{+}15, LKC15, LLSO17, LML^{+}19, LMZZ24, LYWW13, LJL18, LYL19, MS20, NLW^{+}20, NAY24, PFRR17, PSR17, PAF18, RFDT15, SVA^{+}18, SCLW19b, SCLW19a, STA^{+}12, SMT^{+}17, SPR18, TKBM12, TNWD20, VSRR15, ZSXY11, KMD12, KSB^{+}08, KP10, MHL08, MCT10, SMR^{+}12, WOW^{+}10, WCSA10, SM11]. Energy- [STA^{+}12]. energy-adaptive [KMD12]. Energy-Constrained [SCLW19b, SCLW19a]. Energy-Efficiency [LLSO17]. Energy-Efficient [CRKP22, KT20, LKC15, LMZZ24, NAY24, PSR17, SVA^{+}18, SMT^{+}17, SPR18, VSRR15, BY18, GD12, LML^{+}19, LYL19, MS20, SM11]. Energy-Neutral [LPB^{+}15]. enforced [RD22]. Engine [ERGK21, KWG^{+}20, ZK18]. Engineering [BWL^{+}21, CCW18, QCF^{+}16]. Engines [AL17, WWZ^{+}22]. Enhanced [AHS22, PYSJ22]. enhancement [SC06]. Enhancing [FMTP22, KMD12]. Ensembles [WGY21]. Entanglement [ST20]. Environment [RNN^{+}22, OSLT06]. Epidemiology [KKNM22]. Epilepsy [SSN12]. Epileptic [SGR^{+}12]. equation [KTW08]. Era [KDMT22, MAC^{+}21, SMR^{+}21, TSMCB17]. Error [AHHS21, GYM^{+}17, NPS^{+}20, LWX^{+}14, ZXC10]. Error-Aware [NPS^{+}20]. Error-Tolerant [GYM^{+}17, TWL09]. Errors [LPW18, SKRX13]. ESN [GLL^{+}21a]. ESOP [DRSR14]. Estimation [CMJ14, GCJ17]. Eternal [RSD^{+}23a]. Eternal-thing [RSD^{+}23a]. Eucalyptus [AMA^{+}14]. Evaluating [CMZR23, RT07, SJKS20]. Evaluation [CTP14, DRL^{+}19, GRS05, IBO22, JLL^{+}17, MKW^{+}14, WGY21, CDG^{+}12]. Events [KWG^{+}20]. EVHA [AML^{+}23]. Evolutionary [HM14, Sek07]. Evolving [TG07]. Exact [TSÁ^{+}24]. Examples [HMP^{+}22, HYPW22]. Exascale [DMYT15]. Execution [MPM13]. Existing [WWC23]. Exit [XLW^{+}18]. Explainable [AML^{+}23]. Explained [MLP^{+}20]. Explicit [ABR^{+}21]. Exploiting [ACM^{+}20, DK21, JMKM21, KXY16, KDSD24, SLC^{+}17, VDB^{+}16]. Exploits [WYZ^{+}20]. Exploration [LDK^{+}18, WKL16, ZJK22, TJ13a, XLBB06, ZJS09a, ZC07]. Exploring [KAB^{+}21, RD22, SKRX13]. Extended [KDSD24, PPM^{+}13]. Extensible [KAKSP14]. Extension [MMD^{+}20]. Extracting [NPS^{+}20]. Extraction [YL14].

Extractor [RFDT15]. Extreme [KKK22].

F [OYLL24]. F-Bypass [OYLL24]. Fabric [DPB11]. Fabrication [CCH16, VDB^{+}16]. Fabrication-Induced [VDB^{+}16]. Fabrics [NLK^{+}13]. Failure [KYEB15, KKY^{+}20, PFRR17, VAK18]. Failure-Aware [PFRR17]. Fast [KBC^{+}23, SMT21, SKB13, WDG^{+}20, YL14, ZCSG21]. Fault [ABR^{+}21, BBB^{+}16, BKJ19, CBK^{+}22, CVK15, DJ16, FMW^{+}22, HH11, IBO22, JCK23, JWJ^{+}17, LDPPB21, LCK19, MGZ^{+}17, PHS^{+}23, SLC^{+}17, VAK18, DDM^{+}06, SCI^{+}09]. Fault-Based [BBB^{+}16, IBO22]. Fault-Injection [JCK23, PHS^{+}23]. Fault-Tolerant [BKJ19, CVK15, DJ16, HH11, LCK19, MGZ^{+}17, ABR^{+}21]. faults [CHN09]. Feature [SPR18]. Feedback [BY18]. Ferroelectric [LWY22]. Ferroelectric-Based [LWY22]. Field [AGR^{+}23, ATW^{+}22, BMB18, KR18, NPA^{+}12, WWG^{+}19, HZY^{+}12]. Field-coupled [WWG^{+}19]. Field-Coupling [AGR^{+}23]. field-effect [HZY^{+}12]. Fields [CThG15]. File [WX15]. Files [ZCX^{+}17]. Fin [BSS16]. Fine [BLKM23, SMT^{+}17, MLK^{+}08]. Fine-Grain [SMT^{+}17, MLK^{+}08]. Fine-grained [BLKM23]. FinFET [BJ10, BSS16, CMJ14, CJ14a, CJ14b, CJ15, DLTSNA20, FCR23, GJ17, KCWL^{+}16, LJ10, MMJ09, SSF^{+}15, TJ13a, TJ13b, TGCJ16, YJ18, ZJ11]. FinFET-Based [SSF^{+}15, ZJ11, LJ10]. FinFETs [CJ16, GVRR17, GJ17]. Fingerprinting [Bis21]. Firmware [FKM22]. Flash [HC15, YYBK19]. Flexible [BKJ19, KRP^{+}21, MMAAK23, PAF18, HCTK08, LWH14]. Flip [AM18]. Flip-N-Write [AM18]. Floating [HC15, NV14]. Flow [GCB14, WM24, ZPL^{+}20, DMR06, ZJS09c]. Fluidic [WWC23]. fluids [RBGC14]. Fluigi [HD14]. fly [IASK20]. Folding [MMAAK23]. Forest [MPZ21]. Form [CTT^{+}20]. Formal [CCTP08, GYM^{+}17]. Formalizing [FHFK14]. formulation [YYC07]. formulations [ZMT13]. Fortifying [LTM22]. FPGA [AHPC21, AHS22, BRZ21, CKA23, GLL^{+}21a, HON21, LLX^{+}18a, LLX^{+}18b, LYL19, LMM18, RCYB22]. FPGA-Based [LLX^{+}18b, AHS22, HON21, RCYB22]. FPGAs [AAO21, RT07, ZPL^{+}20]. FPIC [JDPH^{+}23]. Framework [CJ15, FKM22, FMW^{+}22, HFLZ22, KAS^{+}22, KBC^{+}23, KPFM16, LPB^{+}15, LH20, LLX^{+}18b, NPH18, PRG^{+}15, RCYB22, SDSS14, XYJ^{+}21, ZJK22, MCT10]. Fredkin [DJ08]. Free [DBG^{+}14, GLMG^{+}15, LKC15, Mog14, WRWW17, YL14]. Frequency [CFK22, EFRB22, JOF^{+}15, KAKSP14, LTKP16]. Frontiers [CLSD18]. Frontside [DDR^{+}16a]. Full [MGMU22]. Fully [CCWCC15, KXY16, TZR20]. Function [LQYL19, TMG^{+}21]. Functional [GCO^{+}11, Sek07]. Functions [DDR^{+}16b, FSD22, LTM22, SWK^{+}16, UMB^{+}18, PT12]. Fundamentals [PLC^{+}13]. Fusing [ORC^{+}24]. Future [BLKM23, DLTSNA20, GMGA23, KAB^{+}21, RSD23b, CA11, MN06].

GALS [PCD^{+}11]. Gap [LDPPB21]. Garbage [DBG^{+}14, Mog14, TR10]. Garbage-Free [DBG^{+}14, Mog14]. GARDENIA [XCS^{+}19]. gas [HYWA09]. Gate [HC15, LSH14, RDM^{+}21]. Gated [BJ10, JRJ22]. Gated-diode [BJ10]. Gated-RRAM [JRJ22]. Gates [AHSZ16, CJ14a, DWK^{+}16, HZSA14, DJ08]. Gating [HN12, ZF15]. Geckos [GSC17]. gem5 [CJ15]. gem5-PVT [CJ15]. Generalized [WHL^{+}21, RBGC14]. Generation [AMH^{+}24, HYPW22, KK23, MKG^{+}23, TW22, WJWM23, XCS^{+}19]. Generator [CKA23]. Generators [LTKP16, ST20]. Genes [AAFM13]. Genetic [ARD24, BBB^{+}22, MZR^{+}14].

Genetic-algorithm-based [BBB^{+}22]. genomics [GN08]. geographically [AMVG12]. Geometric [DSB16]. Global [DDP20, XPD12]. GPGPU [LCY19, TWLL19, WX15]. GPU [BS21, CRKP22, LLX^{+}18a, LML^{+}19, MPZ21]. GPU-Outperforming [LLX^{+}18a]. GPUOPT [BS21]. GPUs [SPS^{+}24, ZCX^{+}17]. Gradient [HYPW22, HHD^{+}23]. Grain [SMT^{+}17, MLK^{+}08]. grained [BLKM23]. Graph [CRKP22, FSD22, MDS21, PP21, XCS^{+}19]. Graph-based [FSD22]. Graphene [DDP20, HC15, KHR^{+}15, WLJC21, WZSC09]. Graphene-Based [WLJC21, WZSC09]. green [AMVG12, PG12]. greener [GD12]. Grid [ZGSA15]. Guarantees [GYM^{+}17]. Guarding [DCSA22]. Guest [CLKG17, CFK22, HVB22, JHPG22, SCLW19a, SLCJ22, Shu09, SK16, TSB15, TSMCB17, ZLWB20, Ano18, CLSD18, FNO^{+}19, GMGA23, SCLW19b]. Guided [MCH22, WDG^{+}20].

H [HCTK08]. Hack [MLP^{+}20]. Hamming [RSBA23, TT20]. Hardening [LBB^{+}18]. Hardware [ASB^{+}21, ABR^{+}21, ALY^{+}21, ASK23, AHS22, ATW^{+}22, BBB^{+}22, BSY^{+}16, BGX^{+}21, BHLD19, BWL^{+}21, BVM^{+}19, CRSSBMR21, CLKG17, CLSD18, CQL21a, CQL21b, CCWCC15, CKWK18, DBS^{+}21, DRG21, DCSA22, EFRB22, FMP^{+}21, GCC^{+}23, GFZ13, HIH18, AML^{+}23, IGGR19, KPM22, KKY^{+}20, KRG^{+}23, KCD15, KKNM22, LQYL19, LFDS22, LJL18, LCK19, MEHT21, MAC^{+}21, MPR^{+}22, MKG^{+}23, PJSM17, SLC^{+}17, SBR19, SCLW19b, SCLW19a, SLCJ22, WYZ^{+}20, WHB^{+}21, XLW^{+}18, ZLGL21, TMM^{+}07]. Hardware-accelerated [KKNM22]. Hardware-Assisted [MPR^{+}22]. Hardware-Aware [SLCJ22].

Hardware-Based [LQYL19, MAC^{+}21]. Hardware-oriented [AHS22]. Hardware-Software [IGGR19]. Harvesting [RSD^{+}23a, RFDT15, ZSXY11, KP10, MCT10, WOW^{+}10, WCSA10]. Hashing [JCK23]. HBM [ARD24]. HDL [OSLT06]. HDLQ [OSLT06]. HDRLPIM [RAC^{+}24]. Heart [XYJ^{+}21]. Hermitian [AHSZ16, HZSA14]. Heterogeneous [AAFM13, BHLD19, CTP14, HTMH18, IASK20, KHR^{+}15, LGL15, LDP^{+}20, MMD^{+}20, SM19, VK18]. Heterojunction [KCWL^{+}16]. Heuristic [BM15, PT12]. Heuristics [RPBA21]. Hexagonal [KDSD24]. Hiding [FHFK14]. Hierarchical [DBS^{+}21, LFDS22, YXW^{+}12]. Hierarchy [PHS^{+}15, VSRR15]. High [ASP^{+}18, BYHT18, BH17, CJ16, CRKP22, GN08, HRR19, LTKP16, LML^{+}19, LZX^{+}21, MTC^{+}08, PAB^{+}17, PSY^{+}18, PAC^{+}22, SC08, WL22b, XLL^{+}18, ZM22, Bea11, BPH^{+}11, CA11, LMC^{+}11, MN06, PFOL07, RMBC12, WWJ09, ZS08, SM11]. high- [MP10]. high-efficiency [ZS08]. High-level [MTC^{+}08, PAC^{+}22, SC08, ZM22]. High-Performance [ASP^{+}18, BH17, CJ16, CRKP22, PSY^{+}18, HRR19, LML^{+}19, Bea11, BPH^{+}11, LMC^{+}11, WWJ09, SM11]. High-Speed [BYHT18, LTKP16, PAB^{+}17, XLL^{+}18]. high-throughput [CA11, RMBC12]. Higher [ST20]. Highlights [DR11]. Highly [HN12, MGMU22]. Highly-Efficient [MGMU22]. History [Ko12]. HMC [HTMH18]. Holistic [BAT^{+}22]. Homogeneous [HRR19]. Horse [DRG21]. hosting [AMVG12]. HPC [KK23]. HTM [ZK19]. Human [KWWI17]. Human-Scale [KWWI17]. Hut [MPZ21]. HW [JRLR15]. HW/SW [JRLR15]. Hybrid [DJ16, DLL^{+}19, HAV^{+}22, HH11, LNL19, LQYL19, LWM^{+}14, LWY22, NAY24, PHS^{+}15, PAP^{+}22, SMT21, WDW13, YLR^{+}23, YYPK17, YYBK19, CB09, CJ14b, LBGR08, LMC^{+}11, RT07, SCI^{+}09, ZJS09a, ZJS09c, ZJS09b, ZJS10]. Hyper [RAC^{+}24]. Hyper-Dimensional [RAC^{+}24]. Hypercolumn [PSL^{+}19]. Hyperdimensional [SBR19]. Hypervectors [SBR19].

IBM [NPS^{+}20]. IC

[BLKM23, CH14, SMZ^{+}19]. iConn [LGL15]. ICs [ARLB18, BLKM23, CKC^{+}18, CBK^{+}22, FRB08, MRG22, RSBA23, SCL17, TZS14, XS14, ZCB^{+}22]. Identification [KRP^{+}21, SMZ^{+}19, CWT14]. identify [PT12]. Idle [SLC^{+}17]. IEEE [Shu09]. IEEE/ACM [Shu09]. II [ZJS09c]. ILP [ZMT13]. Image [BWL^{+}21, CLZ^{+}22, MGMU22, MCH22, PSR17, LM13]. Images [ATW^{+}22, KBC^{+}23, XZL^{+}21]. IMFlexCom [PAF18]. Impact [BSS16, DNHL11, KAKSP14, KKC17, KMC^{+}22, LZX^{+}21, MGK18]. Implantable [BY12, HLH^{+}12, Ko12, MGS^{+}12, SSN12, SCZ^{+}12]. Implants [MLP^{+}20]. Implementation [VDK^{+}21, CRSSBMR21, JOF^{+}15, LFDS22, LLX^{+}18b, LMM18, PCD^{+}11, SSN12, XLW^{+}18]. Implementations [BVM^{+}19, DBG^{+}14]. Implementing [SMR^{+}12]. Implicant [PT14a]. Implicant-Implicit [PT14a]. Implications [KAB^{+}21, VO06]. Implicit [PT14a]. Improve [HSZM17, OYLL24]. Improved [DRSR14, DRG21, PT14a, PSR17, ZJ11]. Improvement [RBHG21, XZR^{+}22, YMWH21, ANR^{+}14, FRB08]. Improving [AHPC21, CMM^{+}18, MPZ21, ON15, RBHG21, SCL17, YWF18]. In-Field [BMB18]. In-Memory [HFLZ22, HHD^{+}23, NHL^{+}17, PAF18, RAC^{+}24, TT20, KMC^{+}22]. In-storage [ORC^{+}24]. Incoherent [YLR^{+}23]. Incremental [XLW^{+}18]. independent [Tah06]. indoor [WOW^{+}10]. Induced [PAC^{+}22, VDB^{+}16]. Inductor [TZS14].

Inductor-Based [TZS14]. inductors [SXL^{+}12]. Inexact [KT14, MGMU22]. Inference [KWG^{+}20, KKNM22, MZZ23, WWZ^{+}22, ZCDD19]. Information [AAFM13, HLS14, PRV^{+}20, DWL10]. information-theoretic [DWL10]. Infrastructure [LGL15, PB21, AMA^{+}14]. Inherently [ABR^{+}21]. Initialization [KKY^{+}20]. Injection [DDR^{+}16a, JCK23, JWJ^{+}17, LDPPB21, PHS^{+}23]. Inkjet [RHB^{+}20]. INspired [VTKT22, CZQK15, DBS^{+}21, BY18, CQZK14]. instability [KCC^{+}14]. Integer [DBG^{+}14, WL22a]. Integer-arithmetic [WL22a]. integral [KTW08]. Integrated [BS15, ISI^{+}18, KKC17, KHC^{+}22, LKK^{+}22, MMAAK23, NLK^{+}13, TJ13b, TW22, XC08, XLL^{+}18, ZXC10, Bea11, BCT^{+}13, HCTK08, MN06, WFCX09, XCF08]. Integration [AAFM13, KWWI17, MLK^{+}08, SX11]. Integrity [CMZR23, PB21]. Intel [RBG^{+}22]. Intellectual [Bis21]. Intelligence [Dea14, KBM21, XYJ^{+}21]. Intelligence-based [XYJ^{+}21]. Inter [CMK^{+}21, CBK^{+}22]. Inter-Layer [CBK^{+}22]. Inter-tier [CMK^{+}21]. Interaction [CV11]. Interactions [NPA^{+}12]. Interconnect [BPS19, KMC^{+}22, MN06, MTC^{+}08]. interconnection [LJ10]. Interconnects [DDP20, KBT24, LLSO17, LKL^{+}18, Bea11]. Interface [WKL16, XDX14]. interference [XHSC07]. Intermittent [RCYB22]. Intermittently [ZPL^{+}20]. Internet [MPR^{+}22]. Interpolation [ZCSG21]. Interposer [KBT24]. Interpreting [GCB14]. interstitial [SC06]. Intracardiac [ZBF^{+}22]. Introduction [Ano18, AD14, BC08, Bah09, BY12, CLKG17, CLSD18, CQL21a, CQL21b, DR11, DHK^{+}23, Edi14, FSDT23, FMP^{+}21, FNO^{+}19, GMGA23, Gui13, HN15, KK23, KP10, LC08, LZBW20, MPR^{+}22, MSW14, PG12, PR13, SCLW19b, SCLW19a, SX11, SS15, SN11, TTS22, WDT14, WD22]. Intrusion [BYHT18]. Investigating [MLK^{+}08]. Investigation [GCTF20]. IoT [EFRB22, KT20, LYEK22, RSD^{+}23a, RDS24, STSG17, TFZ^{+}21, TW22]. IoT-based [RDS24]. IP [ASMK22, GCC^{+}23, OK22, SMZ^{+}19]. IP/IC [SMZ^{+}19]. Iris [TZR20, LMC^{+}11]. irregular [LDL10]. Irreversible [HLS14]. Issue [BY12, CQL21a, CQL21b, DMYT15, DR11, DHK^{+}23, FSDT23, FMP^{+}21, GMGA23, Gui13, HN15, KK23, LZBW20, MPR^{+}22, MSW14, SS15, SLCJ22, SK16, SN11, TTS22, TSB15, TSMCB17, TV17, WDT14, WD22, ZLWB20, AD14, BC08, CS07, Edi14, McK07, PG12, PR13, XCF08].

JETC [BC08, SLCJ22, TV17, ZLWB20]. JETC/TODAES [BC08]. Job [MNT14]. joint [BC08]. JTAG [PB21]. Junction [VDB^{+}16]. Junctionless [BSS16].

Keeping [ABS^{+}12]. Kernels [LCY19]. Ket [CD22]. Key [BBB^{+}16, CKA23, MKMS22]. Key-Sharing [MKMS22]. Kilobyte [TFZ^{+}21]. Kilocore [ACJ17]. KNN [MPZ21]. Kogge [BSL^{+}18]. Kronecker [TFZ^{+}21].

L2 [PAB^{+}17]. lab [ZXC10]. lab-on-chip [ZXC10]. labeling [EWKNW07]. Laboratories [DHK^{+}23]. Language [CTT^{+}20, MMD^{+}20, OBLD14]. Large [Bea11, KCD15, LGYC21, PDL15, SWK^{+}16]. Large-Scale [KCD15, Bea11, LGYC21]. Laser [DDR^{+}16a, LKL^{+}18]. Lasers [FC18]. Last [NLW^{+}20]. Last-level [NLW^{+}20]. Latency [DKK^{+}15, CA11]. Layer [BS15, VRBS16, CBK^{+}22, DKK^{+}15, KST^{+}22, LLSO17, NVW^{+}22, PRG^{+}15, SS15, ZGSA15, KWC^{+}20]. layout [RMBC12]. LC [SXL^{+}12]. LDPC [LPW18]. Leakage [CMJ14, CJ16, MLW^{+}23, CJ14b, GJ17, LSH14]. Leakage/Delay [CMJ14]. learnable [YYBK19]. Learning [ASB^{+}21, ABM21, ASMK22, ARD24, ATW^{+}22, AAO21, AMF^{+}15, BWL^{+}21, CLKG17, CLSD18, CZQK15, CQL21a, CQL21b, CMZR23, DCP^{+}21, DDG^{+}22, DJH^{+}19, DCSA22, FMTP22, GCC^{+}23, GGTG^{+}20, GGHP^{+}24, HON21, KPPB17, KRG^{+}23, KSA^{+}22, KPFM16, LDZ22, LGYC21, MEHT21, MS19, PSR17, PSY^{+}18, RAC^{+}24, Rez23, RBHG21, SAAR20, SCLW19b, SCLW19a, SLCJ22, SMR^{+}21, SPR18, STNP21, TWLL19, WL22a, YLR^{+}23, ZY18, ZK18, ZK19, CQZK14]. Learning-Based [ZY18]. Learning-Part [CQL21a]. less [HYA^{+}20, RSD^{+}23a]. Level [ARLB18, CCW18, CZW^{+}19, DRL^{+}19, JWJ^{+}17, LLX^{+}18b, LZX^{+}21, LCY19, MTC^{+}08, MP10, Moh12, NLW^{+}20, PAC^{+}22, SC08, ZM22]. Leverage [RPBA21]. Leveraging [PRV^{+}20, SMR23]. library [LCJ14]. life [ZJT^{+}14]. light [WOW^{+}10]. Lightweight [ASMK22, LHHZ19, MKMS22, MAC^{+}21, RDS24]. like [HMS^{+}05]. Limit [LCK19]. Limits [RD22, GMM12]. Linear [ZCSG21]. Links [NVW^{+}22]. Literature [AAFM13]. LiwePMS [LHHZ19]. Loads [ZSXY11]. Localization [CBK^{+}22]. Locally [DNHL11]. Lock [ASMK22]. Locking [LZX^{+}21, SMR^{+}21]. Log [LH20]. Log-Structured [LH20]. Logic [AHSZ16, CJ14a, CJ16, CNH12, GCO^{+}11, GUP11, GVRR17, KT20, LCSP14, LP17, LZX^{+}21, LMM18, MS17, NLL^{+}17, PT14a, SJKS20, SMR23, SSP14, SMR^{+}21, SWK^{+}16, VGZ11, ANR^{+}14, CJ14b, DJ08, HMS^{+}05, LJ14, LCT12, MTC^{+}08, PT12, TR13, ZMT13]. Logic-Based [KT20, NLL^{+}17, TR13]. Loihi [RBG^{+}22]. Long [ZSPC19]. Loop [SSN12]. LosPem [LH20]. Loss [HLS14]. Lotus [LGYC21]. Low [AHHS21, BS20, CJ16, Che15, EFRB22, GBLD15, GLMG^{+}15, HHD^{+}23, IBO22, KZW^{+}15, KHR^{+}15, KR18, LTM22, LNL19, MGS^{+}12, MMJ09, OYLL24, PRG^{+}15, PSL^{+}19, QGW20, RMW^{+}17, SLC^{+}17, STSG17, SGR^{+}12, SM19, SSF^{+}15, Tah09, TSB15, ZJS10, ZJ11, ZPL^{+}20, ABS^{+}12, CJ14b, CA11, ERGK21, GGTG^{+}20, GJ17, KT14, LBGR08, LMC^{+}11, MFA^{+}13, WDH^{+}09]. Low-Cost [LNL19, LBGR08]. low-latency [CA11]. Low-overhead [EFRB22, Tah09]. Low-Power [GBLD15, GLMG^{+}15, KHR^{+}15, KR18, OYLL24, PRG^{+}15, PSL^{+}19, QGW20, RMW^{+}17, SGR^{+}12, AHHS21, BS20, MMJ09, SM19, ZJS10, ABS^{+}12, KT14, LBGR08, LMC^{+}11, WDH^{+}09]. Low-Rank [HHD^{+}23]. Low-Swing [SSF^{+}15]. LTPS [LBGR08].

Machine [ABM21, ARD24, BWL^{+}21, CQL21a, CQL21b, DCSA22, GCC^{+}23, GGTG^{+}20, GGHP^{+}24, HON21, KPPB17, KRG^{+}23, KPFM16, LGYC21, MEHT21, MS19, Rez23, SCLW19b, SCLW19a, YP17]. Machine-Learning [KPFM16]. Machines [MDS21, NPS^{+}20]. Magnetic [ATW^{+}22, LBJ^{+}16, VDB^{+}16, XZR^{+}22, AKW^{+}13, DK09]. Main [YYBK19, YXD^{+}17]. majority [HMS^{+}05]. majority-like [HMS^{+}05]. Making [AGD^{+}20]. Management [CMK^{+}21, DJH^{+}19, HTMH18, LHW^{+}17, LHHZ19, PHS^{+}15, PRG^{+}15, SHB^{+}21, ZJ11, FRB08, GMM12, MP10, MCT10]. manager [AMVG12]. Managing [NPH18]. Manufacturing [CZW^{+}19, MRG22]. Manufacturing-Based [CZW^{+}19]. Many [DMYT15, KDMT22, KPFM16]. Many-Core [DMYT15, KDMT22, KPFM16, Rez23]. Manycore [CMK^{+}21, CRKP22, HM21]. Manycores [PHS^{+}15]. Map [RMG15]. Mapping [FC18, GCO^{+}11, GUP11, KDSD24, MZZ23, TSÁ^{+}24, ZMT13]. March [SN10]. Markov [GD12, KR18]. Matching [RDH14]. MATE [PHS^{+}15]. materials [BPH^{+}11]. Matrix [DD14, GCO^{+}11, GUP11]. Maximum [KAKSP14]. means [VOB19]. Measure [BS15]. Measures [HON21]. mechanical [LWH14, WCSA10]. Mechanism [MRG22, RSD23b]. Mechanisms [LBB^{+}18]. MEDA [HRR19]. MEDA-based [HRR19]. Medical [SLCJ22]. Medium [RSD23b]. Memcapacitive [TT21]. Memories [CCWCC15, HH11, KWC^{+}20, LBJ^{+}16, SBT20, SCZ^{+}12, YYPK17, GRS05, RYT^{+}07]. Memory [AIK21, ANT22, DK21, GNY^{+}22, GLMG^{+}15, GG17, GIS^{+}22, HTMH18, HVB22, HFLZ22, HC15, HHD^{+}23, IASK20, JRJ22, KPPB17, KHR^{+}15, KKKK18, KMS^{+}20, KKK22, KHC^{+}22, LKK^{+}22, LHHZ19, LML^{+}19, LH20, LFDS22, LYWW13, MDS21, MRR12, MSB^{+}22, Mit17, NYL^{+}20, NHL^{+}17, PHS^{+}15, PAF18, RAC^{+}24, RMW^{+}17, RCYB22, SMR23, SBZT20, SBR19, TT20, WWZ^{+}22, WDW13, WRWW17, WHB^{+}21, WZL16, WL22b, YYPK17, YYBK19, YXD^{+}17, ZSPC19, ABS^{+}12, AKW^{+}13, CSKM13, KMC^{+}22, KSG14, PR13, SKRX13, TCSV09, VMNI08, WYZ^{+}20, ZLB^{+}22, ZJS09a]. Memory-bound [KHC^{+}22]. Memory-Centric [KPPB17, NYL^{+}20, GNY^{+}22]. Memory-Efficient [HHD^{+}23]. Memory-Storage [YYPK17]. Memristive [ASB^{+}21, ANT22, CZQK15, GLMG^{+}15, KZL15, MRR12, NAY24, SAAR20, TT20, UMB^{+}18, WKL16, YW13, ZJK22, ZK18]. Memristor [ASP^{+}18, BYHT18, MZZ23, QGW20, ZLB^{+}22, dLBHC22]. Memristor-Based [BYHT18, QGW20, ZLB^{+}22]. Memristor-CMOS [ASP^{+}18]. Memristors [CHA20]. mesh [EWKNW07]. meshless [KTW08]. Metamaterial [NVW^{+}22]. Metamaterial-enabled [NVW^{+}22]. Method [BBB^{+}16, DDP20, DLL^{+}19, GCO^{+}11, MZZ23, PP21, XYM18, YYBK19, ZSXY11, MHM^{+}08].

Methodology

[AHHS21, CMJ14, CH14, MLW^{+}23, CB09]. Methods [CZQK15, TSMCB17, CCTP08]. metric [SMR^{+}12]. Metrics [HSZM17]. MFNW [AM18]. MiC [LCY19]. Micro [ABM21, ZLGL21]. Micro-Architectural [ABM21]. Micro-ring [ZLGL21]. Microarchitectural [GOGCK11]. microarchitecture [MLK^{+}08]. microarchitectures [XCF08]. Microarrays [SKB13]. Microdevices [VMV13]. Microfluidic [BMB18, GCB14, HD14, LBB^{+}18, MPM13, OGB18, DJRM09, DDM^{+}06, RMBC12, RBGC14, SC08, XHSC07, XC08, YYC07, ZXC10]. Microfluidics [GHHW19, CZ05, SC06]. microfluidics-based [CZ05, SC06]. Microscope [ATW^{+}22]. Migration [AIK21, IASK20]. Millimeter [MKW^{+}14, KK12]. Millimeter-Wave [MKW^{+}14]. Million [AVK16]. Million-Qubit [AVK16]. Minimization [CCH16, LJL18]. Minimum [LCSP14]. MINLP [BM15]. Mitigate [ZH20]. Mitigation [NLK^{+}13, WYZ^{+}20, ZLGL21]. Mitigations [ASK23]. Mixed [DRL^{+}19, LPM^{+}19]. Mixed-Criticality [DRL^{+}19]. mixing [RBGC14]. MLC [AM18, LHW^{+}17]. MLC/TLC [AM18]. MN [PHS^{+}15]. MN-MATE [PHS^{+}15]. MNEMOSENE [ZLB^{+}22]. mNoC [PDL15]. Mobile [TWLL19, TSMCB17, YYPK17, WDH^{+}09]. Mode [BSS16, PAF18]. Model [BM15, CCWCC15, FYJ^{+}17, FCR23, MZR^{+}14, MMD^{+}20, REL^{+}22, WHL^{+}21, DLWW08, MHL08, MTC^{+}08, ZC07]. Modeling [FMW^{+}22, KBC^{+}23, LYEK22, LTKP16, MN06, SSN12, SCT^{+}22, TGCJ16, TKBM12, ZM22, ZF15, KCC^{+}14, KSG14, PFOL07]. Models [ASMK22, KMS^{+}20, KCD15, KKNM22, MHW14, WJWM23, FRB08]. Modern [MPZ21]. Modular [FKM22, MHW14]. Modularization [FHFK14]. Modulation [MGK18, NVW^{+}22]. Modulator [LZCX22]. Module [MPM13, LCJ14, ZS08]. Module-Based [MPM13]. Modules [TGCJ16]. Modulo [TSÁ^{+}24]. Molecular [AGR^{+}23, CNHL08, DPB11, GPW^{+}15, PDL15, SCT^{+}22, WDW13, KSG14, KTW08, MHL08, RYT^{+}07]. Molecular-Spin-Qubit [SCT^{+}22]. Money [BMTP24]. Monitoring [AUDS22, EFRB22, MGS^{+}12]. Monolayer [RMW^{+}17]. Monolithic [CKC^{+}18, CMK^{+}21, CBK^{+}22, KKC17, LKK^{+}22, TTS22, ZCB^{+}22, BCT^{+}13, XDX14]. MOS [KZW^{+}15]. Motion [MGMU22]. MRAM [AKW^{+}13, DSB16, NLW^{+}20, PAF18, SFD17, SMT21, STSG17, VDB^{+}16, ZCDD19]. MRAM-Based [VDB^{+}16, ZCDD19]. MRAMs [MFA^{+}13, PFRR17]. MTJ [MS19, MS20]. MTJ-based [MS20]. Muller [LJ14]. Multi [BRZ21, CMZR23, DJH^{+}19, DRL^{+}19, KBT24, LLSO17, LLX^{+}18b, LCY19, PRG^{+}15, Rez23, XZL^{+}21, LM13]. Multi-Abstraction-Level [DRL^{+}19]. Multi-core [DJH^{+}19]. Multi-Cycle-Consistent [XZL^{+}21]. Multi-FPGA [BRZ21]. Multi-Layer [LLSO17]. Multi-level [LCY19]. Multi-Level-Optimization [LLX^{+}18b]. multi-peak [LM13]. Multi-Processors [PRG^{+}15]. Multi-tenant [CMZR23]. Multi-Terabit [KBT24]. Multi/Many [Rez23]. Multi/Many-Core [Rez23]. Multicast [CDP17, VOB19]. Multicast-Enabled [CDP17]. Multichannel [KRP^{+}21]. multicomputer [VMNI08]. Multicore [CDP17, HAV^{+}22, PCD^{+}11, WXW^{+}17, ZMC15, KWFH12, SLS^{+}14]. multidiscipline [Moh12]. Multilayer [HC15, MHM^{+}08, BPH^{+}11]. Multilevel [KST^{+}22, MRR12, VSRR15, CWL^{+}13, FGZ14]. Multiparameter [CJ16, GJ17]. Multipath [RNN^{+}22]. Multiple [DDM^{+}06, FMTP22, HZSA14, MMJ09]. Multiple-Controlled [HZSA14]. Multiplexing [XLL^{+}18]. Multiplication [CRSSBMR21, ERGK21, GP17, HYA^{+}20]. Multiplication-less [HYA^{+}20]. Multiplications [GZZ^{+}21]. Multiplier [GZZ^{+}21, LNL19, LQYL19]. Multipliers [Mog14, SVA^{+}18]. Multiprocessor [YXW^{+}12, CJ14b, GMM12, LWX^{+}14]. multiprocessors [BPH^{+}11, CA11]. multiPULPly [ERGK21]. Multistate [KHR^{+}15]. multiwalled [SXL^{+}12]. MV [TWLL19]. MV-Net [TWLL19]. mVLSI [CKB20]. mW [WOW^{+}10]. MWSR [HAV^{+}22]. MWSR/R [HAV^{+}22]. MWSR/R-SWMR [HAV^{+}22].

N [AM18]. NANA [PDLS06]. Nano [GKT^{+}18, YLF^{+}17, CB09, LDL10, MP10, PDLS06, SCI^{+}09, ZMT13, ZJS10, ZC07, MRR12, ZJS09c, ZJS09a, ZJS09b]. nano-architectures [ZMT13]. nano-CMOS [MP10, SCI^{+}09, ZC07]. Nano-Oscillator-Based [YLF^{+}17]. Nano-Oscillators [GKT^{+}18]. nano-scale [LDL10, PDLS06]. nano/CMOS [ZJS10, ZJS09c, ZJS09a, ZJS09b]. NANOARCH [Bah09]. NANOARCH07 [Shu09]. NANOARCH’09 [DR11]. nanoarchitectures [Tah06, Tah09]. Nanoarray [FGZ14, GRS05]. Nanocavities [EDCL^{+}22]. Nanocomputing [AGR^{+}23, WWG^{+}19, WWJ09]. Nanocrossbar [GUP11]. Nanodevice [GCO^{+}11, WZSC09]. Nanodevice-Based [GCO^{+}11]. Nanodevices [CZQK15]. Nanoelectronic [LZBW20, TSMCB17, YL14]. nanofabrics [DMR06]. Nanolectronic [ZLWB20]. Nanomagnet [CNH12]. Nanomagnetic [VGZ11]. nanometer [CCTP08]. nanometer-scale [CCTP08]. Nanophotonic [BS20, ISI^{+}18, LLSO17, LKL^{+}18, NPA^{+}12, PDL15, XYM18, LMC^{+}11, SX11]. Nanopipelined [PT14b]. Nanoribbon [HC15, KHR^{+}15]. Nanoribbon-CMOS [KHR^{+}15]. Nanoribbons [DDP20]. Nanoscale [BOAC^{+}20, JRC^{+}13, NLK^{+}13, Shu09, CQZK14, EWKNW07, Nar05, RT07, RT08, WZSC09]. nanostructures [HYWA09]. nanosystem [TWL09]. Nanotechnology [KBM21]. Nanotube [GRPT13, HC15, LKK^{+}22, DLWW08, HZY^{+}12, MN06, SXL^{+}12]. Nanotube-Based [GRPT13, MN06]. Nanowire [Deh05, RKM15]. Nanowire-based [Deh05]. nanowires [SRD^{+}06]. Navigating [CTT^{+}20]. NBTI [KCC^{+}14, LSH14, SLS^{+}14, YWH^{+}13]. NBTI-aware [YWH^{+}13]. Near [DK21, GVRR17, KXY16, Mit16, NPA^{+}12, ORC^{+}24, SCL17, TNWD20]. Near- [GVRR17]. Near-Field [NPA^{+}12]. Near-Memory [DK21]. Near-storage [ORC^{+}24]. Near-Threshold [Mit16, SCL17]. negative [KCC^{+}14]. Neighborhood [KDSD24]. NEMS [HN12]. Neoteric [SMR23]. NeRF [LMZZ24]. net [BPB^{+}12, HYA^{+}20, TWLL19, WL22a]. net-zero [BPB^{+}12]. Nets [DRG21]. Network [ABR^{+}21, ASK23, ACM^{+}20, BS21, BKJ19, Bis17, Bis21, CMM^{+}18, CDP17, CCWCC15, DJ16, GMGA23, GCTF20, HFLZ22, HM21, IGGR19, KKKK18, KKK22, LDK^{+}18, LQYL19, LMZZ24, LLX^{+}18b, MZZ23, MKMS22, MSCS19, MCH22, MGST22, MSC^{+}21, OK22, OYLL24, PDL15, PHS^{+}23, PAP^{+}22, RDS24, VK18, VSM19, WL19, XYM18, YMWH21, ZBF^{+}22, ZSPC19, BPH^{+}11, CDG^{+}12, LMC^{+}11, PDLS06, PT14b, WVGP13, ZFT13, YXW^{+}12]. Network-based [ABR^{+}21]. network-enabled [WVGP13]. Network-on-Chip [ASK23, BS21, BKJ19, Bis17, Bis21, CMM^{+}18, DJ16, GMGA23, LDK^{+}18, MSCS19, OYLL24, PDL15, VK18, VSM19, ACM^{+}20, BPH^{+}11, CDG^{+}12, WVGP13, YXW^{+}12]. Network-on-chips [PHS^{+}23]. Networks [AHS17, BS20, VRBS16, BRZ21, BHLD19, BVM^{+}19, CKC^{+}18, CQZ^{+}21, CTP14, DK21, Dea14, DEW^{+}23, ERGK21, FNO^{+}19, JCK23, KST^{+}22, KRP^{+}21, KCD15, KMC^{+}22, LPB^{+}15, LZXL22, LLX^{+}18a, LNL19, LDP^{+}20, LJL18, LPM^{+}19, MS20, NAY24, ORC^{+}24, PJSM17, PSL^{+}19, PSY^{+}18, Rez23, RBHG21, RBG^{+}22, SBZT20, SM19, SPR18, TZR20, TT21, XLW^{+}18, XZL^{+}21, ZCDD19, ZH20, GD12, LJ10, LDL10, LWX^{+}14, XPD12]. Networks-on-Chip [VRBS16, CTP14, FNO^{+}19, KST^{+}22, Rez23]. Neural [ABR^{+}21, AHS17, BRZ21, BHLD19, BVM^{+}19, CKC^{+}18, CThG15, CQZ^{+}21, CCWCC15, DK21, DEW^{+}23, ERGK21, HFLZ22, HM21, IGGR19, JCK23, KRP^{+}21, KKKK18, KKK22, KCD15, KMC^{+}22, LZXL22, LLX^{+}18a, LNL19, LQYL19, LDP^{+}20, LMZZ24, LLX^{+}18b, LJL18, LYL19, LPM^{+}19, MZZ23, MS20, MSC^{+}21, NAY24, ORC^{+}24, OK22, PAP^{+}22, PSL^{+}19, RBG^{+}22, SBZT20, SVA^{+}18, SM19, SPR18, WL19, XLW^{+}18, YMWH21, ZH20, ZBF^{+}22]. Neuro [CZQK15, CQZK14]. Neuro-Inspired [CZQK15, CQZK14]. Neuroevolution [AAO21]. Neuroevolutionary [SMR^{+}21]. Neurogenesis [ZK19]. Neuromemrisitive [ZK19]. Neuromemristive [KSA^{+}22]. Neuromorphic [ASB^{+}21, AMH^{+}24, Ano18, AMF^{+}15, BOAC^{+}20, HN15, JRJ22, KZL15, KCD15, MKSW17, PSY^{+}18, RMG15, SAAR20, SHB^{+}21, SPR18, WKL16, ZWL^{+}15]. Neuron [HIH18]. Neurons [VOB19, ZCDD19]. Neutral [LPB^{+}15]. Next [GFZ13, KK23, XCS^{+}19]. Next-Generation [KK23, XCS^{+}19]. NFIC [AGD^{+}20]. NML [DNHL11]. NN [ASMK22]. NN-Lock [ASMK22]. NoC [ACJ17, AGD^{+}20, DRL^{+}19, HAV^{+}22, KYEB15, MKW^{+}14, VOB19, ZF15]. NoC-Based [ZF15]. NoCArc [GMGA23]. NoCs [FC18, GB18, LKC15, LCK19, RSD23b, SHAC19, ZY18, ZLGL21]. Node [PHS^{+}15, GGTG^{+}20, YWH^{+}13]. Nodes [GVRR17, LWM^{+}14, TGCJ16]. Noise [CLZ^{+}22]. Noise-Tolerant [CLZ^{+}22]. Noisy [DDG^{+}22, SCT^{+}22]. Non [GKT^{+}18, MCT18, RCYB22, STSG17, WL22b, YYPK17, YXD^{+}17, ZPL^{+}20]. Non-Boolean [GKT^{+}18]. Non-Restoring [MCT18]. Non-Volatile [STSG17, WL22b, YYPK17, YXD^{+}17, ZPL^{+}20, RCYB22]. Nonhierarchical [PPM^{+}13]. Noninvasive [TFR23]. Nonlinear [KKKK18]. Nonvolatile [HC15, LBJ^{+}16, SCZ^{+}12, SKRX13]. NORM [RCYB22]. Note [WWG^{+}19]. Novel [DJ16, FCR23, JDPH^{+}23, LH20, MGMU22, SKB13, TZS14, ZSXY11, RT08]. NP [WWG^{+}19]. NP-complete [WWG^{+}19]. NTC [CV12]. NTRU [CRSSBMR21]. NUCA [PAB^{+}17]. Number [HH11, LTKP16, WL19]. Numbers [GIS^{+}22]. NxTF [RBG^{+}22].

Obfuscation [CZW^{+}19, GCTF20]. Object [KBC^{+}23]. Objectives [DWK^{+}16]. Observing [TGCJ16]. OFDM [GLL^{+}21a]. Off [KK23, FSDT23, ZFT13]. Off-Chip [KK23, WXW^{+}17, ZFT13]. Offline [BMTP24, LKL^{+}18, MT14]. offs [CDG^{+}12]. Oligo [VTKT22]. On-Chip [CLKG17, CZQK15, CKC^{+}18, KDMT22, KK23, LWM^{+}14, LWY22, SCLW19a, TZS14, WM24, XYM18, CLSD18, CWL^{+}13, KMC^{+}22, LWX^{+}14, SCLW19b, CA11, LMC^{+}11]. On-Device [ZK19]. On-Interposer [KBT24]. on-node [GGTG^{+}20]. On-the-fly [IASK20]. On/Off [WXW^{+}17]. On/Off-Chip [WXW^{+}17]. One [HSZM17, WRWW17]. One-Step [WRWW17]. One-Way [HSZM17]. Online [AUDS22, LJL18, ZPL^{+}20]. Open [BBR^{+}23, FCR23]. OpenQL [KAS^{+}22]. Operation [MPM13]. Operations [CVK15]. Optical [BPS19, VRBS16, DWK^{+}16, GB18, HAV^{+}22, JDPH^{+}23, NPA^{+}12, PAB^{+}17, RDM^{+}21, XLL^{+}18, YLR^{+}23, YXW^{+}12, ZY18, ZLGL21, CA11, EDCL^{+}22]. Optical-Electrical [YLR^{+}23]. Optical-Electronic [YXW^{+}12]. Optimization [DKK^{+}15, DWK^{+}16, LDK^{+}18, LLX^{+}18b, LCY19, LKL^{+}18, NLW^{+}20, PFRR17, Rez23, TGCJ16, YJ18, ZY18, ZGSA15, DLWW08, KKY^{+}20, LWH14, WFCX09, ZJS09c]. Optimizations [SBR19, CWL^{+}13]. Optimize [DJ16]. Optimized [CRSSBMR21, CCWCC15, KKKK18, MS17, MCT18, ON15, WGY21, WL22b]. Optimizing [WL22a, TR10]. Optoelectronic [PAP^{+}22]. Oracle [GCJ17]. Order [CKWK18]. Ordering [WGY21]. Organizing [DK09, RMG15, LDL10, PDL07]. oriented [AHS22]. Oscillation [MSB^{+}22]. Oscillation-Based [MSB^{+}22]. Oscillator [FYJ^{+}17, HON21, YLF^{+}17, SXL^{+}12, ZFT13]. Oscillator-Based [FYJ^{+}17]. Oscillators [GKT^{+}18]. Outlook [BWL^{+}21]. Outperforming [LLX^{+}18a]. outputs [TR10]. Overhead [LTM22, SLC^{+}17, ZJ11, ZPL^{+}20, EFRB22, Tah09]. Overlay [PAB^{+}17]. Overview [AML^{+}23].

p [DPB11]. p-QCA [DPB11]. Packet [BYHT18]. Packets [Bis17]. Page [AIK21, IASK20, KXY16]. PANE [VSM19]. Papers [SN10]. paradigm [LBGR08, WZSC09]. Parallel [BAT^{+}22, Dea14, FSD22, ISI^{+}18, DJRM09, STA^{+}12]. Parallelism [DK21, JOF^{+}15, GN08]. parameter [RYT^{+}07]. Parameterized [REL^{+}22]. Parametric [FRB08]. Parasitic [ZH20]. Parasitics [KCWL^{+}16]. Part [CQL21a, CQL21b, SCLW19a, LZBW20, ZJS09c, ZJS09b, ZLWB20]. Partial [KKK22, LKC15, MEHT21]. Partial-Sum [KKK22]. Partially [AGD^{+}20]. Parties [BMTP24, FMTP22]. Partitioning [LRN05, PP21]. Passive [GSC17]. Path [Bis21, CKWK18, GB18, VAK18, WRWW17, CWT14, ZXC10]. Path-based [Bis21]. Path-Setup [GB18]. Pathology [CLZ^{+}22]. paths [ANR^{+}14]. Pathways [KBT24]. Pattern [HSZM17, NLW^{+}20]. Pauli [HZSA14]. PCA [AHPC21]. PCB [JDPH^{+}23, KBC^{+}23, MLP^{+}20]. PCM [KXY16, LHW^{+}17]. peak [LM13]. PEAL [AHHS21]. Peres [DJ08]. Performance [AVK16, AHS22, ASP^{+}18, BSS16, BH17, CDG^{+}12, CKC^{+}18, CJ16, CRKP22, DNHL11, HTMH18, KDMT22, LYWW13, MKW^{+}14, ON15, PSY^{+}18, RBHG21, SCL17, VAK18, VOB19, YWF18, Bea11, BPH^{+}11, DLWW08, HRR19, LMC^{+}11, LML^{+}19, LCT12, MN06, PFOL07, RT07, STA^{+}12, WWJ09, SM11]. performance-aware [STA^{+}12]. Performance-Temperature [HTMH18]. Permutation [GCTF20]. Persistent [KMS^{+}20, LHHZ19, LH20, WL22b]. Perspective [RSBA23]. Perspectives [BWL^{+}21, TTS22]. Phase [FYJ^{+}17, RNN^{+}22, SMT^{+}17, WZL16, JRC^{+}13]. Phase-Based [RNN^{+}22]. Phase-Change [WZL16]. Photonic [ACJ17, BS21, BPH^{+}11, CDP17, DEW^{+}23, EDCL^{+}22, FC18, KST^{+}22, KBT24, VK18]. Photonics [DCP^{+}21, STNP21, XNK18, Bea11]. Photonics-Based [DCP^{+}21]. Physical [DCSA22, LBB^{+}18, NVW^{+}22, TMG^{+}21, UMB^{+}18, BCT^{+}13, HZY^{+}12]. Physically [LTM22]. PicoServer [KSB^{+}08]. Piezoelectric [RFDT15]. pillar [MFA^{+}13]. PIM [REL^{+}22]. Pin [WXW^{+}17, XHSC07]. pin-constrained [XHSC07]. Pipeline [SM11]. Piracy [SMZ^{+}19]. PLA [CNH12].

Placement

[BM15, VRBS16, BKJ19, CKB20, LWH14, RHB^{+}20, WWG^{+}19, YYC07, LRN05]. plane [LDZ22]. PLAs [CHN09]. Plasticity [AMF^{+}15, LYL19, WLJC21]. Platform [ZM22]. Pluggable [VSM19]. Point [LYWW13, NV14]. Pointing [LBJ^{+}16]. Points [AUDS22]. Polarity [MGZ^{+}17]. Policies [AIK21, ON15]. Polyhedral [HFLZ22]. Polyhedral-Based [HFLZ22]. Polynomial [CRSSBMR21]. Polynomials [LP17]. Pooling [ZMC15]. portability [GN08]. Portable [KAS^{+}22]. Post [CKA23, GCC^{+}23, MRG22, XS14]. post-bond [XS14]. Post-Manufacturing [MRG22]. Post-processing [GCC^{+}23]. Post-quantum [CKA23]. Potential [SJKS20]. Power [BS21, CKC^{+}18, CMK^{+}21, Che15, DJH^{+}19, FC18, GBLD15, GLMG^{+}15, HN12, JRLR15, KZW^{+}15, KHR^{+}15, KR18, LBJ^{+}16, LHW^{+}17, LTO22, LWM^{+}14, LKL^{+}18, MGS^{+}12, MGST22, MSC^{+}21, OYLL24, PRG^{+}15, PSL^{+}19, QGW20, RMW^{+}17, STSG17, SGR^{+}12, TGCJ16, TW22, TSB15, WXW^{+}17, ZJ11, ZSXY11, ZY18, ZGSA15, ZF15, ABS^{+}12, ANR^{+}14, AHHS21, BS20, ERGK21, GMM12, GGTG^{+}20, KT14, KK12, LJ10, LBGR08, LMC^{+}11, MMJ09, MP10, MFA^{+}13, SM19, WDH^{+}09, XDX14, ZS08, ZJS10, ZFT13]. Power-based [LTO22]. Power-efficient [BS21, MSC^{+}21, ANR^{+}14]. Power-Gating [HN12, ZF15]. Power-Utility-Driven [LHW^{+}17]. Powered [JRLR15, ZPL^{+}20, WCSA10]. Powerful [VMV13]. PPU [GYM^{+}17]. Practical [GGHP^{+}24, HAV^{+}22]. Pre [FSDT23, XS14]. pre-bond [XS14]. Pre-silicon [FSDT23]. Precise [WDG^{+}20]. Precision [HM21, LWY22]. Predict [BS20]. Prediction [MKSW17]. Predictions [SMT21]. Predictive [DKK^{+}15, ZC07]. Prefetching [YYBK19]. Pressure [MGS^{+}12]. Prevent [ASMK22].

Prevention [MLP^{+}20]. Primitive [GRPT13]. Primitives [ABR^{+}21, BSY^{+}16, HMS^{+}05]. Principal [AHPC21]. Printed [PB21, RHB^{+}20]. Privacy [FMTP22]. Proactive [PRG^{+}15]. Probabilistic [AHHS21, KSG14, ZCDD19, KT14]. Probability [VAK18]. Probes [SKB13]. Probing [RDM^{+}21]. problem [EWKNW07]. Problems [AAO21]. Process [CMK^{+}21, GPW^{+}15, KAKSP14, LKK^{+}22, MGK18, SCL17, XYM18, ZM22, XPD12]. Process-Variation-Tolerant [XYM18]. Processing [AL17, BH17, CLZ^{+}22, KWG^{+}20, KHC^{+}22, MGMU22, RAC^{+}24, XYJ^{+}21, XCS^{+}19, dLBHC22, GCC^{+}23, Gla14, KT14, LM13]. processing-in-wire [Gla14]. Processor [GKT^{+}18, GYM^{+}17, KRP^{+}21, KZL15, Mit17, STSG17, WXW^{+}17, YJ18]. Processors [HAV^{+}22, KAKSP14, KAB^{+}21, PRG^{+}15, SLC^{+}17, WKL16, ZMC15, ZWL^{+}15]. productivity [SMR^{+}12]. Products [TFZ^{+}21]. Profiled [RBHG21]. Programmable [AMF^{+}15, DPB11, Deh05, WDH^{+}09]. Programming [CD22, KAS^{+}22, KMS^{+}20]. Programs [WWC23]. Project [TMM^{+}07]. Projection [DLTSNA20]. Promises [YW13]. Propagation [BMTP24]. Propagation-of-Provenance [BMTP24]. Property [Bis21]. Prospect [PFOL07]. Prosthesis [SSN12]. Protecting [LCK19]. Protection [Bis21]. Protein [PPM^{+}13]. Protocol [RDS24, YWF18]. Protocols [BBR^{+}23]. PROTON [VRBS16]. Provenance [BMTP24]. Provide [SLC^{+}17]. Providing [KMS^{+}20]. ProWATCh [PRG^{+}15]. Pruning [AHS17]. PUF [AHPC21, BMTP24, DSB16, FMTP22, KT20, LYEK22, MKMS22, RDS24, TMG^{+}21, VDB^{+}16, XZR^{+}22]. PUF-based [BMTP24, LYEK22]. PUF-Cash [FMTP22].

PUFs [AHS22, HON21, IGR^{+}16]. PVFS [JOF^{+}15]. PVT [CMJ14, CJ14a, CJ15, TGCJ16, YJ18].

QCA [CNHL08, CHN09, DPB11, DK09, Gla14, GPW^{+}15, HMS^{+}05, LRN05, MHL08, OSLT06, SWJ07, SDSS14, TCSV09]. QCA-based [CHN09]. QCOR [MMD^{+}20]. QLib [LCJ14]. QPSK [LZCX22]. Quality [AHPC21, MEHT21]. Quality-assured [MEHT21]. Quantifying [HLS14, NPH18]. Quantitative [SPS^{+}24]. Quantization [CQZ^{+}21, KKK22, XLW^{+}18]. Quantized [LNL19]. Quantum [AUDS22, AVK16, AHSZ16, ATW^{+}22, BM15, BH17, CD22, CVK15, CV11, DD14, DPB11, GCJ17, HZSA14, HSZM17, KAS^{+}22, KDSD24, LCSP14, LCJ14, MMD^{+}20, MCT18, NV14, PSY^{+}18, RDH14, SCT^{+}22, ST20, TNWD20, VAK18, WD22, CKA23, CV12, DWL10, MTC^{+}08, TR10, VO06, VMNI08, WDH^{+}09]. Quantum-Classical [MMD^{+}20]. Quantum-Dot [DPB11, TNWD20, DWL10]. Quantum-Logic [AHSZ16]. Quasi [LZCX22]. Quasi-digital [LZCX22]. Quaternary [SJKS20]. Qubit [AVK16, KDSD24, MCT18, NPS^{+}20, SCT^{+}22]. QuickRecall [JRLR15]. Quit [WGY21]. QuTiBench [BHLD19].

R-SWMR [HAV^{+}22]. Racetrack [KWC^{+}20]. Radial [SRD^{+}06]. Radio [CFK22, EFRB22]. Radio-Frequency [CFK22]. Radix [GUP11]. Rail [DLTSNA20]. RAM [CWL^{+}13, LPW18, RKM15, ZCX^{+}17]. Ramifications [JMKM21]. Random [KR18, LTKP16, LBJ^{+}16, MPZ21, AKW^{+}13, CSKM13, SWJ07]. Randomized [SHAC19]. Randomly [CThG15]. Ranging [RNN^{+}22]. Rank [HHD^{+}23]. Rate [QGW20]. Ray [KBC^{+}23]. RCG [ZBF^{+}22]. Read [WRWW17, ZCX^{+}17]. Real [DRL^{+}19, GSC17, JWJ^{+}17, KR18, KPFM16, SPS^{+}24, SM19, TWLL19, ZBF^{+}22, LWX^{+}14]. Real-Time [DRL^{+}19, GSC17, JWJ^{+}17, KR18, KPFM16, TWLL19, SM19, ZBF^{+}22, LWX^{+}14]. Realization [BSL^{+}18, PP21]. Realizing [SDSS14]. reasonable [CNHL08]. Recognition [CKC^{+}18, KSA^{+}22, PSR17, TZR20]. Reconciliation [IASK20]. Reconfigurable [CEW^{+}13, CCH16, CDP17, CNH12, GZZ^{+}21, KPPB17, KZL15, NLL^{+}17, VK18, WHL^{+}21, EWKNW07, Sek07, SC06, Tah06, ZJS09a, ZJS09c, ZJS09b, ZJS10]. Reconfiguration [MEHT21]. Reconstruction [ZBF^{+}22]. Recorder [SCZ^{+}12]. Recovery [KCC^{+}14, Sek07, ZXC10]. Recurrent [LZXL22, LJL18, SBZT20]. Recycle [BS20]. REDELF [LKC15]. Redesign [YXD^{+}17]. Reduced [SGR^{+}12]. Reducing [FC18, LBJ^{+}16]. reduction [LSH14]. Reductions [MMAAK23]. Redundancy [ABR^{+}21, WDW13, SC06, WWJ09]. Redundant [HH11]. Reed [LJ14]. Reference [AMH^{+}24, MGK18]. Reflection [MGST22]. ReG [LDP^{+}20]. Register [CZW^{+}19, WX15, ZCX^{+}17, TCSV09]. Regular [DDR^{+}16b]. Regularization [YMWH21]. Regulators [IBO22]. Reinforcement [AAO21, DJH^{+}19, FMTP22, RAC^{+}24]. rejuvenation [AMA^{+}14, CNPR14, MNT14, ZJT^{+}14]. Relativistic [MJ11]. Release [HLH^{+}12]. Release-on-Demand [HLH^{+}12]. Reliability [ANR^{+}14, BAT^{+}22, HCTK08, KYEB15, LYWW13, LBB^{+}18, SHB^{+}21, CWT14, DJH^{+}19, DK09, Edi14, TMM^{+}07]. Reliable [CKA23, KT20, XZR^{+}22, McK07, WWJ09]. Rematerialization [SBR19]. Rendering [LMZZ24]. Repercussions [SPS^{+}24]. Replacement [ON15]. replicating [TMM^{+}07]. Representation [BBR^{+}23, WL19]. ReRAM [BOAC^{+}20, FSD22, KKY^{+}20, LKK^{+}22, LDP^{+}20, WL19]. ReRAM-Based [WL19, KKY^{+}20, LDP^{+}20]. ReRAMs [MSB^{+}22]. Research [WWG^{+}19]. Reservoir [BY18, LMM18]. Residue [HH11, PAP^{+}22]. Resilience [ACM^{+}20]. Resiliency [SFD17, WWZ^{+}22]. Resilient [ALY^{+}21, LCT12, LZX^{+}21, OK22, RDM^{+}21, RSD^{+}23a]. Resistance [ZJ11, ZH20]. Resistant [VDK^{+}21]. Resistive [BSL^{+}18, DSB16, KKKK18, WRWW17, ZH20]. resonant [LM13]. Resonator [ZLGL21]. Resource [AVK16, NV14, OGB18, PHS^{+}15, TZR20, TSÁ^{+}24, TMG^{+}21]. Resource-Constrained [OGB18, TSÁ^{+}24, TMG^{+}21]. Resource-Efficient [NV14, TZR20]. Response [CMJ14, MGST22]. Restoring [MCT18]. Restricted [YP17]. Resulting [SDSS14]. Retail [KK12]. Rethinking [WZL16]. Retrieval [BBB^{+}16]. Reuse [CH14]. Reverse [BWL^{+}21, CCW18, QCF^{+}16]. Reversible [DRSR14, DBG^{+}14, DDR^{+}16b, DJ08, HM14, LCSP14, Mog14, NV14, SMZ^{+}19, SZSS10, SDSS14, SSP14, SWK^{+}16, WDT14, CW08, LJ14, PSM^{+}06, TR10, TR13]. Review [JLL^{+}17, PYSJ22]. RF [CFK22, BSS16]. RF/Analog [BSS16]. RIMEP2 [HM14]. Ring [HON21, ZFT13, ZLGL21]. Ripple [MGZ^{+}17, RSD^{+}23a]. Ripple-Carry [MGZ^{+}17]. Ripple-less [RSD^{+}23a]. RMDDS [LJ14]. RNNFast [SBZT20]. RNNs [TFZ^{+}21]. RO [AHPC21]. RO-PUF [AHPC21]. Robust [ASB^{+}21, AMH^{+}24, BMB18, CQZK14, GRPT13, GJ17, LZX^{+}21, LDZ22, MGK18, NAY24, PPM^{+}13, SAAR20, CB09, WVGP13]. Robustness [BS15]. Root [MCT18]. Router [Bis17, DRL^{+}19, KPFM16, CA11]. Routing [VRBS16, DDP20, HRR19, LKC15, LDL10, MKW^{+}14, PHS^{+}23, RHB^{+}20, VOB19, WWG^{+}19, RT07, XC08]. RRAM [JRJ22, NHL^{+}17]. RRAMs [FMW^{+}22]. RT [ZBF^{+}22]. RT-RCG [ZBF^{+}22]. RTL [LDPPB21, WJWM23]. Rule [OBLD14]. Rule-Based [OBLD14]. runtime [GMM12].

SaARSP [LZXL22]. SABER [VDK^{+}21]. SAF [MZZ23]. Safety [SPS^{+}24]. Sampling [QGW20]. SAT [TSÁ^{+}24]. SAT-Based [TSÁ^{+}24]. Saviors [ABM21]. Scalable [BPS19, BS21, Che15, DBS^{+}21, GLL^{+}21b, GB18, MT14]. Scale [KCD15, KWWI17, PDL15, Bea11, CCTP08, LDL10, LGYC21, PDLS06]. scaled [LBGR08]. Scaling [BSS16, JOF^{+}15, LYWW13, WCSA10]. SCALPEL [RD22]. Scan [RSBA23, WFCX09, HCTK08, XS14]. Scan-based [RSBA23]. Scan-chain [WFCX09]. ScatterVerif [MGST22]. Scheduling [BM15, MSCS19, OGB18, SM19, TSÁ^{+}24, STA^{+}12, ZJT^{+}14]. Scheme [CHA20, GLMG^{+}15, GB18, MKMS22, MGK18, MSC^{+}21, WRWW17, XS14]. Schemes [NVW^{+}22, SMR^{+}21, GD12]. SCKVdd [Che15]. Scoring [AAFM13]. SCT [RT08]. Search [ZBF^{+}22]. Searches [MT14]. Secret [BBB^{+}16]. Section [SCLW19b, Bah09, LC08, Moh12]. Secure [ALY^{+}21, CFK22, FKM22, MKMS22, RDS24, SK16]. Securing [EFRB22, PHS^{+}23, RSBA23]. Security [AHS22, BSY^{+}16, BGX^{+}21, EDZ^{+}23, FSDT23, FMP^{+}21, GSC17, IGR^{+}16, KRG^{+}23, LTM22, LYEK22, MAC^{+}21, MPR^{+}22, NVW^{+}22, PRV^{+}20, PAC^{+}22, RNN^{+}22, SMR^{+}21, WYZ^{+}20, WJWM23]. Segment [KXY16]. Segment-Based [KXY16]. Segmentation [CLZ^{+}22, MCH22, WL22a]. Seizure [SGR^{+}12]. Self [CBK^{+}22, Che15, DLL^{+}19, IBO22, LCK19, MRG22, MGK18, RMG15, TMM^{+}07, YYBK19, LDL10, PDL07, WCSA10].

Self-aware [DLL^{+}19]. Self-learnable [YYBK19]. Self-Organizing [RMG15, LDL10, PDL07]. Self-Protecting [LCK19]. Self-Reference [MGK18]. Self-replicating [TMM^{+}07]. Self-Stabilized [Che15]. Self-Test [CBK^{+}22, MRG22, IBO22]. self-timed [WCSA10]. Semantic [JDPH^{+}23]. Semi [ZK18]. Semi-Trained [ZK18]. Sense [SFD17]. Sensing [MGK18, ZLGL21]. Sensitive [ZY18]. Sensitivity [LDK^{+}18, WDG^{+}20]. Sensitivity-based [LDK^{+}18]. Sensor [LPB^{+}15, LWM^{+}14, MKMS22, MGS^{+}12, RDS24, GD12, LWX^{+}14, WOW^{+}10]. Sensors [GGTG^{+}20]. Separable [HYA^{+}20, KRP^{+}21]. Separation [KR18]. Sequence [PPM^{+}13]. Sequences [MT14]. Sequential [NLL^{+}17, CW08, TR10]. SerDes [XLL^{+}18]. serial [Gla14]. Series [MKSW17]. server [MNT14]. servers [ABS^{+}12, KSB^{+}08]. Setup [GB18]. SFQ [GGHP^{+}24, PP21]. Share [BS20]. Shared [HAV^{+}22, PAB^{+}17, VK18]. Sharing [CDP17, MKMS22, SSP14]. SHARP [VK18]. Shielding [ZCX^{+}17]. Shift [VSRR15, TCSV09]. Shift-Based [VSRR15]. shift-register-based [TCSV09]. shifter [MP10]. Short [ZSPC19]. Short-Term [ZSPC19]. ShuntFlowPlus [GLL^{+}21b]. Si [HCTK08, KAB^{+}21]. Side [ABM21, ASK23, VDK^{+}21, DDG^{+}22, DCSA22, MLW^{+}23, PYSJ22, PRV^{+}20, RPBA21, RBHG21]. Side-Channel [ABM21, PRV^{+}20, ASK23, DDG^{+}22, DCSA22, MLW^{+}23, RPBA21, RBHG21]. Side-Channel-Resistant [VDK^{+}21]. Sign [FSDT23]. Sign-off [FSDT23]. Signal [LPM^{+}19]. Signaling [KST^{+}22]. Signals [YLR^{+}23]. Signature [TW22, XZR^{+}22]. Signatures [DDG^{+}22]. Signed [GZZ^{+}21]. Silicon [CDP17, KDMT22, MLW^{+}23, STNP21, TZS14, XNK18, BPH^{+}11, FSDT23].

Silicon-correlated [MLW^{+}23]. Silicon-Photonic [CDP17]. SILVerIn [PB21]. SIMD [PDL07]. Similarity [KSA^{+}22]. Simple [DDR^{+}16b]. Simplified [FYJ^{+}17]. simplifying [PSM^{+}06]. Simulated [MPZ21]. Simulation [CJ14a, CJ15, FYJ^{+}17, GHHW19, KKNM22, MLW^{+}23, SWJ07, YLF^{+}17, FGZ14]. Simulation-based [KKNM22]. Simulator [AVK16, RAC^{+}24, VAK18, VSM19, ZLB^{+}22, HYWA09, LJ10]. Simulators [ZF15, KCC^{+}14]. Single [VRBS16, CEW^{+}13, CCH16, WHL^{+}21, SXL^{+}12]. Single-Electron [CEW^{+}13, CCH16, WHL^{+}21]. single-walled [SXL^{+}12]. Situ [MS19, ZK18]. sizing [LSH14]. Sketching [AL17]. Skew [NPA^{+}12]. SkyBridge [BLKM23]. Skyrmion [KWC^{+}20]. Slicing [LDZ22]. Small [MKW^{+}14]. Small-World [MKW^{+}14]. SNE [MPZ21]. Sneak [WRWW17]. Sneak-Path [WRWW17]. SNN [KWG^{+}20]. SoC [GSC17, HLH^{+}12]. soft [LWX^{+}14, SKRX13]. soft-error [LWX^{+}14]. Software [ASK23, AMA^{+}14, BGX^{+}21, IGGR19, KKY^{+}20, LDPPB21, WZL16, ZJT^{+}14, CNPR14, MNT14]. Software-driven [BGX^{+}21]. Solar [RSD^{+}23a]. Solution [BAT^{+}22, BM15, KHC^{+}22]. Solutions [FSDT23, FMP^{+}21, LYEK22, LCK19, Rez23, VGZ11, ZCB^{+}22, MN06, Moh12]. solver [KTW08]. Sort [GUP11]. Sorting [ANT22, MSC^{+}21]. SOT [PAF18]. SOT-MRAM [PAF18]. Source [Bis17, KR18]. Sources [BGX^{+}21]. Space [LDK^{+}18, ZJK22, TJ13a, XLBB06, ZJS09a]. SPARCNet [PJSM17]. Spare [BKJ19]. Sparse [HIH18, HYA^{+}20, PJSM17, RMG15]. Sparsification [PSL^{+}19]. Sparsity [YMWH21]. Spatial [KWFH12, LTO22]. Spatio [MMAAK23]. Spatio-Temporal [MMAAK23]. Special [BY12, CQL21a, CQL21b, DMYT15, DR11, DHK^{+}23, FSDT23, FMP^{+}21, GMGA23, Gui13, HN15, KK23, LZBW20, Moh12, MPR^{+}22, MSW14, SCLW19b, SS15, SLCJ22, SK16, SN11, TTS22, TSB15, TSMCB17, TV17, WD22, ZLWB20, AD14, BC08, Bah09, CS07, Edi14, LC08, McK07, PG12, PR13, WDT14, XCF08]. specialized [BC08]. Specific [DKK^{+}15, FC18]. Specification [MMD^{+}20, OBLD14]. spectrally [KTW08]. Speech [CKC^{+}18]. Speed [BYHT18, LTKP16, PAB^{+}17, XLL^{+}18]. Speedup [KAKSP14]. SPICE [KCC^{+}14]. SPICE-compatible [KCC^{+}14]. Spike [AMF^{+}15, LYL19, MSC^{+}21, ZWL^{+}15]. Spike-Time-Dependent [ZWL^{+}15]. Spike-Timing-Dependent [AMF^{+}15]. Spike-timing-dependent-Plasticity [LYL19]. Spiking [ASB^{+}21, BVM^{+}19, CThG15, HIH18, JRJ22, KCD15, LZXL22, LMZZ24, LJL18, LYL19, MSC^{+}21, NAY24, RBG^{+}22, SPR18, VOB19, LMZZ24]. Spiking-NeRF [LMZZ24]. Spin [AKW^{+}13, MS17, SCT^{+}22, VSRR15, WYZ^{+}20, YLF^{+}17, CSKM13, CWL^{+}13, EWKNW07, MFA^{+}13, XZR^{+}22]. Spin-based [WYZ^{+}20]. Spin-Torque [YLF^{+}17]. Spin-transfer [AKW^{+}13, XZR^{+}22, CWL^{+}13, MFA^{+}13]. spin-transfer-torque [CSKM13]. spin-wave [EWKNW07]. SpiNNaker [PCD^{+}11]. Spintronic [AMH^{+}24, IGR^{+}16, SBT20]. Spintronic-based [AMH^{+}24]. Spintronics [KZW^{+}15]. Split [CZW^{+}19]. Splitters [DWK^{+}16]. Square [MCT18]. SRAM [GJ17, KHC^{+}22, RKM15, TFR23]. SRAMs [RMW^{+}17]. SSS [DLL^{+}19]. Stabilized [Che15]. stacked [KWFH12, MHM^{+}08, SKRX13, ZS08]. stacked-Vdd [ZS08]. stacking [KSB^{+}08, MHM^{+}08]. Stage [ZGSA15]. Stand [RFDT15]. Stand-By [RFDT15]. Standard [CMJ14, KCWL^{+}16, MS17].

STAP [dLBHC22]. State [MDS21, ABS^{+}12]. Static [DLL^{+}19]. Static-dynamic [DLL^{+}19]. Statistical [LTKP16, YJ18]. STDP [SAAR20, SPR18]. STDP-based [SAAR20, SPR18]. Stealthy [CMZR23]. Step [WRWW17]. STIFT [MMAAK23]. Stitch [MDS21]. Stochastic [ACH^{+}17, AL17, EDCL^{+}22, GIS^{+}22, KKNM22, LNL19, LQYL19, LP17, LMM18, MZR^{+}14, MKSW17, MS19, MS20, NLL^{+}17, NPH18, PHS^{+}23, ZCDD19]. stochastically [GRS05]. Stone [BSL^{+}18]. Storage [SCZ^{+}12, VSRR15, VTKT22, YYPK17, ORC^{+}24]. strain [LWH14]. Strategies [SFD17, FRB08, GRS05]. Strategy [MSCS19]. Stream [GLL^{+}21b]. Streaming [GYM^{+}17, KR18, QGW20]. Streams [GGTG^{+}20]. stretching [MRH12]. Structure [DDR^{+}16b, YYPK17]. Structured [AHS17, LH20]. structures [PSM^{+}06]. STT [AKW^{+}13, LPW18, NLW^{+}20, PFRR17, SFD17, SMT21, VDB^{+}16, ZCX^{+}17]. STT-MRAM [AKW^{+}13, SMT21]. STT-MRAM-based [NLW^{+}20]. STT-RAM [LPW18]. STTRAM [GG17, MGK18, WX15]. STTRAM-Based [GG17, WX15]. studies [CNPR14]. Study [DDR^{+}16a, KWC^{+}20, KST^{+}22, PPM^{+}13, PSY^{+}18, SCT^{+}22, YLF^{+}17, CB09, HCTK08]. styles [CJ14b]. Sub [GVRR17, ON15, RFDT15, WOW^{+}10]. Sub- [RFDT15]. Sub-10nm [GVRR17]. Sub-Block [ON15]. sub-mW [WOW^{+}10]. Subcrossbar [YL14]. Substrate [PAB^{+}17]. Success [NPS^{+}20]. Suite [XCS^{+}19]. Sum [KKK22]. Super [GVRR17]. Super-Threshold [GVRR17]. Supercapacitor [LPB^{+}15]. Supercapacitor-Based [LPB^{+}15]. Superconducting [GGHP^{+}24, PP21]. Supervised [CZQK15, LYL19]. Supervision [EFRB22]. Supply [EFRB22, LWM^{+}14, MLP^{+}20, MMJ09].

Support [CTT^{+}20, KMS^{+}20]. Surface [CMJ14, KDMT22, KTW08]. Surface-Wave [KDMT22]. Surgical [XYJ^{+}21]. Survey [BVM^{+}19, EDZ^{+}23, GBLD15, GNY^{+}22, KAB^{+}21, KK12, KRG^{+}23, Mit16, Mit17, QCF^{+}16, SFD17, STNP21, CNPR14]. sustainability [KMD12]. Sustainable [RSD^{+}23a, RSD23b, PG12]. SW [JRLR15]. SWIFTNoC [CDP17]. Swing [SSF^{+}15]. Switched [GB18]. Switching [BSL^{+}18, DRG21, MKG^{+}23]. SWMR [HAV^{+}22]. Symbol [GLL^{+}21a]. Symmetric [DDR^{+}16b]. Symposium [Shu09]. Synapse [LWY22]. Synapses [WKL16, WLJC21, JRC^{+}13]. Synaptic [BOAC^{+}20, KZL15, WKL16]. Synchronous [RFDT15]. Synthesis [AHSZ16, CEW^{+}13, CH14, CCH16, CW08, DRSR14, DWK^{+}16, DDR^{+}16b, GCJ17, HSZM17, HD14, LCSP14, MPM13, PT14a, PAC^{+}22, SMZ^{+}19, SSP14, SKB13, CCTP08, DJ08, LJ14, MMJ09, PT14b, SZSS10, SC08, XC08, ZXC10, VTKT22]. Synthetic [Dea14, FHFK14, HD14, MHW14, MSW14, OBLD14, VMV13]. System [ASB^{+}21, ARLB18, BY18, CJ15, DEW^{+}23, DLL^{+}19, HH11, AML^{+}23, JWJ^{+}17, KSA^{+}22, LBJ^{+}16, LWM^{+}14, MKSW17, PAP^{+}22, QCF^{+}16, RSD^{+}23a, SS15, SN11, SCZ^{+}12, TZR20, TSMCB17, WXW^{+}17, CB09, MP10, YW13, ZJS09c, ZJS09b, PCD^{+}11, YXW^{+}12]. System-Level [ARLB18, JWJ^{+}17, MP10]. System-on-chip [DLL^{+}19, PCD^{+}11, YXW^{+}12]. Systematic [PB21]. Systems [ACJ17, ARD24, BAT^{+}22, CMZR23, DMYT15, DRL^{+}19, EDZ^{+}23, FKM22, FYJ^{+}17, GBLD15, HIH18, IASK20, KZW^{+}15, KDMT22, MSCS19, MGMU22, MAC^{+}21, RMW^{+}17, RNN^{+}22, REL^{+}22, SAAR20, SPS^{+}24, TWLL19, TTS22, TSB15, TKBM12, TV17, WZL16, XNK18, ZSXY11, GMM12, LBGR08, LWX^{+}14, Moh12, MCT10, PG12, STA^{+}12, SLS^{+}14, WOW^{+}10, WCSA10]. systems-on-chip [GMM12, LWX^{+}14]. Systolic [LZXL22]. Systolic-Array [LZXL22].

T [MCT18]. T-count [MCT18]. Tag [RD22]. Tag-enforced [RD22]. Taming [AGR^{+}23]. Targeting [CKC^{+}18, HSZM17, MKG^{+}23]. tasks [STA^{+}12]. Taxonomy [SFD17]. TCAMs [dLBHC22]. Technique [CCW18, Che15, HRR19, HLH^{+}12, TFR23]. Techniques [Bis17, EDZ^{+}23, GBLD15, KPFM16, Mit16, Mit17, RBHG21, TJ13b, WYZ^{+}20, BCT^{+}13, KP10]. Technologies [DLTSNA20, FNO^{+}19, GBLD15, TSB15, BC08, Edi14, PR13, PUBV07, VO06, GN08]. Technology [BLKM23, BSY^{+}16, CTP14, FCR23, GMM12, GVRR17, KZW^{+}15, MGMU22, TGCJ16, TTS22, KSB^{+}08, MHM^{+}08, SX11, XDX14, ZC07]. Technology-Agnostic [CTP14]. Technology-Based [BSY^{+}16]. Technology-driven [GMM12]. Telementoring [XYJ^{+}21]. Temperature [HTMH18, PRG^{+}15, KCC^{+}14, LWH14]. Template [RDH14]. Templates [SM11]. Temporal [LFDS22, MDS21, MMAAK23, KWFH12]. tenant [CMZR23]. Terabit [KBT24]. Term [ZSPC19]. Ternary [GG17]. Test [AUDS22, CBK^{+}22, FMW^{+}22, MRG22, MSCS19, MKG^{+}23, TJ13b, ZCB^{+}22, IBO22, XS14]. Testability [DDR^{+}16b]. Testable [SDSS14, XDX14, LBGR08]. Testing [BMB18, AML^{+}23, KKC17, DJRM09, RT08, XS14, ZJT^{+}14]. TFET [DLTSNA20, KCWL^{+}16]. TFT [HCTK08, LWH14]. TFTs [LBGR08]. Their [KAB^{+}21]. theoretic [DWL10]. Theory [RBGC14]. Thermal [ARLB18, DJH^{+}19, KDMT22, MSCS19, TJ13b, ZY18, ZLGL21, KWFH12, XS14]. Thermal-aware [MSCS19]. thermal-driven [XS14].

Thermal-reliability-aware [DJH^{+}19]. Thermal-Sensitive [ZY18]. thing [RSD^{+}23a]. Things [MPR^{+}22]. Third [FMTP22]. Thread [LML^{+}19]. Threats [ASMK22, GSC17]. three [MLK^{+}08, WFCX09, XS14]. three-dimensional [MLK^{+}08, WFCX09, XS14]. Threshold [GVRR17, Mit16, PT14a, SCL17, MMJ09, PT12, PT14b, WDH^{+}09]. Through-Silicon-Via [TZS14]. Throughput [WL22b, CA11, RMBC12]. tier [CMK^{+}21]. Tile [HMS^{+}05, WWG^{+}19, ZLB^{+}22]. Tile-based [HMS^{+}05, WWG^{+}19]. Tiled [DPB11]. Time [DRL^{+}19, GSC17, JWJ^{+}17, KR18, KPFM16, LMM18, MDS21, MKSW17, NVW^{+}22, SPS^{+}24, SHAC19, TWLL19, XLL^{+}18, ZWL^{+}15, AMH^{+}24, LWX^{+}14, MNT14, SM19, SPR18, ZBF^{+}22]. Time-based [MDS21]. Time-Randomized [SHAC19]. Time-varying [NVW^{+}22]. timed [WCSA10]. Timing [ASK23, AMF^{+}15, CRSSBMR21, TMG^{+}21, LYL19]. Timing-aware [TMG^{+}21]. Timing-based [ASK23]. Timing-Optimized [CRSSBMR21]. TLC [AM18]. TODAES [BC08]. Tolerance [SLC^{+}17, WDW13, DWL10, LSH14, SCI^{+}09, Tah06, Tah09, XC08]. Tolerant [BKJ19, CLZ^{+}22, CVK15, DJ16, GYM^{+}17, HH11, LCK19, MGZ^{+}17, XYM18, ABR^{+}21, LWX^{+}14, PDL07, TWL09, YYC07, ZMT13]. Tolerating [MZZ23]. Tomography [XYJ^{+}21, XZL^{+}21]. Tool [VRBS16, dLBHC22, HZY^{+}12]. Topology [LGYC21]. Torque [XZR^{+}22, YLF^{+}17, AKW^{+}13, CSKM13, CWL^{+}13, MFA^{+}13]. Torus [YXW^{+}12]. Torus-Based [YXW^{+}12]. Tracking [ZPL^{+}20]. trade [CDG^{+}12]. trade-offs [CDG^{+}12]. Tradeoffs [HTMH18, SFD17]. Trading [ACH^{+}17]. Traditional [KMS^{+}20]. Trained [WL19, ZK18]. Training [HHD^{+}23, LDP^{+}20, LWY22, MS19]. Transfer [CZW^{+}19, AKW^{+}13, CSKM13, CWL^{+}13, MFA^{+}13, XZR^{+}22]. Transfer-Level [CZW^{+}19]. Transferable [HMP^{+}22]. Transfers [BMTP24]. Transform [DBG^{+}14]. Transformations [GOGCK11, BCT^{+}13]. transient [ZFT13]. Transiently [JRLR15]. Transistor [CEW^{+}13, CCH16, GLMG^{+}15, HC15, RMW^{+}17, WHL^{+}21, DLWW08, HZY^{+}12]. Transistors [HN12, KAB^{+}21, MGZ^{+}17, WDH^{+}09]. Transparent [KMS^{+}20]. Transporters [PPM^{+}13]. Tree [CH14, GCJ17, MMAAK23, YYC07]. Trends [LZBW20, TGCJ16, ZLWB20]. Triangle [KPM22]. Triangle-based [KPM22]. Trojan [ATW^{+}22, CKWK18, DRG21, GCC^{+}23, MKG^{+}23, OK22, RSD^{+}23a, WHB^{+}21]. Trojans [ZFT13, ZLGL21]. True [LTKP16]. Truly [KHC^{+}22]. Trust [BWL^{+}21, IGR^{+}16]. Trusted [FMTP22]. Trustworthy [JHPG22, SK16]. TSV [AGD^{+}20, KYEB15, MRG22]. TSVs [ARLB18]. Tumor [WL22a]. Tunable [GKT^{+}18, WLJC21]. Tunnel [VDB^{+}16]. tunneling [LM13]. Tutorial [BWL^{+}21]. Two [BMTP24, GUP11]. Two-Dimensional [GUP11]. Type [YJ18].

U [GGHP^{+}24, WL22a]. U-Net-based [WL22a]. U-SFQ [GGHP^{+}24]. ULS [MP10]. Ultra [CJ16, CJ14b, ERGK21, GGTG^{+}20, GJ17, KZW^{+}15, KT14, MGS^{+}12, STSG17, TSB15]. Ultra-Low [TSB15]. Ultra-Low-Leakage [CJ16, CJ14b, GJ17]. Ultra-Low-Power [KZW^{+}15, MGS^{+}12, STSG17, ERGK21, GGTG^{+}20]. Uncertainty [RPBA21]. Unclonable [LTM22, TMG^{+}21, UMB^{+}18]. Underlapped [GVRR17]. UNet [CLZ^{+}22]. Unified [YYPK17]. Uniform [SMT^{+}17]. Unipolar [LP17]. Unit [KHC^{+}22]. Units [BH17, Gla14]. Universal [CZQK15, CVK15, MP10]. Unsupervised [ASB^{+}21, ATW^{+}22, KSA^{+}22, LYL19, SPR18]. Updates [FKM22]. Use [KMS^{+}20]. Using [AVK16, ALY^{+}21, ATW^{+}22, Bis21, CHA20, CJ16, CKWK18, DSB16, DPB11, DLL^{+}19, EDCL^{+}22, GKT^{+}18, GOGCK11, GGTG^{+}20, GGHP^{+}24, GUP11, GSC17, GJ17, HZSA14, HLH^{+}12, JRJ22, KPM22, KR18, KSA^{+}22, KWWI17, LBJ^{+}16, LQYL19, LP17, LMM18, MDS21, MSB^{+}22, Mit17, MGST22, MSC^{+}21, NPS^{+}20, PFRR17, PT14a, PAF18, PB21, RBHG21, RDS24, SCL17, SBZT20, SPS^{+}24, TZR20, TFZ^{+}21, TNWD20, VSRR15, WM24, YLR^{+}23, YJ18, ZCDD19, BSL^{+}18, BPH^{+}11, CMJ14, DRSR14, HMS^{+}05, JRC^{+}13, KRP^{+}21, KT14, KK12, KCD15, LBGR08, LSH14, MMJ09, MHM^{+}08, PAP^{+}22, RBGC14, SZSS10, SPR18, SXL^{+}12, SC06, YYC07, ZFT13, ZJT^{+}14, KSB^{+}08]. Utility [LHW^{+}17]. Utilizing [LDZ22, OYLL24, WDH^{+}09].

Validation [SSN12]. Variability

[GPW^{+}15, NLK^{+}13, VDB^{+}16]. Variable [DKK^{+}15]. Variable-Latency [DKK^{+}15]. Variation [GLMG^{+}15, MZZ23, MGK18, XYM18, ZM22, MRH12, ZMT13]. Variation-Aware [GLMG^{+}15]. variation/defect [ZMT13]. variation/defect-tolerant [ZMT13]. Variations [CMK^{+}21, CMJ14, CJ14a, CJ15, DSB16, KAKSP14, SCL17, TGCJ16, YJ18, RYT^{+}07, XPD12]. Various [KBT24]. varying [NVW^{+}22]. Vdd [ZS08]. Vector [KHC^{+}22]. Vectors [MKG^{+}23]. Vehicular [LTM22]. Verification [EDZ^{+}23, MGST22, PB21, ZM22]. Versus [DDR^{+}16a, AGD^{+}20]. Vertical [BLKM23, KCWL^{+}16, LKC15]. Via [KBC^{+}23, TZS14, PHS^{+}23, WWJ09, YWF18, ZLGL21]. Vias [CBK^{+}22]. vibration [WCSA10]. Victims [ABM21].

Video [QGW20]. Viewpoint [GFZ13]. virtual [Sek07]. virtualized [MNT14]. Virtualizing [WWC23]. Vision [AML^{+}23]. VLIW [SLC^{+}17]. VLSI [AMF^{+}15, DKK^{+}15, MRH12, YP17]. Volatile [KHR^{+}15, RKM15, STSG17, WL22b, YYPK17, YXD^{+}17, ZPL^{+}20, RCYB22]. Voltage [Che15, EFRB22, IBO22, JOF^{+}15, KKKK18, MSB^{+}22, SCL17, SXL^{+}12]. Voltage-Controlled [MSB^{+}22]. voltages [MMJ09, WDH^{+}09]. vs [CJ14a, DWK^{+}16, KWC^{+}20]. Vulnerabilities [PAC^{+}22]. Vulnerability [BGX^{+}21, HON21, SKRX13].

Wafer [KKC17, KWWI17, MHM^{+}08]. Wafer-Bonding [KKC17]. wafer-to-wafer [MHM^{+}08]. Wall [KWC^{+}20, Mit17, SBZT20]. walled [SXL^{+}12]. Warp [MPZ21]. Wave [KDMT22, MKW^{+}14, EWKNW07]. Wavelength [LKL^{+}18]. waves [KK12]. Way [BS20, HSZM17, RPBA21]. Wear [LHHZ19]. Wear-aware [LHHZ19]. Web [AMVG12]. Weight [RSBA23]. Welded [GCJ17]. while [RYT^{+}07]. Wide [MPZ21]. Wide-Warp [MPZ21]. Width [BSS16]. Widths [GZZ^{+}21]. wire [Gla14, SXL^{+}12]. Wireless [ACM^{+}20, CMM^{+}18, DJ16, LPB^{+}15, LWM^{+}14, MKMS22, MGS^{+}12, MKW^{+}14, NVW^{+}22, RSD23b, RDS24, TKBM12, CDG^{+}12, GD12, WOW^{+}10, WVGP13]. wires [DK09]. wise [HYA^{+}20]. Within [KXY16]. without [ABR^{+}21]. Wordline [LYWW13]. Wordline/Bitline [LYWW13]. Workload [PRG^{+}15, SLS^{+}14]. Workload-Aware [PRG^{+}15]. World [MKW^{+}14]. Wormhole [SHAC19]. WoSAR [AD14]. Write [AM18, GLMG^{+}15, KXY16, LHW^{+}17, NLW^{+}20, WX15, WL22b]. Write-Aware [WX15]. Write-Optimized [WL22b].

Writes [VSRR15].

X [DDG^{+}22, KBC^{+}23]. X-Ray [KBC^{+}23].

Yield [PFRR17, SC06, FRB08].

Z [HZSA14]. Zallocator [WL22b]. Zero [KXY16, TNWD20, BPB^{+}12]. Zero-Energy [TNWD20].

References

Abate:2013:ILH

[AAFM13] Francesco Abate, Andrea Acquaviva, Elisa Ficarra, and Enrico Macii. Integration of literature with heterogeneous information for genes correlation scoring. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(4):28:1–28:??, November 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Asseman:2021:ADN

[AAO21] Alexis Asseman, Nicolas Antoine, and Ahmet S. Ozcan. Accelerating deep neuroevolution on distributed FP-GAs for reinforcement learning problems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):21:1–21:17, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                                     /dl.acm.org/doi/10.1145/
                                                     3425500.

Alam:2021:VCS

[ABM21] Manaar Alam, Sarani Bhattacharya, and Debdeep Mukhopadhyay. Victims can be saviors: a machine learning–based detection for micro-architectural side-channel attacks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):14:1– 14:31, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3439189.

Alam:2021:NNB

[ABR^{+}21] Manaar Alam, Arnab Bag, Debapriya Basu Roy, Dirmanto Jap, Jakub Breier, Shivam Bhasin, and Debdeep Mukhopadhyay. Neural network-based inherently fault-tolerant hardware cryptographic primitives without explicit redundancy checks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(1):3:1–3:30, January 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3409594.

Anagnostopoulou:2012:BAM

[ABS^{+}12] Vlasia Anagnostopoulou, Susmit Biswas, Heba Saadeldeen, Alan Savage, Ricardo Bianchini, Tao Yang, Diana Franklin, and Frederic T. Chong. Barely alive memory servers: Keeping data active in a low-power state. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(4):31:1– 31:??, October 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Alaghi:2017:TAE

[ACH^{+}17] Armin Alaghi, Wei-Ting J. Chan, John P. Hayes, Andrew B. Kahng, and Jiajia Li. Trading accuracy for energy in stochastic circuit design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3): 47:1–47:??, May 2017. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Abellan:2017:EPN

[ACJ17] José L. Abellán, Chao Chen, and Ajay Joshi. Electrophotonic NoC designs for kilocore systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):24:1–24:??, March 2017. CODEN ???? ISSN 1550-4832.

Ascia:2020:EDR

[ACM^{+}20] Giuseppe Ascia, Vincenzo Catania, Salvatore Monteleone, Maurizio Palesi, Davide Patti, John Jose, and Valerio Mario Salerno. Exploiting data resilience in wireless network-on-chip architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2):21:1–21:27, April 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3379448.

Avritzer:2014:ISI

[AD14] Alberto Avritzer and Tadashi Dohi. Introduction to special issue on WoSAR 2011. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):7:1–7:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Arka:2020:MCP

[AGD^{+}20] Aqeeb Iqbal Arka, Srinivasan Gopal, Janardhan Rao Doppa, Deukhyoun Heo, and Partha Pratim Pande. Making a case for partially connected 3D NoC: NFIC versus TSV. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(4):41:1–41:17, October 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3394919.

Ardesi:2023:TMF

[AGR^{+}23] Yuri Ardesi, Umberto Garlando, Fabrizio Riente, Giuliana Beretta, Gianluca Piccinini, and Mariagrazia Graziano. Taming molecular field-coupling for nanocomputing design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(1):1:1– 1:??, January 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3552520.

Ayub:2021:PPE

[AHHS21] Muhammad Kamran Ayub, Muhammad Abdullah Hanif, Osman Hasan, and Muhammad Shafique. PEAL: Probabilistic error analysis methodology for low-power approximate adders. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(1):1:1–1:37, January 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3405430.

A:2021:IQF

[AHPC21] Asha K. A., Li En Hsu, Abhishek Patyal, and Hung-Ming Chen. Improving the quality of FPGA RO-PUF by principal component analysis (PCA). ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):34:1–34:25, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3442444.

Anwar:2017:SPD

[AHS17] Sajid Anwar, Kyuyeon Hwang, and Wonyong Sung. Structured pruning of deep convolutional neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):32:1– 32:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Anandakumar:2022:DAF

[AHS22] N. Nalla Anandakumar, Mohammad S. Hashmi, and Somitra Kumar Sanadhya. Design and analysis of FPGAbased PUFs with enhanced performance for hardwareoriented security. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):72:1– 72:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3517813.

Arabzadeh:2016:QLS

[AHSZ16] Mona Arabzadeh, Mahboobeh Houshmand, Mehdi Sedighi, and Morteza Saheb Zamani. Quantum-logic synthesis of Hermitian gates. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):40:1– 40:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Adavally:2021:DAP

[AIK21] Shashank Adavally, Mahzabeen Islam, and Krishna Kavi. Dynamically adapting page migration policies based on applications’ memory access behaviors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):16:1–16:24, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3444750.

Apalkov:2013:STT

[AKW^{+}13] Dmytro Apalkov, Alexey Khvalkovskiy, Steven Watts, Vladimir Nikitin, Xueti Tang, Daniel Lottis, Kiseok Moon, Xiao Luo, Eugene Chen, Adrian Ong, Alexander Driskill-Smith, and Mohamad Krounbi. Spin-transfer torque magnetic random access memory (STT-MRAM). ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(2):13:1–13:??, May 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Alawad:2017:SCS

[AL17] Mohammed Alawad and Mingjie Lin. Sketching computation with stochastic processing engines. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):46:1–46:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Alasad:2021:RSH

[ALY^{+}21] Qutaiba Alasad, Jie Lin, Jiann-Shuin Yuan, Deliang Fan, and Amro Awad. Resilient and secure hardware devices using ASL. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):11:1– 11:26, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3429982.

Alsuwaiyan:2018:MMT

[AM18] Ali Alsuwaiyan and Kartik Mohanram. MFNW: an MLC/TLC flip-n-write architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):28:1–28:??, July 2018. CODEN ???? ISSN 1550-4832.

Araujo:2014:SAE

[AMA^{+}14] Jean Araujo, Rubens Matos, Vandi Alves, Paulo Maciel, F. Vieira de Souza, Rivalino Matias Jr., and Kishor S. Trivedi. Software aging in the Eucalyptus cloud computing infrastructure: Characterization and rejuvenation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):11:1– 11:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Azghadi:2015:PST

[AMF^{+}15] Mostafa Rahimi Azghadi, Saber Moradi, Daniel B. Fasnacht, Mehmet Sirin Ozdas, and Giacomo Indiveri. Programmable spike-timing-dependent plasticity learning circuits in neuromorphic VLSI architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(2):17:1– 17:??, August 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Ahmed:2024:DTR

[AMH^{+}24] Soyed Tuhin Ahmed, Mahta Mayahinia, Michael Hefenbrock, Christopher Münch, and Mehdi B. Tahoori. Design-time reference current generation for robust spintronic-based neuromorphic architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(1):2:1–2:??, January 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3625556.

Hasan:2023:EEV

[AML^{+}23] Md. Mahfuz Al Hasan, Mohammad Tahsin Mostafiz, Thomas An Le, Jake Julia, Nidish Vashistha, Shayan Taheri, and Navid Asadizanjani. EVHA: Explainable vision system for hardware testing and assurance — an overview. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(3):24:1–24:??, July 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3590772.

Abbasi:2012:DGD

[AMVG12] Zahra Abbasi, Tridib Mukherjee, Georgios Varsamopoulos, and Sandeep K. S. Gupta. DAHM: a green and dynamic Web application hosting manager across geographically distributed data centers. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(4):34:1– 34:??, October 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Anonymous:2018:GEI

[Ano18] Anonymous. Guest Editor introduction: Neuromorphic computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4):39:1–39:??, December 2018. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3283217.

Arasu:2014:RIL

[ANR^{+}14] Senthil Arasu, Mehrdad Nourani, Vijay Reddy, John M. Carulli Jr., Gautam Kapila, and Min Chen. Reliability improvement of logic and clock paths in powerefficient designs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):3:1–3:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Alam:2022:SMM

[ANT22] Mohsen Riahi Alam, M. Hassan Najafi, and Nima Taherinejad. Sorting in memristive memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):69:1–69:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3517181.

Amouzegar:2024:GCM

[ARD24] Morteza Amouzegar, Morteza Rezaalipour, and Masoud Dehyadegari. Genetic cache: a machine learning approach to designing DRAM cache controllers in HBM systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(3):11:1– 11:??, July 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3676966.

Alqahtani:2018:SLA

[ARLB18] Ayed Alqahtani, Zongqing Ren, Jaeho Lee, and Nader Bagherzadeh. System-level analysis of 3D ICs with thermal TSVs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(3):37:1–37:??, October 2018. CODEN ???? ISSN 1550-4832.

Adnan:2021:DRM

[ASB^{+}21] Md Musabbir Adnan, Sagarvarma Sayyaparaju, Samuel D. Brown, Mst Shamim Ara Shawkat, Catherine D. Schuman, and Garrett S. Rose. Design of a robust memristive spiking neuromorphic system with unsupervised learning in hardware. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):56:1–56:26, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3451210.

Ali:2023:CTB

[ASK23] Usman Ali, Sheikh Abdul Rasheed Sahni, and Omer Khan. Characterization of timing-based software sidechannel attacks and mitigations on network-on-chip hardware. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(3):21:1–21:??, July 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3585519.

Alam:2022:NLL

[ASMK22] Manaar Alam, Sayandeep Saha, Debdeep Mukhopadhyay, and Sandip Kundu. NN-Lock: a lightweight authorization to prevent IP threats of deep learning models. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):51:1– 51:19, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3505634.

Athreyas:2018:MCA

[ASP^{+}18] Nihar Athreyas, Wenhao Song, Blair Perot, Qiangfei Xia, Abbie Mathew, Jai Gupta, Dev Gupta, and J. Joshua Yang. Memristor-CMOS analog coprocessor for acceleration of highperformance computing applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(3):38:1–38:??, October 2018. CODEN ???? ISSN 1550-4832.

Ashok:2022:HTD

[ATW^{+}22] Maitreyi Ashok, Matthew J. Turner, Ronald L. Walsworth, Edlyn V. Levine, and Anantha P. Chandrakasan. Hardware Trojan detection using unsupervised deep learning on quantum diamond microscope magnetic field images. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):67:1– 67:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3531010.

Acharya:2022:TPO

[AUDS22] Nikita Acharya, Miroslav Urbanek, Wibe A. De Jong, and Samah Mohamed Saeed. Test points for online monitoring of quantum circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):14:1– 14:19, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3477928.

Ahsan:2016:DMQ

[AVK16] Muhammad Ahsan, Rodney Van Meter, and Jungsang Kim. Designing a millionqubit quantum computer using a resource performance simulator. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):39:1–39:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Bahar:2009:ISS

[Bah09] R. Iris Bahar. Introduction to special section: Best of NANOARCH 2008. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(2):6:1–6:??, July 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Bagherzadeh:2022:HSR

[BAT^{+}22] Javad Bagherzadeh, Aporva Amarnath, Jielun Tan, Subhankar Pal, and Ronald G. Dreslinski. A holistic solution for reliability of 3D parallel systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):23:1–23:27, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3488900.

Barenghi:2016:FBS

[BBB^{+}16] Alessandro Barenghi, Guido M. Bertoni, Luca Breveglieri, Gerardo Pelosi, Stefano Sanfilippo, and Ruggero Susella. A fault-based secret key retrieval method for ECDSA: Analysis and countermeasure. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1):8:1– 8:??, December 2016. CO-DEN ???? ISSN 1550-4832.

Barbareschi:2022:GAB

[BBB^{+}22] Mario Barbareschi, Salvatore Barone, Alberto Bosio, Jie Han, and Marcello Traiola. A genetic-algorithm-based approach to the design of DCT hardware accelerators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):50:1– 50:25, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3501772.

Bartley:2023:BOR

[BBR^{+}23] Bryan Bartley, Jacob Beal, Miles Rogers, Daniel Bryce, Robert P. Goldman, Benjamin Keller, Peter Lee, Vanessa Biggers, Joshua Nowak, and Mark Weston. Building an open representation for biological protocols. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(3):28:1– 28:??, July 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3604568.

Bahar:2008:IJA

[BC08] R. Iris Bahar and Krishnendu Chakrabarty. Introduction to joint ACM JETC/TODAES special issue on new, emerging, and specialized technologies. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(2): 5:1–5:??, April 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Bobba:2013:CTP

[BCT^{+}13] Shashikanth Bobba, Ashutosh Chakraborty, Olivier Thomas, Perrine Batude, and Giovanni de Micheli. Cell transformations and physical design techniques for 3D monolithic integrated circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(3):19:1–19:??, September 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Beausoleil:2011:LSI

[Bea11] Raymond G. Beausoleil. Large-scale integrated photonics for high-performance interconnects. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(2):6:1–6:??, June 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Biernacki:2021:SDS

[BGX^{+}21] Lauren Biernacki, Mark Gallagher, Zhixing Xu, Misiker Tadesse Aga, Austin Harris, Shijia Wei, Mohit Tiwari, Baris Kasikci, Sharad Malik, and Todd Austin. Software-driven security attacks: From vulnerability sources to durable hardware defenses. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):42:1– 42:38, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3456299.

Britt:2017:HPC

[BH17] Keith A. Britt and Travis S. Humble. High-performance computing with quantum processing units. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):39:1–39:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Blott:2019:QBN

[BHLD19] Michaela Blott, Lisa Halder, Miriam Leeser, and Linda Doyle. QuTiBench: Benchmarking neural networks on heterogeneous hardware. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):37:1– 37:??, December 2019. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3358700.

Biswas:2017:SAT

[Bis17] Arnab Kumar Biswas. Source authentication techniques for network-on-chip router configuration packets. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):28:1– 28:??, March 2017. CODEN ???? ISSN 1550-4832.

Biswas:2021:NCI

[Bis21] Arnab Kumar Biswas. Networkon-chip intellectual property protection using circular path-based fingerprinting. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(1):4:1– 4:22, January 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3410024.

Bhoj:2010:GDF

[BJ10] Ajay N. Bhoj and Niraj K. Jha. Gated-diode FinFET DRAMs: Device and circuit design-considerations. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(4):12:1– 12:??, December 2010. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Bhanu:2019:FTN

[BKJ19] P. Veda Bhanu, Pranav Venkatesh Kulkarni, and Soumya J. Fault-tolerant network-on-chip design with flexible spare core placement. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):5:1–5:??, February 2019. CODEN ???? ISSN 1550-4832.

Bhat:2023:SFG

[BLKM23] Sachin Bhat, Mingyu Li, Sourabh Kulkarni, and Csaba Andras Moritz. SkyBridge 2.0: a fine-grained vertical 3D-IC technology for future ICs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(4):33:1– 33:??, October 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3617501.

Bahreini:2015:MMS

[BM15] Tayebeh Bahreini and Naser Mohammadzadeh. An MINLP model for scheduling and placement of quantum circuits with a heuristic solution approach. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3):29:1–29:??, September 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Bhattacharjee:2018:RFT

[BMB18] Sukanta Bhattacharjee, Debasis Mitra, and Bhargab B. Bhattacharya. Robust in-field testing of digital microfluidic biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1):5:1–5:??, March 2018. CODEN ???? ISSN 1550-4832.

Bean:2024:PBD

[BMTP24] Benjamin Bean, Cyrus Minwalla, Eirini Eleni Tsiropoulou, and Jim Plusquellic. PUFbased digital money with propagation-of-provenance and offline transfers between two parties. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(3):9:1–9:??, July 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3663676.

Beckmann:2020:TSB

[BOAC^{+}20] Karsten Beckmann, Wilkie Olin-Ammentorp, Gangotree Chakma, Sherif Amer, Garrett S. Rose, Chris Hobbs and Joseph Van Nostrand, Martin Rodgers, and Nathaniel C. Cady. Towards synaptic behavior of nanoscale ReRAM devices for neuromorphic computing applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2):23:1– 23:18, April 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3381859.

Banerjee:2012:TNZ

[BPB^{+}12] Prithviraj Banerjee, Chandrakant Patel, Cullen Bash, Amip Shah, and Martin Arlitt. Towards a net-zero data center. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(4):27:1–27:??, October 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Biberman:2011:PNC

[BPH^{+}11] Aleksandr Biberman, Kyle Preston, Gilbert Hendry, Nicolás Sherwood-Droz, Johnnie Chan, Jacob S. Levy, Michal Lipson, and Keren Bergman. Photonic networkon-chip architectures using multilayer deposited silicon materials for highperformance chip multiprocessors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(2):7:1–7:??, June 2011. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Bashir:2019:BSO

[BPS19] Janibul Bashir, Eldhose Peter, and Smruti R. Sarangi. BigBus: a scalable optical interconnect. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):8:1–8:??, February 2019. CODEN ???? ISSN 1550-4832.

Biookaghazadeh:2021:TMF

[BRZ21] Saman Biookaghazadeh, Pravin Kumar Ravi, and Ming Zhao. Toward multi-FPGA acceleration of the neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):25:1– 25:23, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3432816.

Barke:2015:CLA

[BS15] Martin Barke and Ulf Schlichtmann. A cross-layer approach to measure the robustness of integrated circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3):24:1– 24:??, September 2015. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Bashir:2020:PSR

[BS20] Janibul Bashir and Smruti Ranjan Sarangi. Predict, share, and recycle your way to low-power nanophotonic networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):4:1–4:26, February 2020. CODEN ???? ISSN 1550-4832. URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3356585.

Bashir:2021:GPE

[BS21] Janibul Bashir and Smruti R. Sarangi. GPUOPT: Powerefficient photonic networkon-chip for a scalable GPU. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(1):8:1– 8:26, January 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3416850.

Bhattacharjee:2018:KSA

[BSL^{+}18] Debjyoti Bhattacharjee, Anne Siemon, Eike Linn, Stephan Menzel, and Anupam Chattopadhyay. Kogge–Stone adder realization using 1S1R resistive switching crossbar arrays. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):30:1–30:??, July 2018.

CODEN ???? ISSN 1550-4832.

Biswas:2016:IFW

[BSS16] Kalyan Biswas, Angsuman Sarkar, and Chandan Kumar Sarkar. Impact of fin width scaling on RF/ analog performance of junctionless accumulation-mode bulk FinFET. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):36:1–36:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Bi:2016:ETB

[BSY^{+}16] Yu Bi, Kaveh Shamsi, Jiann-Shiun Yuan, Pierre-Emmanuel Gaillardon, Giovanni De Micheli, Xunzhao Yin, X. Sharon Hu, Michael Niemier, and Yier Jin. Emerging technologybased design of primitives for hardware security. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1):3:1–3:??, December 2016. CODEN ???? ISSN 1550-4832.

Bouvier:2019:SNN

[BVM^{+}19] Maxence Bouvier, Alexandre Valentian, Thomas Mesquida, François Rummens, Marina Reyboz, Elisa Vianello, and Edith Beigne. Spiking neural networks hardware implementations and challenges: a survey. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(2):

22:1–22:??, June 2019. CO-DEN ???? ISSN 1550-

                                        4832. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3304103.

Botero:2021:HTA

[BWL^{+}21] Ulbert J. Botero, Ronald Wilson, Hangwei Lu, Mir Tanjidur Rahman, Mukhil A. Mallaiyan, Fatemeh Ganji, Navid Asadizanjani, Mark M. Tehranipoor, Damon L. Woodard, and Domenic Forte. Hardware trust and assurance through reverse engineering: a tutorial and outlook from image analysis and machine learning perspectives. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):62:1– 62:53, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3464959.

Bhunia:2012:ISI

[BY12] Swarup Bhunia and Darrin J. Young. Introduction to special issue on implantable electronics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(2):7:1–7:??, June 2012. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Bai:2018:DEE

[BY18] Kangjun Bai and Yang Yi. DFR: an energy-efficient analog delay feedback reservoir computing system for braininspired computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4):45:1– 45:??, December 2018. CO-DEN ???? ISSN 1550-4832.

Bontupalli:2018:EMB

[BYHT18] Venkataramesh Bontupalli, Chris Yakopcic, Raqibul Hasan, and Tarek M. Taha. Efficient memristor-based architecture for intrusion detection and high-speed packet classification. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4):41:1–41:??, December 2018. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3264819.

Cianchetti:2011:LLH

[CA11] Mark J. Cianchetti and David H. Albonesi. A lowlatency, high-throughput onchip optical router architecture for future chip multiprocessors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(2):9:1–9:??, June 2011. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chakraborty:2009:SAD

[CB09] Rajat Subhra Chakraborty and Swarup Bhunia. A study of asynchronous design methodology for robust CMOS-nano hybrid system design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(3):12:1–12:??, August 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chaudhuri:2022:BST

[CBK^{+}22] Arjun Chaudhuri, Sanmitra Banerjee, Jinwoo Kim, Heechun Park, Bon Woong Ku, Sukeshwar Kannan, Krishnendu Chakrabarty, and Sung Kyu Lim. Built-in self-test and fault localization for inter-layer vias in monolithic 3D ICs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):22:1– 22:37, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3464430.

Chen:2016:AMS

[CCH16] Yi-Hang Chen, Jian-Yu Chen, and Juinn-Dar Huang. Area minimization synthesis for reconfigurable single-electron transistor arrays with fabrication constraints. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):37:1– 37:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Carmona:2008:FMA

[CCTP08] Josep Carmona, Jordi Cortadella, Yousuke Takada, and Ferdinand Peper. Formal methods for the analysis and synthesis of nanometer-scale cellular arrays. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(2):8:1–8:??, April 2008. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chen:2018:CLA

[CCW18] Shuai Chen, Junlin Chen, and Lei Wang. A chiplevel anti-reverse engineering technique. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):29:1–29:??, July 2018. CODEN ???? ISSN 1550-4832.

Coussy:2015:FBN

[CCWCC15] Philippe Coussy, Cyrille Chavet, Hugues Nono Wouafo, and Laura Conde-Canencia. Fully binary neural network model and optimized hardware architectures for associative memories. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4):35:1–35:??, April 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

ChagasRibeiroDaRosa:2022:KQP

[CD22] Evandro Chagas Ribeiro Da Rosa and Rafael De Santiago. Ket quantum programming. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):12:1– 12:25, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3474224.

Chang:2012:PED

[CDG^{+}12] Kevin Chang, Sujay Deb, Amlan Ganguly, Xinmin Yu, Suman Prasad Sah, Partha Pratim Pande, Benjamin Belzer, and Deukhyoun Heo. Performance evaluation and design trade-offs for wireless network-on-chip architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):23:1–23:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chittamuru:2017:SRS

[CDP17] Sai Vineel Reddy Chittamuru, Srinivas Desai, and Sudeep Pasricha. SWIFTNoC: a reconfigurable silicon-photonic network with multicast-enabled channel sharing for multicore architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):58:1– 58:??, August 2017. CODEN ???? ISSN 1550-4832.

Chen:2013:SAR

[CEW^{+}13] Yung-Chih Chen, Soumya Eachempati, Chun-Yao Wang, Suman Datta, Yuan Xie, and Vijaykrishnan Narayanan. A synthesis algorithm for reconfigurable single-electron transistor arrays. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(1):5:1–5:??, February 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chen:2022:GES

[CFK22] Vanessa Chen, Mohammad AL Faruque, and Fadi Kurdahi. Guest editorial: Secure radiofrequency (RF)-Analog electronics and electromagnetics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):63:1– 63:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3564261.

Chen:2014:CTS

[CH14] Fu-Wei Chen and Tingting Hwang. Clock-tree synthesis with methodology of reuse in 3D-IC. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(3):22:1–22:??, April 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chakrabarty:2010:E

[Cha10] Krishnendu Chakrabarty. Editorial. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(1):1:1–1:??, March 2010. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Cambou:2020:CAS

[CHA20] Bertrand Cambou, David Hély, and Sareh Assiri. Cryptography with analog scheme using memristors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(4):40:1– 40:30, October 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3412439.

Cheng:2015:SSC

[Che15] Ching-Hwa Cheng. SCKVdd: a scalable clock-controlled self-stabilized voltage technique for low power CMOS digital circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(1):10:1–10:??, July 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Crocker:2009:DFQ

[CHN09] Michael Crocker, X. Sharon Hu, and Michael Niemier. Defects and faults in QCAbased PLAs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(2):8:1–8:??, July 2009. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chaudhuri:2014:VDS

[CJ14a] Sourindra M. Chaudhuri and Niraj K. Jha. 3D vs. 2D device simulation of FinFET logic gates under PVT variations. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(3):26:1– 26:??, April 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chen:2014:ULL

[CJ14b] Xianmin Chen and Niraj K.

Jha. Ultra-low-leakage chip multiprocessor design with hybrid FinFET logic styles. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(1):3:1– 3:??, September 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chen:2015:GPF

[CJ15] Xianmin Chen and Niraj K.

Jha. gem5-PVT: a framework for FinFET system simulation under PVT variations. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3):28:1– 28:??, September 2015. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chaudhuri:2016:ULL

[CJ16] Sourindra M. Chaudhuri and Niraj K. Jha. Ultra-low-leakage and high-performance logic circuit design using multiparameter asymmetric FinFETs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):43:1–43:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Canto:2023:RCK

[CKA23] Alvaro Cintas Canto, Mehran Mozaffari Kermani, and Reza Azarderakhsh. Reliable constructions for the key generator of code-based postquantum cryptosystems on FPGA. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(1):5:1–5:??, January 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3544921.

Crites:2020:DPM

[CKB20] Brian Crites, Karen Kong, and Philip Brisk. Directed placement for mVLSI devices. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2):14:1– 14:26, April 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3369585.

Chang:2018:PPA

[CKC^{+}18] Kyungwook Chang, Deepak Kadetotad, Yu Cao, Jae-Sun Seo, and Sung Kyu Lim. Power, performance, and area benefit of monolithic 3D ICs for on-chip deep neural networks targeting speech recognition. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4):42:1–42:??, December 2018. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3273956.

Cui:2018:HTD

[CKWK18] Xiaotong Cui, Elnaz Koopahi, Kaijie Wu, and Ramesh Karri. Hardware Trojan detection using the order of path delay. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(3):33:1– 33:??, October 2018. CODEN ???? ISSN 1550-4832.

Cao:2017:GEI

[CLKG17] Yu Cao, Xin Li, Taemin Kim, and Suyog Gupta. Guest editors’ introduction: Hardware and algorithms for onchip learning. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):30:1–30:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Cao:2018:GEI

[CLSD18] Yu Cao, Xin Li, Jae-Sun Seo, and Ganesh Dasika. Guest Editors’ introduction: Frontiers of hardware and algorithms for on-chip learning. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2): 14:1–14:??, July 2018. CO-DEN ???? ISSN 1550-4832.

Chen:2022:AUA

[CLZ^{+}22] Yufei Chen, Tingtao Li, Qinming Zhang, Wei Mao, Nan Guan, Mei Tian, Hao Yu, and Cheng Zhuo. ANT-UNet: Accurate and noise-tolerant segmentation for pathology image processing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):27:1–27:17, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3451213.

Chaudhuri:2014:ALD

[CMJ14] Sourindra M. Chaudhuri, Prateek Mishra, and Niraj K. Jha. Accurate leakage/delay estimation for FinFET standard cells under PVT variations using the response surface methodology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):19:1– 19:??, November 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chatterjee:2021:PMM

[CMK^{+}21] Anwesha Chatterjee, Shouvik Musavvir, Ryan Gary Kim, Janardhan Rao Doppa, and Partha Pratim Pande.

Power management of monolithic 3D manycore chips with inter-tier process variations. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):13:1– 13:19, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3430765.

Catania:2018:IEE

[CMM^{+}18] Vincenzo Catania, Andrea Mineo, Salvatore Monteleone, Maurizio Palesi, and Davide Patti. Improving energy efficiency in wireless networkon-chip architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1):9:1–9:??, March 2018. CODEN ???? ISSN 1550-4832.

Chen:2023:AED

[CMZR23] Xiangru Chen, Maneesh Merugu, Jiaqi Zhang, and Sandip Ray. AroMa: Evaluating deep learning systems for stealthy integrity attacks on multi-tenant accelerators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(2):13:1– 13:??, April 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3579033.

Crocker:2012:RPA

[CNH12] Michael Crocker, Michael Niemier, and X. Sharon Hu.

A reconfigurable PLA architecture for nanomagnet logic. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(1):1:1– 1:??, February 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Crocker:2008:MQD

[CNHL08] Michael Crocker, Michael Niemier, X. Sharon Hu, and Marya Lieberman. Molecular QCA design with chemically reasonable constraints. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(2):9:1–9:??, April 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Cotroneo:2014:SSA

[CNPR14] Domenico Cotroneo, Roberto Natella, Roberto Pietrantuono, and Stefano Russo. A survey of software aging and rejuvenation studies. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):8:1–8:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chen:2021:ISIa

[CQL21a] Yiran Chen, Qinru Qiu, and Yingyan Lin. Introduction of special issue on hardware and algorithms for efficient machine learning-part 1. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):18:1– 18:2, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3449045.

Chen:2021:ISIb

[CQL21b] Yiran Chen, Qinru Qiu, and Yingyan Lin. Introduction to the special issue on hardware and algorithms for efficient machine learning — part 2. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):45:1–45:2, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3464917.

Chen:2021:QDN

[CQZ^{+}21] Wentao Chen, Hailong Qiu, Jian Zhuang, Chutong Zhang, Yu Hu, Qing Lu, Tianchen Wang, Yiyu Shi, Meiping Huang, and Xiaowe Xu. Quantization of deep neural networks for accurate edge computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):54:1–54:11, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3451211.

Chabi:2014:RLA

[CQZK14] Djaafar Chabi, Damien Querlioz, Weisheng Zhao, and Jacques-Olivier Klein. Robust learning approach for neuroinspired nanoscale crossbar architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):5:1–5:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Choudhury:2022:HPE

[CRKP22] Dwaipayan Choudhury, Aravind Sukumaran Rajam, Ananth Kalyanaraman, and Partha Pratim Pande. Highperformance and energyefficient 3D manycore GPU architecture for accelerating graph analytics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):18:1– 18:19, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3482880.

Camacho-Ruiz:2021:TOH

[CRSSBMR21] Eros Camacho-Ruiz, Santiago Sánchez-Solano, Piedad Brox, and Macarena C. Martínez-Rodríguez. Timing-optimized hardware implementation to accelerate polynomial multiplication in the NTRU algorithm. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):35:1–35:16, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3445979.

Chakrabarty:2007:ESI

[CS07] Krishnendu Chakrabarty and Sachin Sapatnekar. Editorial to special issue DAC 2006. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(3):11:1– 11:??, November 2007. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chatterjee:2013:EAS

[CSKM13] Subho Chatterjee, Sayeef Salahuddin, Satish Kumar, and Saibal Mukhopadhyay. Electrothermal analysis of spin-transfer-torque random access memory arrays. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(2):15:1– 15:??, May 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

ChappetDeVangel:2015:RSD

[CThG15] Benoît Chappet De Vangel, Cesar Torres-huitzil, and Bernard Girau. Randomly spiking dynamic neural fields. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4):37:1– 37:??, April 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chung:2014:DET

[CTP14] Haera Chung, Christof Teuscher, and Partha Pande. Design and evaluation of technologyagnostic heterogeneous networkson-chip. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(3):20:1–20:??, April 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Cui:2020:LSN

[CTT^{+}20] Weilong Cui, Georgios Tzimpragos, Yu Tao, Joseph McMahan, Deeksha Dangwal, Nestan Tsiskaridze, George Michelogiannakis, Dilip P. Vasudevan, and Timothy Sherwood. Language support for navigating architecture design in closed form. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):9:1–9:28, February 2020. CODEN ???? ISSN 1550-4832.

URL https://dl.acm.org/
doi/abs/10.1145/3360047.

Choi:2011:EQI

[CV11] Byung-Soo Choi and Rodney Van Meter. On the effect of quantum interaction distance on quantum addition circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(3):11:1–11:17, August 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Choi:2012:DQA

[CV12] Byung-Soo Choi and Rodney Van Meter. A Θ(\sqrt{n})- depth quantum adder on the 2D NTC quantum computer architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):24:1–24:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chien:2015:FTO

[CVK15] Chia-Hung Chien, Rodney Van Meter, and Sy-Yen Kuo. Fault-tolerant operations for universal blind quantum computation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(1):9:1–9:??, July 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chuang:2008:SRS

[CW08] Min-Lun Chuang and Chun-Yao Wang. Synthesis of reversible sequential elements. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(4):4:1– 4:??, January 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chen:2013:CCB

[CWL^{+}13] Yiran Chen, Weng-Fai Wong, Hai Li, Cheng-Kok Koh, Yaojun Zhang, and Wujie Wen. On-chip caches built on multilevel spin-transfer torque RAM cells and its optimizations. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(2):16:1–16:??, May 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chen:2014:CRP

[CWT14] Jifeng Chen, Shuo Wang, and Mohammad Tehranipoor. Critical-reliability path identification and delay analysis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(2): 12:1–12:??, February 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chakrabarty:2005:DAM

[CZ05] Krishnendu Chakrabarty and Jun Zeng. Design automation for microfluidics-based biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 1(3):186–223, October 2005. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Chabi:2015:CUS

[CZQK15] Djaafar Chabi, Weisheng Zhao, Damien Querlioz, and Jacques-Olivier Klein. Onchip universal supervised learning methods for neuroinspired block of memristive nanodevices. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4):34:1–34:??, April 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Cui:2019:SMB

[CZW^{+}19] Xiaotong Cui, Jeff (Jun)

Zhang, Kaijie Wu, Siddharth Garg, and Ramesh Karri. Split manufacturing-based register transfer-level obfuscation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1): 11:1–11:??, February 2019. CODEN ???? ISSN 1550-4832.

DeVos:2014:DGF

[DBG^{+}14] Alexis De Vos, Stéphane Burignat, Robert Glück, Torben Ægidius Mogensen, Holger Bock Axelsen, Michael Kirkedal Thomsen, Eva Rotenberg, and Tetsuo Yokoyama. Designing garbage-free reversible implementations of the integer cosine transform. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):11:1– 11:??, November 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Dey:2021:SCB

[DBS^{+}21] Sumon Dey, Lee Baker, Joshua Schabel, Weifu Li, and Paul D. Franzon. A scalable cluster-based hierarchical hardware accelerator for a cortically inspired algorithm. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):52:1– 52:29, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3447777.

Dang:2021:BCP

[DCP^{+}21] Dharanidhar Dang, Sai Vineel Reddy Chittamuru, Sudeep Pasricha, Rabi Mahapatra, and Debashis Sahoo. BPLight-CNN: a photonicsbased backpropagation accelerator for deep learning. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):49:1– 49:26, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3446212.

Dubey:2022:GML

[DCSA22] Anuj Dubey, Rosario Cammarota, Vikram Suresh, and Aydin Aysu. Guarding machine learning hardware against physical side-channel attacks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):56:1–56:31, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3465377.

DeVos:2014:MCC

[DD14] Alexis De Vos and Stijn De Baerdemacker. Matrix calculus for classical and quantum circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):9:1–9:??, November 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Danial:2022:EXD

[DDG^{+}22] Josef Danial, Debayan Das, Anupam Golder, Santosh Ghosh, Arijit Raychowdhury, and Shreyas Sen. EM-X-DL: Efficient cross-device deep learning side-channel attack with noisy EM signatures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):4:1– 4:17, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3465380.

Davids:2006:MFD

[DDM^{+}06] Daniel Davids, Siddhartha Datta, Arindam Mukherjee, Bharat Joshi, and Arun Ravindran. Multiple fault diagnosis in digital microfluidic biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(4):262–276, October 2006. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Das:2020:GRM

[DDP20] Subrata Das, Debesh Kumar Das, and Soumya Pandit. A global routing method for graphene nanoribbons based circuits and interconnects. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3): 31:1–31:28, July 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3384214.

Castro:2016:FVB

[DDR^{+}16a] Stephan De Castro, Jean-Max Dutertre, Bruno Rouzeyre, Giorgio Di Natale, and Marie-Lise Flottes. Frontside versus backside laser injection: a comparative study. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1):7:1–7:??, December 2016. CODEN ???? ISSN 1550-4832.

Deb:2016:RSS

[DDR^{+}16b] Arighna Deb, Debesh K. Das, Hafizur Rahaman, Robert Wille, Rolf Drechsler, and Bhargab B. Bhattacharya. Reversible synthesis of symmetric functions with a simple regular structure and easy testability. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):34:1–34:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Deans:2014:PNS

[Dea14] Tara L. Deans. Parallel networks: Synthetic biology and artificial intelligence. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3):21:1– 21:??, December 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Dehon:2005:NBP

[Deh05] André Dehon. Nanowirebased programmable architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 1(2):109–162, 2005. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Demirkiran:2023:EPS

[DEW^{+}23] Cansu Demirkiran, Furkan Eris, Gongyu Wang, Jonathan Elmhurst, Nick Moore, Nicholas C. Harris, Ayon Basumallik, Vijay Janapa Reddi, Ajay Joshi, and Darius Bunandar. An electro-photonic system for accelerating deep neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(4):30:1–30:??, October 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3606949.

Densmore:2023:ISI

[DHK^{+}23] Douglas Densmore, Nathan J. Hillson, Eric Klavins, Chris Myers, Jean Peccoud, and Giovanni Stracquadanio. Introduction to the special issue on BioFoundries and cloud laboratories. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(3):26:1–26:??, July 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3609485.

Donald:2008:RLS

[DJ08] James Donald and Niraj K. Jha. Reversible logic synthesis with Fredkin and Peres gates. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(1):2:1– 2:??, March 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Dehghani:2016:NAO

[DJ16] Abbas Dehghani and Kamal Jamshidi. A novel approach to optimize fault-tolerant hybrid wireless network-on-chip architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):45:1–45:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Dinakarrao:2019:ATR

[DJH^{+}19] Sai Manoj Pudukotai Dinakarrao, Arun Joseph, Anand Haridass, Muhammad Shafique, Jörg Henkel, and Houman Homayoun. Application and thermal-reliability-aware reinforcement learning based multi-core power management. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):33:1–33:??, December 2019. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3323055.

Datta:2009:EPT

[DJRM09] Siddhartha Datta, Bharat Joshi, Arun Ravindran, and Arindam Mukherjee. Efficient parallel testing and diagnosis of digital microfluidic biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(2):10:1–10:??, July 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Dysart:2009:OWR

[DK09] Timothy J. Dysart and Peter M. Kogge. Organizing wires for reliability in magnetic QCA. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(4):19:1–19:??, November 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Das:2021:CNM

[DK21] Palash Das and Hemangee K.

Kapoor. CLU: a near-memory accelerator exploiting the parallelism in convolutional neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):22:1–22:25, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3427472.

De:2015:ASC

[DKK^{+}15] Vivek K. De, Andrew B. Kahng, Tanay Karnik, Bao Liu, Milad Maleki, and Lu Wang. Application-specific cross-layer optimization based on predictive variable-latency VLSI design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3):21:1–21:??, September 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

deLima:2022:SAD

[dLBHC22] João Paulo Cardoso de Lima, Marcelo Brandalero, Michael Hübner, and Luigi Carro. STAP: an architecture and design tool for automata processing on memristor TCAMs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):39:1– 39:22, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450769.

Du:2019:SSA

[DLL^{+}19] Gaoming Du, Guanyu Liu, Zhenmin Li, Yifan Cao, Duoli Zhang, Yiming Ouyang, Minglun Gao, and Zhonghai Lu. SSS: Self-aware system-on-chip using a staticdynamic hybrid method. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(3): 28:1–28:??, June 2019. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3313869.

Delgado-Lozano:2020:PDR

[DLTSNA20] I. M. Delgado-Lozano, E. Tena-Sánchez, J. Núñez, and A. J. Acosta. Projection of dualrail DPA countermeasures in future FinFET and emerging TFET technologies. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3):30:1– 30:16, July 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3381857.

Deng:2008:CNT

[DLWW08] Jie Deng, Albert Lin, Gordon C. Wan, and H.-S. Philip Wong. Carbon nanotube transistor compact model for circuit design and performance optimization. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(2):7:1–7:??, April 2008. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Datta:2006:ADF

[DMR06] Kushal Datta, Arindam Mukherjee, and Arun Ravindran. Automated design flow for diode-based nanofabrics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(3):219– 241, July 2006. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Daneshtalab:2015:SIE

[DMYT15] Masoud Daneshtalab, Farhad Mehdipour, Zhiyi Yu, and Hannu Tenhunen. Special issue on emerging many-core systems for exascale computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4): 39:1–39:??, April 2015. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Dingler:2011:PEI

[DNHL11] Aaron Dingler, Michael T. Niemier, Xiaobo Sharon Hu, and Evan Lent. Performance and energy impact of locally controlled NML circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(1):2:1– 2:??, January 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Devadoss:2011:PQT

[DPB11] Rajeswari Devadoss, Kolin Paul, and M. Balakrishnan. p-QCA: a tiled programmable fabric architecture using molecular quantum-dot cellular automata. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(3):13:1–13:??, August 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Das:2011:ISI

[DR11] Shamik Das and Garrett S.

Rose. Introduction to special issue: Highlights of NANOARCH’09. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(1):1:1–1:??, January 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Dhar:2021:HTH

[DRG21] Tapobrata Dhar, Surajit Kumar Roy, and Chandan Giri. Hardware Trojan horse detection through improved switching of dormant nets. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):33:1– 33:22, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3439951.

Dridi:2019:DMA

[DRL^{+}19] Mourad Dridi, Stéphane Rubini, Mounir Lallali, Martha Johanna Sepúlveda Flórez, Frank Singhoff, and Jean-Philippe Diguet. Design and multi-abstraction-level evaluation of a NoC router for mixed-criticality realtime systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):2:1–2:??, February 2019. CODEN ???? ISSN 1550-4832.

Datta:2014:IRC

[DRSR14] Kamalika Datta, Gaurav Rathi, Indranil Sengupta, and Hafizur Rahaman. An improved reversible circuit synthesis approach using clustering of ESOP cubes. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):15:1– 15:??, November 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Das:2016:MPU

[DSB16] Jayita Das, Kevin Scott, and Sanjukta Bhanja. MRAM PUF: Using geometric and resistive variations in MRAM cells. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1):2:1–2:??, December 2016. CODEN ???? ISSN 1550-4832.

Deb:2016:GVS

[DWK^{+}16] Arighna Deb, Robert Wille, Oliver Keszöcze, Stefan Hillmich, and Rolf Drechsler. Gates vs. splitters: Contradictory optimization objectives in the synthesis of optical circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1):11:1– 11:??, December 2016. CO-DEN ???? ISSN 1550-4832.

Dai:2010:ITA

[DWL10] Jianwei Dai, Lei Wang, and Fabrizio Lombardi. An information-theoretic analysis of quantum-dot cellular automata for defect tolerance. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(3):9:1– 9:??, August 2010. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

El-Derhalli:2022:TAO

[EDCL^{+}22] Hassnaa El-Derhalli, Léa Constans, Sébastien Le Beux, Alfredo De Rossi, Fabrice Raineri, and Sofiène Tahar. Towards all-optical stochastic computing using photonic crystal nanocavities. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):16:1– 16:25, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3484871.

Editors:2014:ISI

[Edi14] Editors. Introduction to special issue on reliability and device degradation in emerging technologies. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):1:1–1:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Erata:2023:SAT

[EDZ^{+}23] Ferhat Erata, Shuwen Deng, Faisal Zaghloul, Wenjie Xiong, Onur Demir, and Jakub Szefer. Survey of approaches and techniques for security verification of computer systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(1):6:1– 6:??, January 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3564785.

Elkanishy:2022:LOH

[EFRB22] Abdelrahman Elkanishy, Paul M. Furth, Derrick T. Rivera, and Abdel-Hameed A. Badawy. Low-overhead hardware supervision for securing an IoT Bluetooth-enabled device: Monitoring radio frequency and supply voltage. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):6:1– 6:28, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3468064.

Eliahu:2021:MME

[ERGK21] Adi Eliahu, Ronny Ronen, Pierre-Emmanuel Gaillardon, and Shahar Kvatinsky. multi-PULPly: a multiplication engine for accelerating neural networks on ultra-low-power architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):24:1–24:27, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3432815.

Eshaghian-Wilner:2007:SWN

[EWKNW07] Mary M. Eshaghian-Wilner, Alex Khitun, Shiva Navab, and Kang L. Wang. The spin-wave nanoscale reconfigurable mesh and the labeling problem. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(2):5:1–5:??, July 2007. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Fusella:2018:RPC

[FC18] Edoardo Fusella and Alessandro Cilardo. Reducing power consumption of lasers in photonic NoCs through application-specific mapping. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):23:1– 23:??, July 2018. CODEN ???? ISSN 1550-4832.

Forero:2023:BOD

[FCR23] Freddy Forero, Victor Champac, and Michel Renovell. Bopen defect: a novel defect model in FinFET technology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(1):3:1– 3:??, January 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3564244.

Frache:2014:NAM

[FGZ14] Stefano Frache, Mariagrazia Graziano, and Maurizio Zamboni. Nanoarray architectures multilevel simulation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):6:1–6:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Fellermann:2014:FMD

[FHFK14] Harold Fellermann, Maik Hadorn, Rudolf M. Füchslin, and Natalio Krasnogor. Formalizing modularization and data hiding in synthetic biology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3): 24:1–24:??, December 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Falas:2022:MEE

[FKM22] Solon Falas, Charalambos Konstantinou, and Maria K. Michael. A modular endto-end framework for secure firmware updates on embedded systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):3:1–3:19, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3460234.

Forte:2021:ISI

[FMP^{+}21] Domenic Forte, Debdeep Mukhopadhyay, Ilia Polian, Yunsi Fei, and Rosario Cammarota. Introduction to the special issue on emerging challenges and solutions in hardware security. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):29:1–29:4, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3464326.

Fragkos:2022:EPP

[FMTP22] Georgios Fragkos, Cyrus Minwalla, Eirini Eleni Tsiropoulou, and Jim Plusquellic. Enhancing privacy in PUF-Cash through multiple trusted third parties and reinforcement learning. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):7:1–7:26, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3441139.

Fieback:2022:DFM

[FMW^{+}22] Moritz Fieback, Guilherme Cardoso Medeiros, Lizhou Wu, Hassen Aziza, Rajendra Bishnoi, Mottaqiallah Taouil, and Said Hamdioui. Defects, fault modeling, and test development framework for RRAMs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):52:1– 52:26, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3510851.

Fusella:2019:GEI

[FNO^{+}19] Edoardo Fusella, Mahdi Nikdast, Ian O’Connor, José Flich, and Sudeep Pasricha. Guest Editors’ introduction: Emerging networkson-chip designs, technologies, and applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):1:1–1:??, February 2019. CODEN ???? ISSN 1550-4832.

Ferri:2008:PYM

[FRB08] Cesare Ferri, Sherief Reda, and R. Iris Bahar. Parametric yield management for 3D ICs: Models and strategies for improvement. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(4):19:1–19:??, October 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Froehlich:2022:PCG

[FSD22] Saman Froehlich, Saeideh Shirinzadeh, and Rolf Drechsler. Parallel computing of graph-based functions in ReRAM. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):41:1–41:24, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3453163.

Farahmandi:2023:ISI

[FSDT23] Farimah Farahmandi, Ankur Srivastava, Giorgio Di Natale, and Mark Tehranipoor. Introduction to the special issue on CAD for security: Presilicon security sign-off solutions through design cycle. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(1):4:1– 4:??, January 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3584317.

Fang:2017:SPM

[FYJ^{+}17] Yan Fang, Victor V. Yashin, Brandon B. Jennings, Donald M. Chiarulli, and Steven P.

Levitan. A simplified phase model for simulation of oscillator-based computing systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):14:1–14:??, March 2017. CODEN ???? ISSN 1550-4832.

Grani:2018:SPS

[GB18] Paolo Grani and Sandro Bartolini. Scalable path-setup scheme for all-optical dynamic circuit switched NoCs in cache coherent CMPs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1):12:1– 12:??, March 2018. CODEN ???? ISSN 1550-4832.

Gaillardon:2015:SLP

[GBLD15]  Pierre-Emmanuel  Gaillar-

don, Edith Beigne, Suzanne Lesecq, and Giovanni De Micheli. A survey on lowpower techniques with emerging technologies: From devices to systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(2):12:1– 12:??, August 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Grissom:2014:IAC

[GCB14] Daniel Grissom, Christopher Curtis, and Philip Brisk. Interpreting assays with control flow on digital microfluidic biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(3):24:1–24:??, April 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Gaikwad:2023:HIA

[GCC^{+}23] Pravin Gaikwad, Jonathan Cruz, Prabuddha Chakraborty, Swarup Bhunia, and Tamzidul Hoque. Hardware IP assurance against Trojan attacks with machine learning and post-processing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(3):25:1– 25:??, July 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592795.

Ghosh:2017:AQC

[GCJ17] Mrityunjay Ghosh, Amlan Chakrabarti, and Niraj K. Jha. Automated quantum circuit synthesis and cost estimation for the binary welded tree oracle. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):51:1– 51:??, August 2017. CODEN ???? ISSN 1550-4832.

Gaillardon:2011:MNB

[GCO^{+}11] P.-E. Gaillardon, F. Clermidy, I. O’Connor, J. Liu, M. Amadou, and G. Nicolescu. Matrix nanodevicebased logic architectures and associated functional mapping method. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(1):3:1–3:??, January 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Guo:2020:PNO

[GCTF20] Zimu Guo, Sreeja Chowdhury, Mark M. Tehranipoor, and Domenic Forte. Permutation network de-obfuscation: a delay-based attack and countermeasure investigation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2):17:1– 17:25, April 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3371407.

Ghidini:2012:EEM

[GD12] Giacomo Ghidini and Sajal K. Das. Energy-efficient Markov chain-based duty cycling schemes for greener wireless sensor networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(4):29:1–29:??, October 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Graziano:2013:HVB

[GFZ13] Mariagrazia Graziano, Stefano Frache, and Maurizio Zamboni. A hardware viewpoint on biosequence analysis: What’s next? ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(4):29:1–29:??, November 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Govindaraj:2017:DAS

[GG17] Rekha Govindaraj and Swaroop Ghosh. Design and analysis of STTRAM-based ternary content addressable memory cell. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):52:1–52:??, August 2017. CODEN ???? ISSN 1550-4832.

Gonzalez-Guerrero:2024:TPS

[GGHP^{+}24] Patricia Gonzalez-Guerrero, Kylie Huch, Nirmalendu Patra, Thom Popovici, and George Michelogiannakis. Toward practical superconducting accelerators for machine learning using U-SFQ. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(2):7:1–7:??, April 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3653073.

Gonzalez-Guerrero:2020:TNM

[GGTG^{+}20] Patricia Gonzalez-Guerrero, Tommy Tracy II, Xinfei Guo, Rahul Sreekumar, Marzieh Lenjani, Kevin Skadron, and Mircea R. Stan. Towards on-node machine learning for ultra-low-power sensors using asynchronous ΣΔ streams. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(4):44:1– 44:20, October 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3404975.

Grimmer:2019:ASD

[GHHW19] Andreas Grimmer, Medina Hamidović, Werner Haselmayr, and Robert Wille. Advanced simulation of droplet microfluidics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(3):26:1–26:??, June 2019. CODEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3313867.

Gupta:2022:CCS

[GIS^{+}22] Saransh Gupta, Mohsen Imani, Joonseop Sim, Andrew Huang, Fan Wu, Jaeyoung Kang, Yeseong Kim, and Tajana Simunić Rosing. COSMO: Computing with stochastic numbers in memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2): 37:1–37:25, April 2022. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3484731.

Guler:2017:ULL

[GJ17] Abdullah Guler and Niraj K.

Jha. Ultra-low-leakage, robust FinFET SRAM design using multiparameter asymmetric FinFETs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):26:1–26:??, March 2017. CODEN ???? ISSN 1550-4832.

Gala:2018:ATN

[GKT^{+}18] Neel Gala, Sarada Krithivasan, Wei-Yu Tsai, Xueqing Li, Vijaykrishnan Narayanan, and V. Kamakoti. An accuracy tunable non-Boolean coprocessor using coupled nanooscillators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1):1:1–1:??, March 2018. CODEN ???? ISSN 1550-4832.

Gladshtein:2014:DBP

[Gla14] Michael Gladshtein. Delaybased processing-in-wire for design of QCA serial decimal arithmetic units. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(2):13:1– 13:??, February 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Gan:2021:CED

[GLL^{+}21a] Victor M. Gan, Yibin Liang, Lianjun Li, Lingjia Liu, and Yang Yi. A cost-efficient digital ESN architecture on FPGA for OFDM symbol detection. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):47:1–47:15, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3440017.

Gong:2021:SES

[GLL^{+}21b] Shijun Gong, Jiajun Li, Wenyan Lu, Guihai Yan, and Xiaowei Li. ShuntFlow-Plus: an efficient and scalable dataflow accelerator architecture for stream applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):59:1– 59:24, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3453164.

Ghofrani:2015:LPV

[GLMG^{+}15] Amirali Ghofrani, Miguel-Angel Lastras-Montaño, Siddharth Gaba, Melika Payvand, Wei Lu, Luke Theogarajan, and Kwang-Ting Cheng. A low-power variation-aware adaptive write scheme for access-transistor-free memristive memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(1):3:1–3:??, July 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Ganguly:2023:GEI

[GMGA23] Amlan Ganguly, Salvatore Monteleone, Diana Goehringer, and Cristinel Ababei. Guest Editors introduction: Special issue on network-on-chip architectures of the future (NoCArc). ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(3):19:1–19:??, July 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3609500.

Garg:2012:TDL

[GMM12] Siddharth Garg, Diana Marculescu, and Radu Marculescu. Technology-driven limits on runtime power management algorithms for multiprocessor systems-on-chip. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(4):28:1– 28:??, October 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Guiducci:2008:HPP

[GN08] Carlotta Guiducci and Christine Nardini. High parallelism, portability, and broad accessibility: Technologies for genomics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(1):3:1–3:??, March 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Gebregiorgis:2022:SMC

[GNY^{+}22] Anteneh Gebregiorgis, Hoang Anh Du Nguyen, Jintao Yu, Rajendra Bishnoi, Mottaqiallah Taouil, Francky Catthoor, and Said Hamdioui. A survey on memory-centric computer architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):79:1– 79:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3544974.

Galceran-Oms:2011:MTU

[GOGCK11] Marc Galceran-Oms, Alexander Gotmanov, Jordi Cortadella, and Mike Kishinevsky. Microarchitectural transformations using elasticity. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(4):18:1– 18:??, December 2011. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Gorantla:2017:DAC

[GP17] Anusha Gorantla and Deepa P. Design of approximate compressors for multiplication. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3): 44:1–44:??, May 2017. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Graziano:2015:PVE

[GPW^{+}15] Mariagrazia Graziano, Azzurra Pulimeno, Ruiyu Wang, Xiang Wei, Massimo Ruo Roch, and Gianluca Piccinini. Process variability and electrostatic analysis of molecular QCA. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(2):18:1– 18:??, August 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Ghavami:2013:DAR

[GRPT13] Behnam Ghavami, Mohsen Raji, Hossein Pedram, and Mehdi B. Tahoori. Design and analysis of a robust carbon nanotube-based asynchronous primitive circuit. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(1):4:1–4:??, February 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Gojman:2005:EDS

[GRS05] Benjamin Gojman, Eric Rachlin, and John E. Savage. Evaluation of design strategies for stochastically assembled nanoarray memories. ACM Journal on Emerging Technologies in Computing Systems (JETC), 1(2): 73–108, 2005. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Guha:2017:RTS

[GSC17] Krishnendu Guha, Debasri Saha, and Amlan Chakrabarti. Real-time SoC security against passive threats using crypsis behavior of geckos. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):41:1– 41:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Guiducci:2013:ISI

[Gui13] Carlotta Guiducci. Introduction to special issue on bioinformatics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(4):26:1–26:??, November 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Goren:2011:DAN

[GUP11] Sezer Gören, H. Fatih Ugurdag, and Okan Palaz. Defect-aware nanocrossbar logic mapping through matrix canonization using two-dimensional radix sort. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(3): 12:1–12:??, August 2011. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Goud:2017:AUF

[GVRR17] A. Arun Goud, Rangharajan Venkatesan, Anand Raghunathan, and Kaushik Roy.

Asymmetric underlapped Fin-FETs for near- and superthreshold logic at sub-10nm technology nodes. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):23:1– 23:??, March 2017. CODEN ???? ISSN 1550-4832.

Golnari:2017:PCE

[GYM^{+}17] Pareesa Ameneh Golnari, Yavuz Yetim, Margaret Martonosi, Yakir Vizel, and Sharad Malik. PPU: a control errortolerant processor for streaming applications with formal guarantees. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):43:1–43:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Guo:2021:RMS

[GZZ^{+}21] Chuliang Guo, Li Zhang, Xian Zhou, Grace Li Zhang, Bing Li, Weikang Qian, Xunzhao Yin, and Cheng Zhuo. A reconfigurable multiplier for signed multiplications with asymmetric bit-widths. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):48:1– 48:16, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3446213.

Han:2022:PSO

[HAV^{+}22] Haiyang Han, Theoni Alexoudi, Chris Vagionas, Nikos Pleros, and Nikos Hardavellas. A practical shared optical cache with hybrid MWSR/R-SWMR NoC for multicore processors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):76:1– 76:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3531012.

Hossain:2015:MGN

[HC15] Nahid M. Hossain and Masud H. Chowdhury. Multilayer graphene nanoribbon and carbon nanotube based floating gate transistor for nonvolatile flash memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(1):2:1–2:??, July 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Huang:2008:RAF

[HCTK08] Tsung-Ching Huang, Kwang-Ting (Tim) Cheng, Huai-Yuan Tseng, and Chen-Pang Kung. Reliability analysis for flexible electronics: Case study of integrated a-Si:H TFT scan driver. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(3):12:1–12:??, August 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Huang:2014:FMD

[HD14] Haiyao Huang and Douglas Densmore. Fluigi: Microfluidic device synthesis for synthetic biology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3):26:1–26:??, December 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Han:2022:PBC

[HFLZ22] Jianhui Han, Xiang Fei, Zhaolin Li, and Youhui Zhang. Polyhedral-based compilation framework for inmemory neural network accelerators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):15:1–15:23, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3469847.

Haron:2011:RRN

[HH11] Nor Zaidi Haron and Said Hamdioui. Redundant residue number system code for faulttolerant hybrid memories. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(1):4:1– 4:??, January 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Huang:2023:LRG

[HHD^{+}23] Siyuan Huang, Brian D.

Hoskins, Matthew W. Daniels, Mark D. Stiles, and Gina C. Adam. Low-rank gradient descent for memory-efficient training of deep in-memory arrays. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(2):16:1–16:??, April 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3577214.

Hamilton:2018:SHE

[HIH18] Kathleen E. Hamilton, Neena Imam, and Travis S. Humble. Sparse hardware embedding of spiking neuron systems for community detection. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4): 40:1–40:??, December 2018. CODEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3223048.

Huang:2012:IRD

[HLH^{+}12] Yu-Jie Huang, Hsin-Hung Liao, Pen-Li Huang, Tao Wang, Yao-Joe Yang, Yao-Hong Wang, and Shey-Shi Lu. An implantable releaseon-demand CMOS drug delivery SoC using electrothermal activation technique. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(2):12:1– 12:??, June 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Hanninen:2014:QII

[HLS14] Ismo K. Hänninen, Craig S.

Lent, and Gregory L. Snider. Quantifying irreversible information loss in digital circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):10:1– 10:??, November 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Hadjam:2014:RED

[HM14] Fatima Zohra Hadjam and Claudio Moraga. RIMEP2: Evolutionary design of reversible digital circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3):27:1– 27:??, December 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Hosseini:2021:BPN

[HM21] Morteza Hosseini and Tinoosh Mohsenin. Binary precision neural network manycore accelerator. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):19:1–19:27, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3423136.

Huang:2022:DAA

[HMP^{+}22] Tianjin Huang, Vlado Menkovski, Yulong Pei, Yuhao Wang, and Mykola Pechenizkiy. Direction-aggregated attack for transferable adversarial examples. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):60:1–60:22, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3501769.

Huang:2005:TBQ

[HMS^{+}05] J. Huang, M. Momenzadeh, L. Schiano, M. Ottavi, and F. Lombardi. Tile-based QCA design using majoritylike logic primitives. ACM Journal on Emerging Technologies in Computing Systems (JETC), 1(3):163–185, October 2005. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Henry:2012:TNH

[HN12] Michael B. Henry and Leyla Nazhandali. From transistors to NEMS: Highly efficient power-gating of CMOS circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(1):2:1–2:??, February 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Hammerstrom:2015:ISI

[HN15] Dan Hammerstrom and Vijaykrishnan Narayanan. Introduction to special issue on neuromorphic computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4):32:1– 32:??, April 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Hazari:2021:MLV

[HON21] Noor Ahmad Hazari, Ahmed Oun, and Mohammed Niamat. Machine learning vulnerability analysis of FPGAbased ring oscillator PUFs and counter measures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):36:1– 36:20, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3445978.

Howladar:2019:HPH

[HRR19] Pampa Howladar, Pranab Roy, and Hafizur Rahaman. A high-performance homogeneous droplet routing technique for MEDA-based biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):38:1– 38:??, December 2019. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3327965.

Houshmand:2017:QCS

[HSZM17] Mahboobeh Houshmand, Mehdi Sedighi, Morteza Saheb Zamani, and Kourosh Marjoei. Quantum circuit synthesis targeting to improve one-way quantum computation pattern cost metrics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):55:1– 55:??, August 2017. CODEN ???? ISSN 1550-4832.

Hajkazemi:2018:HHM

[HTMH18] Mohammad Hossein Hajkazemi, Mohammad Khavari Tavana, Tinoosh Mohsenin, and Houman Homayoun. Heterogeneous HMC + DDRx memory management for performance-temperature tradeoffs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1): 4:1–4:??, March 2018. CO-DEN ???? ISSN 1550-4832.

Hamdioui:2022:GEC

[HVB22] Said Hamdioui, Elena-Ioana Vatajelu, and Alberto Bosio. Guest editorial: Computationin-memory (CIM): from device to applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):31:1– 31:3, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3503263.

He:2020:SBN

[HYA^{+}20] Zhezhi He, Li Yang, Shaahin Angizi, Adnan Siraj Rakin, and Deliang Fan. Sparse BD-Net: a multiplication-less DNN with sparse binarized depth-wise separable convolution. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2): 15:1–15:24, April 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3369391.

Huang:2022:GBB

[HYPW22] Po-Hao Huang, Honggang Yu, Max Panoff, and Ting-Chi Wang. Generation of blackbox audio adversarial examples based on gradient approximation and autoencoders. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):59:1– 59:19, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3491220.

Huo:2009:SBN

[HYWA09] Dennis Huo, Qiaoyan Yu, David Wolpert, and Paul Ampadu. A simulator for ballistic nanostructures in a 2-D electron gas. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(1):5:1–5:??, January 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Houshmand:2014:DDH

[HZSA14] Mahboobeh Houshmand, Morteza Saheb Zamani, Mehdi Sedighi, and Mona Arabzadeh. Decomposition of diagonal Hermitian quantum gates using multiple-controlled Pauli Z gates. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3):28:1–28:??, December 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Huang:2012:PDT

[HZY^{+}12] Jiale Huang, Minhao Zhu, Shengqi Yang, Pallav Gupta, Wei Zhang, Steven M. Rubin, Gilda Garretón, and Jin He. A physical design tool for carbon nanotube field-effect transistor circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):25:1–25:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Islam:2020:FPM

[IASK20] Mahzabeen Islam, Shashank Adavally, Marko Scrbak, and Krishna Kavi. On-the-fly page migration and address reconciliation for heterogeneous memory systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):10:1– 10:27, February 2020. CO-DEN ???? ISSN 1550-4832.

URL https://dl.acm.org/
doi/abs/10.1145/3364179.

Ince:2022:DFB

[IBO22] Mehmet Ince, Bora Bilgic, and Sule Ozev. Digital faultbased built-in self-test and evaluation of low dropout voltage regulators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):54:1– 54:20, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3510852.

Imani:2019:HSC

[IGGR19] Mohsen Imani, Ricardo Garcia, Saransh Gupta, and Tajana Rosing. Hardwaresoftware co-design to accelerate neural network applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(2):21:1–21:??, June 2019. CODEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3304086.

Iyengar:2016:SPS

[IGR^{+}16] Anirudh Iyengar, Swaroop Ghosh, Kenneth Ramclam, Jae-Won Jang, and Cheng-Wei Lin. Spintronic PUFs for security, trust, and authentication. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1):4:1–4:??, December 2016. CODEN ???? ISSN 1550-4832.

Irwin:2005:E

[IN05] Mary Jane Irwin and Vijaykrishnan Narayanan. Editorial. ACM Journal on Emerging Technologies in Computing Systems (JETC), 1(1):1– 6, April 2005. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Ishihara:2018:INP

[ISI^{+}18] Tohru Ishihara, Akihiko Shinya, Koji Inoue, Kengo Nozaki, and Masaya Notomi. An integrated nanophotonic parallel adder. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):26:1–26:??, July 2018. CODEN ???? ISSN 1550-4832.

Javaheripi:2023:AAH

[JCK23] Mojan Javaheripi, Jung-Woo Chang, and Farinaz Koushanfar. AccHashtag: Accelerated hashing for detecting fault-injection attacks on embedded neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(1):7:1– 7:??, January 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3555808.

Jessurun:2023:FNS

[JDPH^{+}23] Nathan Jessurun, Olivia P. Dizon-Paradis, Jacob Harrison, Shajib Ghosh, Mark M. Tehranipoor, Damon L. Woodard, and Navid Asadizanjani. FPIC: a novel semantic dataset for optical PCB assurance. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(2):17:1–17:??, April 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3588032.

Jin:2022:GET

[JHPG22] Yier Jin, Tsung-Yi Ho, Stjepan Picek, and Siddharth Garg. Guest editorial: Trustworthy AI. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):55:1–55:3, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3534957.

Jiang:2017:RCC

[JLL^{+}17] Honglan Jiang, Cong Liu, Leibo Liu, Fabrizio Lombardi, and Jie Han. A review, classification, and comparative evaluation of approximate arithmetic circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):60:1– 60:??, August 2017. CODEN ???? ISSN 1550-4832.

Jha:2021:DED

[JMKM21] Nandan Kumar Jha, Sparsh Mittal, Binod Kumar, and Govardhan Mattela. DeepPeep:

Exploiting design ramifications to decipher the architecture of compact DNNs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(1):5:1– 5:25, January 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414552.

Jafri:2015:AID

[JOF^{+}15] Syed M. A. H. Jafri, Ozan Ozbag, Nasim Farahini, Kolin Paul, Ahmed Hemani, Juha Plosila, and Hannu Tenhunen. Architecture and implementation of dynamic parallelism, voltage and frequency scaling (PVFS) on CGRAs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4):40:1– 40:??, April 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Jackson:2013:NES

[JRC^{+}13] Bryan L. Jackson, Bipin Rajendran, Gregory S. Corrado, Matthew Breitwisch, Geoffrey W. Burr, Roger Cheek, Kailash Gopalakrishnan, Simone Raoux, Charles T. Rettner, Alvaro Padilla, Alex G. Schrott, Rohit S. Shenoy, Bülent N. Kurdi, Chung H. Lam, and Dharmendra S. Modha. Nanoscale electronic synapses using phase change devices. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(2):12:1–12:??, May 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Jones:2022:SNA

[JRJ22] Alexander Jones, Aaron Ruen, and Rashmi Jha. A spiking neuromorphic architecture using Gated-RRAM for associative memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):36:1– 36:22, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3461667.

Jayakumar:2015:QHS

[JRLR15] Hrishikesh Jayakumar, Arnab Raha, Woo Suk Lee, and Vijay Raghunathan. QuickRecall: a HW/SW approach for computing across power cycles in transiently powered computers. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(1):8:1–8:??, July 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Jiang:2017:SLD

[JWJ^{+}17] Wei Jiang, Liang Wen, Ke Jiang, Xia Zhang, Xiong Pan, and Keran Zhou. System-level design to detect fault injection attacks on embedded real-time applications.

ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):22:1– 22:??, March 2017. CODEN ???? ISSN 1550-4832.

Kim:2021:SDB

[KAB^{+}21] Heewoo Kim, Aporva Amarnath, Javad Bagherzadeh, Nishil Talati, and Ronald G. Dreslinski. A survey describing beyond Si transistors and exploring their implications for future processors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):27:1– 27:44, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3453143.

Kamal:2014:IPV

[KAKSP14] Mehdi Kamal, Ali Afzali-Kusha, Saeed Safari, and Massoud Pedram. Impact of process variations on speedup and maximum achievable frequency of extensible processors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(3): 19:1–19:??, April 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Karri:2020:E

[Kar20] Ramesh Karri. Editorial. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):1:1, February 2020. CODEN ????

ISSN 1550-4832. URL https:

//dl.acm.org/doi/abs/10.
1145/3378024.

Khammassi:2022:OPQ

[KAS^{+}22] N. Khammassi, I. Ashraf, J. V. Someren, R. Nane, A. M. Krol, M. A. Rol, L. Lao, K. Bertels, and C. G. Almudever. OpenQL: a portable quantum programming framework for quantum accelerators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18 (1):13:1–13:24, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3474222.

Koblah:2023:FOD

[KBC^{+}23] David Selasi Koblah, Ulbert J. Botero, Sean P. Costello, Olivia P. Dizon-Paradis, Fatemeh Ganji, Damon L. Woodard, and Domenic Forte. A fast object detection-based framework for via modeling on PCB X-ray CT images. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(4):34:1– 34:??, October 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3606948.

Kulkarni:2021:AAI

[KBM21] Sourabh Kulkarni, Sachin Bhat, and Csaba Andras Moritz. Architecting for artificial intelligence with emerging nanotechnology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):26:1– 26:33, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3445977.

Karempudi:2024:AVD

[KBT24] Venkata Sai Praneeth Karempudi, Janibul Bashir, and Ishan G. Thakkar. An analysis of various design pathways towards multi-terabit photonic on-interposer interconnects. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(2):6:1–6:??, April 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3635031.

Kufluoglu:2014:RMN

[KCC^{+}14] Haldun Küflüoglu, Cathy Chancellor, Min Chen, Claude Cirba, and Vijay Reddy. Recovery modeling of negative bias temperature instability (NBTI) for SPICEcompatible circuit aging simulators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):2:1–2:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Krichmar:2015:LSS

[KCD15] Jeffrey L. Krichmar, Philippe Coussy, and Nikil Dutt. Large-scale spiking neural networks using neuromorphic hardware compatible models. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4):36:1– 36:??, April 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kim:2016:CAP

[KCWL^{+}16] Moon Seok Kim, William Cane-Wissing, Xueqing Li, Jack Sampson, Suman Datta, Sumeet Kumar Gupta, and Vijaykrishnan Narayanan. Comparative area and parasitics analysis in FinFET and heterojunction vertical TFET standard cells. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):38:1–38:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Karkar:2022:TPE

[KDMT22] Ammar Karkar, Nizar Dahir, Terrence Mak, and Kin-Fai Tong. Thermal and performance efficient on-chip surface-wave communication for many-core systems in dark silicon era. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):49:1–49:18, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3501771.

Kole:2024:EEN

[KDSD24] Abhoy Kole, Kamalika Datta, Indranil Sengupta, and Rolf Drechsler. Exploiting the extended neighborhood of hexagonal qubit architecture for mapping quantum circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(4):13:1– 13:??, October 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3688391.

Kooli:2022:TTI

[KHC^{+}22] Maha Kooli, Antoine Heraud, Henri-Pierre Charles, Bastien Giraud, Roman Gauchi, Mona Ezzadeen, Kevin Mambu, Valentin Egloff, and Jean-Philippe Noel. Towards a truly integrated vector processing unit for memory-bound applications based on a costcompetitive computational SRAM design solution. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):40:1– 40:26, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3485823.

Khasanvis:2015:LPH

[KHR^{+}15] Santosh Khasanvis, K. M. Masum Habib, Mostafizur Rahman, Roger Lake, and Csaba Andras Moritz. Lowpower heterogeneous graphene nanoribbon-CMOS multistate volatile memory circuit. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(2):15:1– 15:??, August 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Komerath:2012:RBP

[KK12] Narayanan Komerath and Aravinda Kar. Retail beamed power using millimeter waves: Survey. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):18:1–18:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kim:2023:ISI

[KK23] John Kim and Tushar Krishna. Introduction to the special issue on next-generation on-chip and off-chip communication architectures for edge, cloud and HPC. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(4):31:1– 31:??, October 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3631144.

Koneru:2017:IEC

[KKC17] Abhishek Koneru, Sukeshwar Kannan, and Krishnendu Chakrabarty. Impact of electrostatic coupling and wafer-bonding defects on delay testing of monolithic 3D integrated circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):54:1– 54:??, August 2017. CODEN ???? ISSN 1550-4832.

Kim:2022:EPS

[KKK22] Yulhwa Kim, Hyungjun Kim, and Jae-Joon Kim. Extreme partial-sum quantization for analog computing-in-memory neural network accelerators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):75:1– 75:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3528104.

Kim:2018:DNN

[KKKK18] Hyungjun Kim, Taesu Kim, Jinseok Kim, and Jae-Joon Kim. Deep neural network optimized to resistive memory with nonlinear currentvoltage characteristics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):15:1– 15:??, July 2018. CODEN ???? ISSN 1550-4832.

Kulkarni:2022:HAS

[KKNM22] Sourabh Kulkarni, Mario Michael Krell, Seth Nabarro, and Csaba Andras Moritz. Hardwareaccelerated simulation-based inference of stochastic epidemiology models for COVID-19. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2): 30:1–30:24, April 2022. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3471188.

Kim:2020:HSC

[KKY^{+}20] Youngseok Kim, Seyoung Kim, Chun-Chen Yeh, Vijay Narayanan, and Jungwook Choi. Hardware and software co-optimization for the initialization failure of the ReRAMbased cross-bar array. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(4):36:1– 36:19, October 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3393669.

Krishnan:2022:ICI

[KMC^{+}22] Gokul Krishnan, Sumit K.

Mandal, Chaitali Chakrabarti, Jae-Sun Seo, Umit Y. Ogras, and Yu Cao. Impact of on-chip interconnect on inmemory acceleration of deep neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):34:1– 34:22, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3460233.

Kant:2012:EDC

[KMD12] Krishna Kant, Muthukumar Murugan, and David H. C. Du. Enhancing data center sustainability through energy-adaptive computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(4):33:1– 33:??, October 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kim:2020:PSA

[KMS^{+}20] J. Hyun Kim, Young Je Moon, Hyunsub Song, Jay H. Park, and Sam H. Noh. On providing OS support to allow transparent use of traditional programming models for persistent memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3):33:1–33:24, July 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3388637.

Ko:2012:EHC

[Ko12] Wen H. Ko. Early history and challenges of implantable electronics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(2):8:1–8:??, June 2012. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kocak:2010:IDT

[KP10] Taskin Kocak and Dhiraj Pradhan. Introduction to design techniques for energy harvesting. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(2):4:1–4:??, June 2010. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kulkarni:2016:RTA

[KPFM16] Amey Kulkarni, Youngok Pino, Matthew French, and Tinoosh Mohsenin. Real-time anomaly detection framework for many-core router through machine-learning techniques. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1):10:1– 10:??, December 2016. CO-DEN ???? ISSN 1550-4832.

K:2022:CHA

[KPM22] Amal Thomas K., Soumyajit Poddar, and Hemanta Kumar Mondal. A CNN hardware accelerator using trianglebased convolution. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):78:1– 78:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3544975.

Karam:2017:MCR

[KPPB17] Robert Karam, Somnath Paul, Ruchir Puri, and Swarup Bhunia. Memorycentric reconfigurable accelerator for classification and machine learning applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):34:1– 34:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Ko:2018:RTL

[KR18] Glenn G. Ko and Rob A. Rutenbar. Real-time and lowpower streaming source separation using Markov random field. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2): 17:1–17:??, July 2018. CO-DEN ???? ISSN 1550-4832.

Koylu:2023:SML

[KRG^{+}23] Troya Çagil Köylü, Cezar Rodolfo Wedig Reinbrecht, Anteneh Gebregiorgis, Said Hamdioui, and Mottaqiallah Taouil. A survey on machine learning in hardware security. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(2):18:1– 18:??, April 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3589506.

Khatwani:2021:FME

[KRP^{+}21] Mohit Khatwani, Hasib-Al Rashid, Hirenkumar Paneliya, Mark Horton, Nicholas Waytowich, W. David Hairston, and Tinoosh Mohsenin. A flexible multichannel EEG artifact identification processor using depthwise-separable convolutional neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):23:1–23:21, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3427471.

Ku:2022:UDR

[KSA^{+}22] Bon Woong Ku, Catherine D. Schuman, Md Musabbir Adnan, Tiffany M. Mintz, Raphael Pooser, Kathleen E. Hamilton, Garrett S. Rose, and Sung Kyu Lim. Unsupervised digit recognition using cosine similarity in a neuromemristive competitive learning system. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):38:1– 38:20, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3473036.

Kgil:2008:PUS

[KSB^{+}08] Taeho Kgil, Ali Saidi, Nathan Binkert, Steve Reinhardt, Krisztian Flautner, and Trevor Mudge. PicoServer: Using 3D stacking technology to build energy efficient servers. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(4):16:1– 16:??, October 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kumawat:2014:PMA

[KSG14] Renu Kumawat, Vineet Sahula, and Manoj S. Gaur. Probabilistic modeling and analysis of molecular memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(1):6:1–6:??, September 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Karempudi:2022:PNC

[KST^{+}22] Venkata Sai Praneeth Karempudi, Febin Sunny, Ishan G. Thakkar, Sai Vineel Reddy Chittamuru, Mahdi Nikdast, and Sudeep Pasricha. Photonic networks-on-chip employing multilevel signaling: a cross-layer comparative study. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):45:1– 45:36, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3487365.

Kim:2014:ICU

[KT14] Jaeyoon Kim and Sandip Tiwari. Inexact computing using probabilistic circuits: Ultra low-power digital processing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(2):

16:1–16:??, February 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kumar:2020:DAL

[KT20] S. Dinesh Kumar and Himanshu Thapliyal. Design of adiabatic logic-based energyefficient and reliable PUF for IoT devices. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3):34:1–34:18, July 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3390771.

Kuo:2008:MSA

[KTW08] Shih-Hsien Kuo, Bruce Tidor, and Jacob White. A meshless, spectrally accurate, integral equation solver for molecular surface electrostatics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(2):6:1–6:??, April 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kang:2020:CCL

[KWC^{+}20] Wang Kang, Bi Wu, Xing Chen, Daoqian Zhu, Zhaohao Wang, Xichao Zhang, Yan Zhou, Youguang Zhang, and Weisheng Zhao. A comparative cross-layer study on racetrack memories: Domain wall vs Skyrmion. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):2:1–2:17, February 2020. CODEN ???? ISSN 1550-4832.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3333336.

Kursun:2012:STT

[KWFH12] Eren Kursun, Jamil Wakil, Mukta Farooq, and Robert Hannon. Spatial and temporal thermal characterization of stacked multicore architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):21:1–21:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kang:2020:AAS

[KWG^{+}20] Ziyang Kang, Lei Wang, Shasha Guo, Rui Gong, Shiming Li, Yu Deng, and Weixia Xu. ASIE: an asynchronous SNN inference engine for AER events processing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(4):39:1– 39:22, October 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3404992.

Kumar:2017:THS

[KWWI17] Arvind Kumar, Zhe Wan, Winfried W. Wilcke, and Subramanian S. Iyer. Toward human-scale brain computing using 3D wafer scale integration. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3): 45:1–45:??, May 2017. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Khouzani:2016:FEP

[KXY16] Hoda Aghaei Khouzani, Yuan Xue, and Chengmo Yang. Fully exploiting PCM write capacity within near zero cost through segment-based page allocation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):31:1–31:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Khayambashi:2015:ARA

[KYEB15] Misagh Khayambashi, Pooria M. Yaghini, Ashkan Eghbal, and Nader Bagherzadeh. Analytical reliability analysis of 3D NoC under TSV failure. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4): 43:1–43:??, April 2015. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kim:2015:RDN

[KZL15] Yongtae Kim, Yong Zhang, and Peng Li. A reconfigurable digital neuromorphic processor with memristive synaptic crossbar for cognitive computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4): 38:1–38:??, April 2015. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Kang:2015:SEU

[KZW^{+}15] Wang Kang, Yue Zhang, Zhaohao Wang, Jacques-Olivier Klein, Claude Chappert, Dafiné Ravelosona, Gefei Wang, Youguang Zhang, and Weisheng Zhao. Spintronics: Emerging ultra-low-power circuits and systems beyond MOS technology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(2):16:1– 16:??, August 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Lu:2018:RHM

[LBB^{+}18] Guan-Ruei Lu, Ansuman Banerjee, Bhargab B. Bhattacharya, Tsung-Yi Ho, and Hung-Ming Chen. Reliability hardening mechanisms in cyber-physical digitalmicrofluidic biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(3):34:1– 34:??, October 2018. CODEN ???? ISSN 1550-4832.

Li:2008:ADP

[LBGR08] Jing Li, Aditya Bansal, Swarop Ghosh, and Kaushik Roy. An alternate design paradigm for low-power, lowcost, testable hybrid systems using scaled LTPS TFTs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(3):13:1– 13:??, August 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Layer:2016:RSP

[LBJ^{+}16] Christophe Layer, Laurent Becker, Kotb Jabeur, Sylvain Claireux, Bernard Dieny, Guillaume Prenat, Gregory Di Pendina, Stephane Gros, Pierre Paoli, Virgile Javerliac, Fabrice Bernard-Granger, and Loic Decloedt. Reducing system power consumption using check-pointing on nonvolatile embedded magnetic random access memories. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):32:1– 32:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Lebeck:2008:IDS

[LC08] Alvin R. Lebeck and Krishnendu Chakrabarty. Introduction to DAC 2007 special section. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(3):10:1–10:??, August 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Lin:2014:QQM

[LCJ14] Chia-Chun Lin, Amlan Chakrabarti, and Niraj K. Jha. QLib: Quantum module library. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(1):7:1– 7:??, September 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Louri:2019:LHS

[LCK19] Ahmed Louri, Jacques Collet, and Avinash Karanth. Limit of hardware solutions for self-protecting faulttolerant NoCs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):4:1–4:??, February 2019. CODEN ???? ISSN 1550-4832.

Li:2014:SAB

[LCSP14] Zhiqiang Li, Hanwu Chen, Xiaoyu Song, and Marek Perkowski. A synthesis algorithm for 4-bit reversible logic circuits with minimum quantum cost. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3):29:1–29:??, December 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Liu:2012:RAP

[LCT12] Bao Liu, Xuemei Chen, and Fiona Teshome. Resilient and adaptive performance logic. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):22:1– 22:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Liu:2019:MML

[LCY19] Qixiao Liu, Zhifeng Chen, and Zhibin Yu. MiC: Multilevel characterization and optimization of GPGPU kernels. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(3): 25:1–25:??, June 2019. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3304108.

Lee:2018:DSE

[LDK^{+}18] Dongjin Lee, Sourav Das, Dae Hyun Kim, Janardhan Rao Doppa, and Partha Pratim Pande. Design space exploration of 3D network-on-chip: a sensitivity-based optimization approach. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(3):32:1– 32:??, October 2018. CODEN ???? ISSN 1550-4832.

Liu:2010:RSO

[LDL10] Yang Liu, Chris Dwyer, and Alvin R. Lebeck. Routing in self-organizing nano-scale irregular networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(1):3:1–3:??, March 2010. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Li:2020:RRB

[LDP^{+}20] Bing Li, Janardhan Rao Doppa, Partha Pratim Pande, Krishnendu Chakrabarty, Joe X. Qiu, and Hai (Helen) Li. 3D-ReG: a 3D ReRAM-based heterogeneous architecture for training deep neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2):20:1– 20:24, April 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3375699.

Laurent:2021:BGB

[LDPPB21] J. Laurent, C. Deleuze, F. Pebay-Peyroula, and V. Beroulle. Bridging the gap between RTL and software fault injection. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3): 38:1–38:24, July 2021. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3446214.

Liu:2022:DAA

[LDZ22] Yuan Liu, Jinxin Dong, and Pingqiang Zhou. Defending against adversarial attacks in deep learning with robust auxiliary classifiers utilizing bit-plane slicing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):61:1– 61:17, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3510855.

Li:2022:HIH

[LFDS22] Weifu Li, Paul Franzon, Sumon Dey, and Joshua Schabel. Hardware implementation of hierarchical temporal memory algorithm. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):17:1– 17:23, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3479430.

Li:2015:ICI

[LGL15] Zhongqi Li, Nilanjan Goswami, and Tao Li. iConn: a communication infrastructure for heterogeneous computing architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4):42:1–42:??, April 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Lu:2021:LNT

[LGYC21] Yunfeng Lu, Huaxi Gu, Xiaoshan Yu, and Krishnendu Chakrabarty. Lotus: a new topology for large-scale distributed machine learning. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(1):7:1– 7:21, January 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3415749.

Li:2020:LNL

[LH20] Sumin Li and Linpeng Huang. LosPem: a novel log-structured framework for persistent memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3):27:1–27:17, July 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3379932.

Li:2019:LLP

[LHHZ19] Sumin Li, Kaixin Huang, Linpeng Huang, and Jiashun Zhu. LiwePMS: a lightweight persistent memory with wearaware memory management. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(3): 30:1–30:??, June 2019. CO-DEN ???? ISSN 1550-

                                        4832. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3327963.

Li:2017:PUD

[LHW^{+}17] Bing Li, Yu Hu, Ying Wang, Jing Ye, and Xiaowei Li. Power-utility-driven write management for MLC PCM. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):50:1– 50:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Lee:2010:FBP

[LJ10] Chun-Yi Lee and Niraj K.

Jha. FinFET-based power simulator for interconnection networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(1):2:1–2:??, March 2010. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Lin:2014:RRM

[LJ14] Chia-Chun Lin and Niraj K.

Jha. RMDDS: Reed–Muller decision diagram synthesis of reversible logic circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(2):14:1– 14:??, February 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Liu:2018:OAE

[LJL18] Yu Liu, Yingyezhe Jin, and Peng Li. Online adaptation and energy minimization for hardware recurrent spiking neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1):11:1– 11:??, March 2018. CODEN ???? ISSN 1550-4832.

Lee:2015:REE

[LKC15] Jinho Lee, Kyungsu Kang, and Kiyoung Choi. REDELF: an energy-efficient deadlockfree routing for 3D NoCs with partial vertical connections. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3):26:1– 26:??, September 2015. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Lee:2022:RMC

[LKK^{+}22] Edward Lee, Daehyun Kim, Jinwoo Kim, Sung Kyu Lim, and Saibal Mukhopadhyay. A ReRAM memory compiler for monolithic 3D integrated circuits in a carbon nanotube process. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):20:1– 20:20, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3466681.

Luo:2018:OOW

[LKL^{+}18] Jiating Luo, Cedric Killian, Sebastien Le Beux, Daniel Chillet, Olivier Sentieys, and Ian O’Connor. Offline optimization of wavelength allocation and laser power in nanophotonic interconnects. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):24:1– 24:??, July 2018. CODEN ???? ISSN 1550-4832.

Li:2017:EEC

[LLSO17] Hui Li, Sébastien Le Beux, Martha Johanna Sepulveda, and Ian O’Connor. Energyefficiency comparison of multilayer deposited nanophotonic crossbar interconnects. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):61:1– 61:??, August 2017. CODEN ???? ISSN 1550-4832.

Li:2018:GOF

[LLX^{+}18a] Yixing Li, Zichuan Liu, Kai Xu, Hao Yu, and Fengbo Ren. A GPU-outperforming FPGA accelerator architecture for binary convolutional neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):18:1–18:??, July 2018. CODEN ???? ISSN 1550-4832.

Liu:2018:MLO

[LLX^{+}18b] Zhongyang Liu, Shaoheng Luo, Xiaowei Xu, Yiyu Shi, and Cheng Zhuo. A multilevel-optimization framework for FPGA-based cellular neural network implementation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4):47:1– 47:??, December 2018. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3273957.

Lee:2013:CIP

[LM13] Woo Hyung Lee and Pinaki Mazumder. Color image processing with multi-peak resonant tunneling diodes. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(3):18:1– 18:??, September 2013. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Li:2011:IHN

[LMC^{+}11] Zheng Li, Moustafa Mohamed, Xi Chen, Hongyu Zhou, Alan Mickelson, Li Shang, and Manish Vachharajani. Iris: a hybrid nanophotonic network design for highperformance and low-power on-chip communication. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(2):8:1–8:??, June 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Li:2019:TBH

[LML^{+}19] Bing Li, Mengjie Mao, Xiaoxiao Liu, Tao Liu, Zihao Liu, Wujie Wen, Yiran Chen, and Hai (Helen) Li. Thread batching for high-performance energy-efficient GPU memory design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):39:1–39:??, December 2019. CODEN ???? ISSN 1550-4832. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3330152.

Loomis:2018:FIT

[LMM18] Lisa Loomis, Nathan McDonald, and Cory Merkel. An FPGA implementation of a time delay reservoir using stochastic logic. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4):46:1–46:??, December 2018. CODEN ???? ISSN 1550-4832.

Li:2024:SNS

[LMZZ24] Ziwen Li, Yu Ma, Jindong Zhou, and Pingqiang Zhou. Spiking-NeRF: Spiking neural network for energy-efficient neural rendering. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(3):10:1– 10:??, July 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3675808.

Li:2019:LCS

[LNL19] Bingzhe Li, M. Hassan Najafi, and David J. Lilja. Lowcost stochastic hybrid multiplier for quantized neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(2):18:1–18:??, June 2019. CODEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3309882.

Liu:2017:CPU

[LP17] Yin Liu and Keshab K. Parhi.

Computing polynomials using unipolar stochastic logic. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):42:1– 42:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Le:2015:END

[LPB^{+}15] Trong Nhan Le, Alain Pegatoquet, Olivier Berder, Olivier Sentieys, and Arnaud Carer. Energy-neutral design framework for supercapacitorbased autonomous wireless sensor networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(2):19:1– 19:??, August 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Lou:2019:MSA

[LPM^{+}19] Qiuwen Lou, Chenyun Pan, John McGuinness, Andras Horvath, Azad Naeemi, Michael Niemier, and X. Sharon Hu. A mixed signal architecture for convolutional neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(2):19:1–19:??, June 2019. CODEN ???? ISSN 1550-

                                        4832. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3304110.

Li:2018:ELC

[LPW18] Bohua Li, Yukui Pei, and Wujie Wen. Efficient LDPC code design for combating asymmetric errors in STT-RAM. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1):10:1– 10:??, March 2018. CODEN ???? ISSN 1550-4832.

Li:2019:NNC

[LQYL19] Bingzhe Li, Yaobin Qin, Bo Yuan, and David J. Lilja. Neural network classifiers using a hardware-based approximate activation function with a hybrid stochastic multiplier. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):12:1– 12:??, February 2019. CO-DEN ???? ISSN 1550-4832.

Lim:2005:PPB

[LRN05] Sung Kyu Lim, Ramprasad Ravichandran, and Mike Niemier. Partitioning and placement for buildable QCA circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 1(1):50–72, April 2005. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Lin:2014:NTL

[LSH14] Ing-Chao Lin, Shun-Ming Syu, and Tsung-Yi Ho. NBTI tolerance and leakage reduction using gate sizing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(1):4:1–4:??, September 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Lao:2016:BFD

[LTKP16] Yingjie Lao, Qianying Tang, Chris H. Kim, and Keshab K. Parhi. Beat frequency detector-based highspeed true random number generators: Statistical modeling and analysis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1):9:1–9:??, December 2016. CODEN ???? ISSN 1550-4832.

Labrado:2022:FVS

[LTM22] Carson Labrado, Himanshu Thapliyal, and Saraju P. Mohanty. Fortifying vehicular security through low overhead physically unclonable functions. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):8:1–8:18, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3442443.

Li:2022:PBA

[LTO22] Ge Li, Mohit Tiwari, and Michael Orshansky. Powerbased attacks on spatial DNN accelerators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):58:1–58:18, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3491219.

Lin:2014:POF

[LWH14] Jiun-Li Lin, Po-Hsun Wu, and Tsung-Yi Ho. Placement optimization of flexible TFT circuits with mechanical strain and temperature consideration. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(1):1:1–1:??, September 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Liu:2014:CHP

[LWM^{+}14] Wulong Liu, Yu Wang, Yuchun Ma, Yuan Xie, and Huazhong Yang. On-chip hybrid power supply system for wireless sensor nodes. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(3):23:1– 23:??, April 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Liu:2014:CSN

[LWX^{+}14] Weichen Liu, Xuan Wang, Jiang Xu, Wei Zhang, Yaoyao Ye, Xiaowen Wu, Mahdi Nikdast, and Zhehui Wang. On-chip sensor networks for soft-error tolerant real-time multiprocessor systems-on-chip. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(2): 15:1–15:??, February 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Luo:2022:ACT

[LWY22] Yandong Luo, Panni Wang, and Shimeng Yu. Accelerating on-chip training with ferroelectric-based hybrid precision synapse. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):35:1–35:20, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3473461.

Lalouani:2022:CMA

[LYEK22] Wassila Lalouani, Mohamed Younis, Mohammad Ebrahimabadi, and Naghmeh Karimi. Countering modeling attacks in PUF-based IoT security solutions. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):46:1–46:28, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3491221.

Liu:2019:EEF

[LYL19] Yu Liu, Sai Sourabh Yenamachintala, and Peng Li. Energy-efficient FPGA spiking neural accelerators with supervised and unsupervised spike-timing-dependent-plasticity. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(3):27:1–27:??, June 2019. CODEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3313866.

Liang:2013:EWB

[LYWW13] Jiale Liang, Stanley Yeh, S. Simon Wong, and H.- S. Philip Wong. Effect of wordline/bitline scaling on the performance, energy consumption, and reliability of cross-point memory array. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(1):9:1– 9:??, February 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Li:2020:ISI

[LZBW20] Helen Li, Wei Zhang, Swarup Bhunia, and Wujie Wen. Introduction to the special issue on new trends in nanoelectronic device, circuit, and architecture design,Part 1. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3): 24:1–24:3, July 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3392080.

Li:2022:QDQ

[LZCX22] Dawei Li, Yang Zhou, Shaopin Chen, and Xiaowei Xu. A quasi-digital QPSK modulator design for biomedical devices. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):28:1–28:16, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3465379.

Liu:2021:RAR

[LZX^{+}21] Yuntao Liu, Michael Zuzak, Yang Xie, Abhishek Chakraborty, and Ankur Srivastava. Robust and attack resilient logic locking with a high applicationlevel impact. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):37:1–37:22, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3446215.

Lee:2022:SAS

[LZXL22] Jeong-Jun Lee, Wenrui Zhang, Yuan Xie, and Peng Li. SaARSP: an architecture for systolic-array acceleration of recurrent spiking neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):68:1–68:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3510854.

Mexis:2021:LAH

[MAC^{+}21] Nico Mexis, Nikolaos Athanasios Anagnostopoulos, Shuai Chen, Jan Bambach, Tolga Arul, and Stefan Katzenbeisser. A lightweight architecture for hardware-based security in the emerging era of systems of systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):43:1– 43:25, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3458824.

Mishra:2022:ICG

[MCH22] Suraj Mishra, Danny Z. Chen, and X. Sharon Hu. Image complexity guided network compression for biomedical image segmentation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):26:1– 26:23, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3471190.

McKee:2007:ESI

[McK07] Sally A. McKee. Editorial to special issue on reliable computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(2):4:1–4:??, July 2007. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Moser:2010:EMF

[MCT10] Clemens Moser, Jian-Jia Chen, and Lothar Thiele. An energy management framework for energy harvesting embedded systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(2):7:1–7:??, June 2010. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Munoz-Coreas:2018:CQO

[MCT18] Edgard Muñoz-Coreas and Himanshu Thapliyal. Tcount and qubit optimized quantum circuit design of the non-restoring square root algorithm. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(3):36:1–36:15, October 2018. CODEN ???? ISSN 1550-4832.

Madhavan:2021:TSM

[MDS21] Advait Madhavan, Matthew W. Daniels, and Mark D. Stiles. Temporal state machines: Using temporal memory to stitch time-based graph computations. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):28:1–28:27, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3451214.

Masadeh:2021:QAA

[MEHT21] Mahmoud Masadeh, Yassmeen Elderhalli, Osman Hasan, and Sofiene Tahar. A quality-assured approximate hardware accelerators-based on machine learning and dynamic partial reconfiguration. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):57:1– 57:19, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3462329.

Mojumder:2013:DPS

[MFA^{+}13] Niladri N. Mojumder, Xuanyao Fong, Charles Augustine, Sumeet K. Gupta, Sri Harsha Choday, and Kaushik Roy. Dual pillar spin-transfer torque MRAMs for low power applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(2):14:1– 14:??, May 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Motaman:2018:IPV

[MGK18] Seyedhamidreza Motaman, Swaroop Ghosh, and Jaydeep Kulkarni. Impact of process variation on self-reference sensing scheme and adaptive current modulation for robust STTRAM sensing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1):8:1–8:??, March 2018. CODEN ???? ISSN 1550-4832.

Mehrabani:2022:NHE

[MGMU22] Yavar Safaei Mehrabani, Samaneh Goldani Gigasari, Mohammad Mirzaei, and Hamidreza Uoosefian. A novel highly-efficient inexact full adder cell for motion and edge detection systems of image processing in CNFET technology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18 (4):74:1–74:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3524061.

Majerus:2012:WUL

[MGS^{+}12] Steve J. A. Majerus, Steven L. Garverick, Michael A. Suster, Paul C. Fletter, and Margot S. Damaser. Wireless, ultra-low-power implantable sensor for chronic bladder pressure monitoring. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(2):11:1– 11:??, June 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mosavirik:2022:SVE

[MGST22] Tahoura Mosavirik, Fatemeh Ganji, Patrick Schaumont, and Shahin Tajik.

ScatterVerif: Verification

of electronic boards using reflection response of power distribution network. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):65:1– 65:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3513087.

Mohammadi:2017:FTR

[MGZ^{+}17] Hassan Ghasemzadeh Mohammadi, Pierre-Emmanuel Gaillardon, Jian Zhang, Giovanni De Micheli, Ernesto Sanchez, and Matteo Sonza Reorda. A fault-tolerant ripple-carry adder with controllablepolarity transistors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):16:1– 16:??, March 2017. CODEN ???? ISSN 1550-4832.

Ma:2008:MCE

[MHL08] Xiaojun Ma, Jing Huang, and Fabrizio Lombardi. A model for computing and energy dissipation of molecular QCA devices and circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(4):3:1–3:??, January 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Miyakawa:2008:MST

[MHM^{+}08] Nobuaki Miyakawa, Eiri Hashimoto, Takanori Maebashi, Natsuo Nakamura, Yutaka Sacho, Shigeto Nakayama, and Shinjiro Toyoda. Multilayer stacking technology using wafer-to-wafer stacked method. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(4):20:1–20:??, October 2008.

CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Misirli:2014:CMM

[MHW14] Goksel Misirli, Jennifer Hallinan, and Anil Wipat. Composable modular models for synthetic biology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3):22:1– 22:??, December 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mittal:2016:SAT

[Mit16] Sparsh Mittal. A survey of architectural techniques for near-threshold computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):46:1– 46:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mittal:2017:STA

[Mit17] Sparsh Mittal. A survey of techniques for architecting processor components using domain-wall memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):29:1– 29:??, March 2017. CODEN ???? ISSN 1550-4832.

Matherat:2011:RCC

[MJ11] Philippe Matherat and Marc-Thierry Jaekel. Relativistic causality and clockless circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(4):20:1–20:??, December 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mondal:2023:TGT

[MKG^{+}23] Anindan Mondal, Debasish Kalita, Archisman Ghosh, Suchismita Roy, and Bibhash Sen. Toward the generation of test vectors for the detection of hardware Trojan targeting effective switching activity. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(4):29:1– 29:??, October 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3597497.

Mahalat:2022:PBS

[MKMS22] Mahabub Hasan Mahalat, Dipankar Karmakar, Anindan Mondal, and Bibhash Sen. PUF based secure and lightweight authentication and key-sharing scheme for wireless sensor network. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):9:1– 9:23, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3466682.

Merkel:2017:SCB

[MKSW17] Cory Merkel, Dhireesha Kudithipudi, Manan Suri, and Bryant Wysocki. Stochastic CBRAM-based neuromorphic time series prediction system. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3): 37:1–37:??, May 2017. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Murray:2014:PEC

[MKW^{+}14] Jacob Murray, Ryan Kim, Paul Wettin, Partha Pratim Pande, and Behrooz Shirazi. Performance evaluation of congestion-aware routing with DVFS on a millimeter-wave small-world wireless NoC. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):17:1– 17:??, November 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Ma:2008:IEF

[MLK^{+}08] Yuchun Ma, Yongxiang Liu, Eren Kursun, Glenn Reinman, and Jason Cong. Investigating the effects of finegrain three-dimensional integration on microarchitecture design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(4):17:1–17:??, October 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mehta:2020:BHE

[MLP^{+}20] Dhwani Mehta, Hangwei Lu, Olivia P. Paradis, Mukhil Azhagan M. S., M. Tanjidur Rahman, Yousef Iskander, Praveen Chawla, Damon L. Woodard, Mark Tehranipoor, and Navid Asadizanjani. The big hack explained: Detection and prevention of PCB supply chain implants. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(4):42:1– 42:25, October 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3401980.

Monta:2023:SCS

[MLW^{+}23] Kazuki Monta, Lang Lin, Jimin Wen, Harsh Shrivastav, Calvin Chow, Hua Chen, Joao Geada, Sreeja Chowdhury, Nitin Pundir, Norman Chang, and Makoto Nagata. Silicon-correlated simulation methodology of EM side-channel leakage analysis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(1):9:1– 9:??, January 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3568957.

Munoz-Martinez:2023:SST

[MMAAK23] Francisco Muñoz-Martínez, José L. Abellán, Manuel E. Acacio, and Tushar Krishna.

STIFT: a spatio-temporal integrated folding tree for efficient reductions in flexible DNN accelerators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(4):32:1– 32:??, October 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3531011.

Mintz:2020:QLE

[MMD^{+}20] Tiffany M. Mintz, Alexander J. McCaskey, Eugene F. Dumitrescu, Shirley V. Moore, Sarah Powers, and Pavel Lougovski. QCOR: a language extension specification for the heterogeneous quantumclassical model of computation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2): 22:1–22:17, April 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3380964.

Mishra:2009:LPF

[MMJ09] Prateek Mishra, Anish Muttreja, and Niraj K. Jha. Lowpower FinFET circuit synthesis using multiple supply and threshold voltages. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(2):7:1–7:??, July 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Massoud:2006:MDC

[MN06] Yehia Massoud and Arthur Nieuwoudt. Modeling and design challenges and solutions for carbon nanotube-based interconnect in future high performance integrated circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(3):155– 196, July 2006. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Machida:2014:JCT

[MNT14] Fumio Machida, Victor F.

Nicola, and Kishor S. Trivedi. Job completion time on a virtualized server with software rejuvenation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):10:1– 10:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mogensen:2014:GFR

[Mog14] Torben Ægidius Mogensen.

Garbage-free reversible multipliers for arbitrary constants. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):12:1– 12:??, November 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mohanty:2012:SSN

[Moh12] Saraju P. Mohanty. Special section on new circuit and architecture-level solutions for multidiscipline systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):14:1– 14:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mohanty:2010:UDS

[MP10] Saraju P. Mohanty and Dhiraj K. Pradhan. ULS: a dual-V_{th} /high-κ nano-CMOS universal level shifter for systemlevel power management. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(2):8:1– 8:??, June 2010. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Maftei:2013:MBS

[MPM13] Elena Maftei, Paul Pop, and Jan Madsen. Module-based synthesis of digital microfluidic biochips with dropletaware operation execution. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(1):2:1– 2:??, February 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mohanty:2022:ISI

[MPR^{+}22] Saraju P. Mohanty, Jim Plusquellic, Garrett S. Rose, Wei Zhang, and Maria K. Michael. Introduction to the special issue on hardwareassisted security for emerging Internet of Things. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):1:1–1:3, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3475952.

Meyer:2021:IBH

[MPZ21] Bruno Henrique Meyer, Aurora Trinidad Ramirez Pozo, and Wagner M. Nunan Zola. Improving Barnes–Hut t-SNE algorithm in modern GPU architectures with random forest KNN and simulated wide-warp. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):53:1–53:26, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3447779.

Maity:2022:CEB

[MRG22] Dilip Kumar Maity, Surajit Kumar Roy, and Chandan Giri. A cost-effective built-in self-test mechanism for post-manufacturing TSV defects in 3D ICs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):70:1– 70:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3517808.

Mahalingam:2012:DCS

[MRH12] Venkataraman Mahalingam, Nagarajan Ranganathan, and Ransford Hyman, Jr. Dynamic clock stretching for variation compensation in VLSI circuit design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):16:1– 16:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Manem:2012:DCM

[MRR12] H. Manem, J. Rajendran, and G. S. Rose. Design considerations for multilevel CMOS/ Nano memristive memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(1):6:1– 6:??, February 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mankalale:2017:OSC

[MS17] Meghna G. Mankalale and Sachin S. Sapatnekar. Optimized standard cells for allspin logic. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):21:1–21:??, March 2017. CODEN ???? ISSN 1550-4832.

Mondal:2019:SST

[MS19] Ankit Mondal and Ankur Srivastava. In situ stochastic training of MTJ crossbars with machine learning algorithms. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(2):16:1–16:??, June 2019. CODEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3309880.

Mondal:2020:EED

[MS20] Ankit Mondal and Ankur Srivastava. Energy-efficient design of MTJ-based neural networks with stochastic computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):7:1–7:27, February 2020. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/doi/abs/10.
1145/3359622.

Mayahinia:2022:VCO

[MSB^{+}22] Mahta Mayahinia, Abhairaj Singh, Christopher Bengel, Stefan Wiefels, Muath A. Lebdeh, Stephan Menzel, Dirk J. Wouters, Anteneh Gebregiorgis, Rajendra Bishnoi, Rajiv Joshi, and Said Hamdioui. A voltage-controlled, oscillation-based ADC design for computation-in-memory architectures using emerging ReRAMs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):32:1–32:25, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3451212.

Mukhopadhyay:2021:PES

[MSC^{+}21] Anand Kumar Mukhopadhyay, Atul Sharma, Indrajit Chakrabarti, Arindam Basu, and Mrigank Sharad. Powerefficient spike sorting scheme using analog spiking neural network classifier. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):12:1– 12:29, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3432814.

Manna:2019:TAT

[MSCS19] Kanchan Manna, Chatla Swami Sagar, Santanu Chattopadhyay, and Indranil Sengupta. Thermal-aware test scheduling strategy for network-on-chip based systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):6:1–6:??, February 2019. CODEN ???? ISSN 1550-4832.

Myers:2014:ISI

[MSW14] Chris J. Myers, Herbert Sauro, and Anil Wipat. Introduction to the special issue on computational synthetic biology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3): 20:1–20:??, December 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Mohanty:2014:SOS

[MT14] Pragyan (Sheela) Mohanty and Spyros Tragoudas. Scalable offline searches in DNA sequences. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):18:1–18:??, November 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Metodi:2008:HLI

[MTC^{+}08] Tzvetan S. Metodi, Darshan D. Thaker, Andrew W. Cross, Isaac L. Chuang, and Frederic T. Chong. High-level interconnect model for the quantum logic array architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(1): 1:1–1:??, March 2008. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Madsen:2014:SMC

[MZR^{+}14] Curtis Madsen, Zhen Zhang, Nicholas Roehner, Chris Winstead, and Chris Myers. Stochastic model checking of genetic circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3):23:1–23:??, December 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Ma:2023:MMT

[MZZ23] Yu Ma, Linfeng Zheng, and Pingqiang Zhou. A mapping method tolerating SAF and variation for memristor crossbar array based neural network inference on edge devices. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(2):15:1–15:??, April 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3585518.

Narendra:2005:CDC

[Nar05] Siva G. Narendra. Challenges and design choices in nanoscale CMOS. ACM Journal on Emerging Technologies in Computing Systems (JETC), 1(1):7–49, April 2005. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Narayanan:2008:E

[Nar08] Vijaykrishnan Narayanan. Editorial. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(2):4:1–4:??, April 2008. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Nowshin:2024:TEE

[NAY24] Fabiha Nowshin, Hongyu An, and Yang Yi. Towards energy-efficient spiking neural networks: a robust hybrid CMOS-memristive accelerator. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(1):

5:1–5:??, January 2024. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3635165.

Ni:2017:DMC

[NHL^{+}17] Leibin Ni, Hantao Huang, Zichuan Liu, Rajiv V. Joshi, and Hao Yu. Distributed in-memory computing on binary RRAM crossbar. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):36:1– 36:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Narayanan:2013:VNF

[NLK^{+}13] Pritish Narayanan, Michael Leuchtenburg, Jorge Kina, Prachi Joshi, Pavan Panchapakeshan, Chi On Chui, and C. Andras Moritz. Variability in nanoscale fabrics: Bottomup integrated analysis and mitigation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(1):8:1–8:??, February 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Najafi:2017:RAS

[NLL^{+}17] M. Hassan Najafi, Peng Li, David J. Lilja, Weikang Qian, Kia Bazargan, and Marc Riedel. A reconfigurable architecture with sequential logic-based stochastic computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):57:1–57:??, August 2017. CODEN ???? ISSN 1550-4832.

Ni:2020:WBE

[NLW^{+}20] Jiacheng Ni, Keren Liu, Bi Wu, Weisheng Zhao, Yuanqing Cheng, Xiaolong Zhang, and Ying Wang. Write back energy optimization for STT-MRAM-based last-level cache with data pattern characterization. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3):29:1–29:18, July 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3381860.

Naruse:2012:SDN

[NPA^{+}12] Makoto Naruse, Ferdinand Peper, Kouichi Akahane, Naokatsu Yamamoto, Tadashi Kawazoe, Naoya Tate, and Motoichi Ohtsu. Skew dependence of nanophotonic devices based on optical nearfield interactions. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(1):4:1–4:??, February 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Neugebauer:2018:FQM

[NPH18] Florian Neugebauer, Ilia Polian, and John P. Hayes.

Framework for quantifying and managing accuracy in stochastic circuit design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):31:1– 31:??, July 2018. CODEN ???? ISSN 1550-4832.

Nishio:2020:ESI

[NPS^{+}20] Shin Nishio, Yulu Pan, Takahiko Satoh, Hideharu Amano, and Rodney Van Meter. Extracting success from IBM’s 20-qubit machines using error-aware compilation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3): 32:1–32:25, July 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386162.

Nguyen:2014:RED

[NV14] Trung Duc Nguyen and Rodney Van Meter. A resourceefficient design for a reversible floating point adder in quantum computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):13:1– 13:??, November 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Nooraiepour:2022:TVM

[NVW^{+}22] Alireza Nooraiepour, Shaghayegh Vosoughitabar, Chung-Tse Michael Wu, Waheed U. Bajwa, and Narayan B. Mandayam. Time-varying metamaterialenabled directional modulation schemes for physical layer security in wireless communication links. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):64:1–64:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3513088.

Nguyen:2020:CMC

[NYL^{+}20] Hoang Anh Du Nguyen, Jintao Yu, Muath Abu Lebdeh, Mottaqiallah Taouil, Said Hamdioui, and Francky Catthoor. A classification of memory-centric computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2): 13:1–13:26, April 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3365837.

Oberortner:2014:RBD

[OBLD14] Ernst Oberortner, Swapnil Bhatia, Erik Lindgren, and Douglas Densmore. A rulebased design specification language for synthetic biology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3):25:1– 25:??, December 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Oneal:2018:RCS

[OGB18] Kenneth O’neal, Daniel Grissom, and Philip Brisk. Resource-constrained scheduling for digital microfluidic biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1):7:1–7:??, March 2018. CODEN ???? ISSN 1550-4832.

Olney:2022:DNT

[OK22] Brooks Olney and Robert Karam. Diverse, neural Trojan resilient ecosystem of neural network IP. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):57:1–57:23, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3471189.

Olorode:2015:IPS

[ON15] Oluleye Olorode and Mehrdad Nourani. Improving performance in sub-block caches with optimized replacement policies. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4):41:1–41:??, April 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Okafor:2024:FSN

[ORC^{+}24] Ikenna Okafor, Akshay Krishna Ramanathan, Nagadastagiri Reddy Challapalle, Zheyu Li, and Vijaykrishnan Narayanan. Fusing in-storage and near-storage acceleration of convolutional neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(1):1:1–1:??, January 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3597496.

Ottavi:2006:HHE

[OSLT06] Marco Ottavi, Luca Schiano, Fabrizio Lombardi, and Douglas Tougaw. HDLQ: A HDL environment for QCA design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(4): 243–261, October 2006. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Ouyang:2024:FBL

[OYLL24] Yiming Ouyang, Shuaijie Yuan, Jianhua Li, and Huaguo Liang. F-Bypass: a low-power network-on-chip design utilizing bypass to improve network connectivity. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(4):14:1– 14:??, October 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3695874.

Peter:2017:OON

[PAB^{+}17] Eldhose Peter, Anuj Arora, Janibul Bashir, Akriti Bagaria, and Smruti R. Sarangi. Optical overlay NUCA: a highspeed substrate for shared L2 caches. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):53:1–53:??, August 2017. CODEN ???? ISSN 1550-4832.

Pundir:2022:ASV

[PAC^{+}22] Nitin Pundir, Sohrab Aftabjahani, Rosario Cammarota, Mark Tehranipoor, and Farimah Farahmandi. Analyzing security vulnerabilities induced by high-level synthesis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):47:1– 47:22, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3492345.

Parveen:2018:IEE

[PAF18] Farhana Parveen, Shaahin Angizi, and Deliang Fan. IM-FlexCom: Energy efficient in-memory flexible computing using dual-mode SOT-MRAM. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(3):35:1–35:??, October 2018. CODEN ???? ISSN 1550-4832.

Peng:2022:DNN

[PAP^{+}22] Jiaxin Peng, Yousra Alkabani, Krunal Puri, Xiaoxuan Ma, Volker Sorger, and Tarek El-Ghazawi. A deep neural network accelerator using residue arithmetic in a hybrid optoelectronic system. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):81:1– 81:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550273.

Paul:2021:SSI

[PB21] Shubhra Deb Paul and Swarup Bhunia. SILVerIn: Systematic integrity verification of printed circuit board using JTAG infrastructure. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):44:1– 44:28, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3460232.

Plana:2011:SDI

[PCD^{+}11] Luis A. Plana, David Clark, Simon Davidson, Steve Furber, Jim Garside, Eustace Painkras, Jeffrey Pepper, Steve Temple, and John Bainbridge. SpiN-Naker: Design and implementation of a GALS multicore System-on-Chip. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(4):17:1–17:??, December 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Patwardhan:2007:SOD

[PDL07] Jaidev Patwardhan, Chris Dwyer, and Alvin R. Lebeck. A self-organizing defect tolerant SIMD architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(2):10:1– 10:??, July 2007. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Pang:2015:MLN

[PDL15] Jun Pang, Christopher Dwyer, and Alvin R. Lebeck. mNoC: Large nanophotonic networkon-chip crossbars with molecular scale devices. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(1):1:1–1:??, July 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Patwardhan:2006:NNS

[PDLS06] Jaidev P. Patwardhan, Chris Dwyer, Alvin R. Lebeck, and Daniel J. Sorin. NANA: a nano-scale active network architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(1):1–30, January 2006. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Paul:2007:PBC

[PFOL07] Bipul C. Paul, Shinobu Fujita, Masaki Okajima, and Thomas Lee. Prospect of ballistic CNFET in high performance applications: Modeling and analysis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(3):12:1–12:??, November 2007. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Pajouhi:2017:YAE

[PFRR17] Zoha Pajouhi, Xuanyao Fong, Anand Raghunathan, and Kaushik Roy. Yield, area, and energy optimization in STT–MRAMs using failureaware ECC. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):20:1–20:??, March 2017. CODEN ???? ISSN 1550-4832.

Pande:2012:ISI

[PG12] Partha Pratim Pande and Amlan Ganguly. Introduction to the special issue on sustainable and green computing systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(4):26:1–26:??, October 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Park:2015:MME

[PHS^{+}15] Kyu Ho Park, Woomin Hwang, Hyunchul Seok, Chulmin Kim, Dong jae Shin, Dong Jin Kim, Min Kyu Maeng, and Seong Min Kim. MN-MATE: Elastic resource management of manycores and a hybrid memory hierarchy for a cloud node. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(1):5:1–5:??, July 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Patooghy:2023:SNC

[PHS^{+}23] Ahmad Patooghy, Mahdi Hasanzadeh, Amin Sarihi, Mostafa Abdelrehim, and Abdel-Hameed A. Badawy. Securing network-on-chips against fault-injection and crypto-analysis attacks via stochastic anonymous routing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(3): 22:1–22:??, July 2023. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592798.

Page:2017:SHA

[PJSM17] Adam Page, Ali Jafari, Colin Shea, and Tinoosh Mohsenin. SPARCNet: a hardware accelerator for efficient deployment of sparse convolutional networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):31:1–31:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Peper:2013:BCF

[PLC^{+}13] Ferdinand Peper, Jia Lee, Josep Carmona, Jordi Cortadella, and Kenichi Morita. Brownian circuits: Fundamentals. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(1):3:1–3:??, February 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Pasandi:2021:DBG

[PP21] Ghasem Pasandi and Massoud Pedram. Depth-bounded graph partitioning algorithm and dual clocking method for realization of superconducting SFQ circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(1):2:1–2:22, January 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3412389.

Piovesan:2013:ERP

[PPM^{+}13] Damiano Piovesan, Giuseppe Profiti, Pier Luigi Martelli, Piero Fariselli, and Rita Casadio. Extended and robust protein sequence annotation over conservative nonhierarchical clusters: The case study of the ABC transporters. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(4):27:1– 27:??, November 2013. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Paul:2013:ISI

[PR13] Bipul C. Paul and Arijit Raychowdhury. Introduction to the special issue on memory technologies. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(2):10:1–10:??, May 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Patnaik:2015:PPC

[PRG^{+}15] Milan Patnaik, Chidhambaranathan R., Chirag Garg, Arnab Roy, V. R. Devanathan, Shankar Balachandran, and V. Kamakoti. ProWATCh: a proactive cross-layer workload-aware temperature management framework for low-power chip multi-processors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3):22:1–22:??, September 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Park:2020:LSC

[PRV^{+}20] Jungmin Park, Fahim Rahman, Apostol Vassilev, Domenic Forte, and Mark Tehranipoor. Leveraging sidechannel information for disassembly and security. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):6:1–6:21, February 2020. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/doi/abs/10.
1145/3359621.

Pilly:2019:HSL

[PSL^{+}19] Praveen K. Pilly, Nigel D. Stepp, Yannis Liapis, David W. Payton, and Narayan Srinivasa. Hypercolumn sparsification for low-power convolutional neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(2): 20:1–20:??, June 2019. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3304104.

Prasad:2006:DSA

[PSM^{+}06] Aditya K. Prasad, Vivek V. Shende, Igor L. Markov, John P. Hayes, and Ketan N. Patel. Data structures and algorithms for simplifying reversible circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(4):277–293, October 2006. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Panda:2017:EEI

[PSR17] Priyadarshini Panda, Abhronil Sengupta, and Kaushik Roy. Energy-efficient and improved image recognition with conditional deep learning. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):33:1– 33:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Potok:2018:SCD

[PSY^{+}18] Thomas E. Potok, Catherine Schuman, Steven Young, Robert Patton, Federico Spedalieri, Jeremy Liu, Ke-Thia Yao, Garrett Rose, and Gangotree Chakma. A study of complex deep learning networks on high-performance, neuromorphic, and quantum computers. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):19:1–19:??, July 2018. CODEN ???? ISSN 1550-4832.

Palaniswamy:2012:EHI

[PT12] Ashok Kumar Palaniswamy and Spyros Tragoudas. An efficient heuristic to identify threshold logic functions. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):19:1– 19:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Palaniswamy:2014:ITL

[PT14a] Ashok Kumar Palaniswamy and Spyros Tragoudas. Improved threshold logic synthesis using implicant-implicit algorithms. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(3):21:1–21:??, April 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Pierce:2014:NTN

[PT14b] Luke Pierce and Spyros Tragoudas. Nanopipelined threshold network synthesis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(2):17:1– 17:??, February 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Prodan:2007:DDE

[PUBV07] Lucian Prodan, Mihai Udrescu, Oana Boncalo, and Mircea Vladutiu. Design for dependability in emerging technologies. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(2): 6:1–6:??, July 2007. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Panoff:2022:RCA

[PYSJ22] Max Panoff, Honggang Yu, Haoqi Shan, and Yier Jin. A review and comparison of AIenhanced side channel analysis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3): 62:1–62:20, July 2022. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3517810.

Quadir:2016:SCS

[QCF^{+}16] Shahed E. Quadir, Junlin Chen, Domenic Forte, Navid Asadizanjani, Sina Shahbazmohamadi, Lei Wang, John Chandy, and Mark Tehranipoor. A survey on chip to system reverse engineering. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1): 6:1–6:??, December 2016. CO-DEN ???? ISSN 1550-4832.

Qian:2020:MBC

[QGW20] Fengyu Qian, Yanping Gong, and Lei Wang. A memristorbased compressive sampling encoder with dynamic rate control for low-power video streaming. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2):12:1–12:16, April 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3365836.

Rakka:2024:HSH

[RAC^{+}24] Mariam Rakka, Walaa Amer, Hanning Chen, Mohsen Imani, and Fadi Kurdahi. HDRLPIM: a simulator for hyper-dimensional reinforcement learning based on processing in-memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(4):15:1– 15:??, October 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3695875.

Rueckauer:2022:NAC

[RBG^{+}22] Bodo Rueckauer, Connor Bybee, Ralf Goettsche, Yashwardhan Singh, Joyesh Mishra, and Andreas Wild. NxTF: an API and compiler for deep spiking neural networks on Intel Loihi. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):48:1–48:22, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3501770.

Roy:2014:TAG

[RBGC14] Sudip Roy, Bhargab B. Bhattacharya, Sarmishtha Ghoshal, and Krishnendu Chakrabarty. Theory and analysis of generalized mixing and dilution of biochemical fluids using digital microfluidic biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(1):2:1–2:??, September 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Robissout:2021:IDL

[RBHG21] Damien Robissout, Lilian Bossuet, Amaury Habrard, and Vincent Grosso. Improving deep learning networks for profiled side-channel analysis using performance improvement techniques. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):41:1– 41:30, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3453162.

Ruffini:2022:NFB

[RCYB22] Simone Ruffini, Luca Caronti, Kasim Sinan Yildirim, and Davide Brunelli. NORM: an FPGA-based non-volatile memory emulation framework for intermittent computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):73:1– 73:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3517812.

Roessler:2022:SEL

[RD22] Nick Roessler and André De-Hon. SCALPEL: Exploring the limits of tag-enforced compartmentalization. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):2:1–2:28, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3461673.

Rahman:2014:AQT

[RDH14] Md. Mazder Rahman, Gerhard W. Dueck, and Joseph D. Horton. An algorithm for quantum template matching. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3):31:1– 31:??, December 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Rahman:2021:CGO

[RDM^{+}21] M. Tanjidur Rahman, Nusrat Farzana Dipu, Dhwani Mehta, Shahin Tajik, Mark Tehranipoor, and Navid Asadizanjani. CONCEALING-Gate: Optical contactless probing resilient design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):39:1– 39:25, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3446998.

Roy:2024:SLA

[RDS24] Sourav Roy, Dipnarayan Das, and Bibhash Sen. Secure and lightweight authentication protocol using PUF for the IoT-based wireless sensor network. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(1):3:1–3:??, January 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3624477.

Ronen:2022:BMP

[REL^{+}22] Ronny Ronen, Adi Eliahu, Orian Leitersdorf, Natan Peled, Kunal Korgaonkar, Anupam Chattopadhyay, Ben Perach, and Shahar Kvatinsky. The Bitlet model: a parameterized analytical model to compare PIM and CPU systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):43:1–43:29, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3465371.

Reza:2023:MLE

[Rez23] Md Farhadur Reza. Machine learning enabled solutions for design and optimization challenges in networks-on-chip based multi/many-core architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(3):23:1–23:??, July 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3591470.

Romani:2015:SSC

[RFDT15] Aldo Romani, Matteo Filippi, Michele Dini, and Marco Tartagni. A sub-μ a stand-by current synchronous electric charge extractor for piezoelectric energy harvesting. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(1):7:1–7:??, July 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Rasheed:2020:CAP

[RHB^{+}20] Farhan Rasheed, Michael Hefenbrock, Rajendra Bishnoi, Michael Beigl, Jasmin Aghassi-Hagmann, and Mehdi B. Tahoori. Crossoveraware placement and routing for inkjet printed circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2):19:1– 19:22, April 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3375461.

Rahman:2015:NVR

[RKM15] Mostafizur Rahman, Santosh Khasanvis, and Csaba Andras Moritz. Nanowire volatile RAM as an alternative to SRAM. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3):30:1–30:??, September 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Roy:2012:CAL

[RMBC12] Sudip Roy, Debasis Mitra, Bhargab B. Bhattacharya, and Krishnendu Chakrabarty.

Congestion-aware layout design for high-throughput digital microfluidic biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):17:1– 17:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Rodriguez:2015:TSS

[RMG15] Laurent Rodriguez, Benoît Miramond, and Bertrand Granado. Toward a sparse self-organizing map for neuromorphic architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(4):33:1– 33:??, April 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Rakshit:2017:MTS

[RMW^{+}17] Joydeep Rakshit, Kartik Mohanram, Runlai Wan, Kai Tak Lam, and Jing Guo. Monolayer transistor SRAMs: Toward low-power, denser memory systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):18:1–18:??, March 2017. CODEN ???? ISSN 1550-4832.

Riaz:2022:SAP

[RNN^{+}22] Arslan Riaz, Dylan Nash, Jonathan Ngo, Chiraag Juvekar, Phillip Nadeau, Tao Yu, and Rabia Tugce Yazicigil. Security assessment of phase-based ranging systems in a multipath environment. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):66:1– 66:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3517809.

Rioja:2021:USC

[RPBA21] Unai Rioja, Servio Paguada, Lejla Batina, and Igor Armendariz. The uncertainty of side-channel analysis: a way to leverage from heuristics. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):40:1– 40:27, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3446997.

Ray:2023:SCI

[RSBA23] Dipojjwal Ray, Yogendra Sao, Santosh Biswas, and Sk Subidh Ali. On securing cryptographic ICs against scan-based attacks: a Hamming weight distribution perspective. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(2):10:1–10:??, April 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3577215.

Ram:2023:ETA

[RSD^{+}23a] Saswat Kumar Ram, Sauvagya Ranjan Sahoo, Banee Bandana Das, Kamalakanta Mahapatra, and Saraju P. Mohanty. Eternal-thing 2.0: Analog-Trojan-resilient ripple-less solar harvesting system for sustainable IoT. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(2):12:1–12:??, April 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3575800.

Rout:2023:MSE

[RSD23b] Sidhartha Sankar Rout, Mitali Sinha, and Sujay Deb. 2DMAC: a sustainable and efficient medium access control mechanism for future wireless NoCs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(3):20:1–20:??, July 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3570727.

Rad:2007:EAP

[RT07] Reza M. P. Rad and Mohammad Tehranipoor. Evaluating area and performance of hybrid FPGAs with nanoscale clusters and CMOS routing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(3):15:1– 15:??, November 2007. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Rad:2008:SNA

[RT08] Reza Rad and Mohammad Tehranipoor. SCT: a novel approach for testing and configuring nanoscale devices. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(3):14:1– 14:??, August 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Rose:2007:DCM

[RYT^{+}07] Garrett S. Rose, Yuxing Yao, James M. Tour, Adam C. Cabe, Nadine Gergel-Hackett, Nabanita Majumdar, John C. Bean, Lloyd R. Harriott, and Mircea R. Stan. Designing CMOS/molecular memories while considering device parameter variations. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(1):3:1–3:??, April 2007. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Sayyaparaju:2020:DAC

[SAAR20] Sagarvarma Sayyaparaju, Md Musabbir Adnan, Sherif Amer, and Garrett S. Rose. Deviceaware circuit design for robust memristive neuromorphic systems with STDP-based learning. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3): 28:1–28:25, July 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3380969.

Schmuck:2019:HOD

[SBR19] Manuel Schmuck, Luca Benini, and Abbas Rahimi. Hardware optimizations of dense binary hyperdimensional computing: Rematerialization of hypervectors, binarized bundling, and combinational associative memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):32:1–32:??, December 2019. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3314326.

Sayed:2020:ASM

[SBT20] Nour Sayed, Rajendra Bishnoi, and Mehdi B. Tahoori. Approximate spintronic memories. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(4):43:1–43:22, October 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3404980.

Samavatian:2020:RAR

[SBZT20] Mohammad Hossein Samavatian, Anys Bacha, Li Zhou, and Radu Teodorescu. RN-NFast: an accelerator for recurrent neural networks using domain-wall memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(4):38:1– 38:27, October 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3399670.

Su:2006:YER

[SC06] Fei Su and Krishnendu Chakrabarty. Yield enhancement of reconfigurable microfluidics-based biochips using interstitial redundancy. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(2):104– 128, April 2006. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Su:2008:HLS

[SC08] Fei Su and Krishnendu Chakrabarty. High-level synthesis of digital microfluidic biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(4):1:1–1:??, January 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Simsir:2009:HNC

[SCI^{+}09] Muzaffer O. Simsir, Srihari Cadambi, Franjo Ivanvčić, Martin Roetteler, and Niraj K. Jha. A hybrid nano-CMOS architecture for defect and fault tolerance. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(3):14:1– 14:??, August 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Samal:2017:IPU

[SCL17] Sandeep Kumar Samal, Guoqing Chen, and Sung Kyu Lim. Improving performance under process and voltage variations in near-threshold computing using 3D ICs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):59:1– 59:??, August 2017. CODEN ???? ISSN 1550-4832.

Seo:2019:GEIb

[SCLW19a] Jae-Sun Seo, Yu Cao, Xin Li, and Paul Whatmough. Guest editors’ introduction: Hardware and algorithms for energy-constrained on-chip machine learning (part 2). ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):31:1– 31:??, December 2019. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3359336.

Seo:2019:GEIa

[SCLW19b] Jae-Sun Seo, Yu Cao, Xin Li, and Paul Whatmough. Guest Editors’ introduction to the special section on hardware and algorithms for energyconstrained on-chip machine learning. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(2):14:1–14:??, June 2019. CODEN ???? ISSN 1550-

                                        4832. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3322433.

Simoni:2022:TCM

[SCT^{+}22] Mario Simoni, Giovanni Amedeo Cirillo, Giovanna Turvani, Mariagrazia Graziano, and Maurizio Zamboni. Towards compact modeling of noisy quantum computers: a molecular-spin-qubit case of study. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):11:1–11:26, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3474223.

Sun:2012:NMD

[SCZ^{+}12] Zhenyu Sun, Xiang Chen, Yaojun Zhang, Hai Li, and Yiran Chen. Nonvolatile memories as the data storage system for implantable ECG recorder. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(2):13:1–13:??, June 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Sen:2014:RRC

[SDSS14] Bibhash Sen, Manojit Dutta, Samik Some, and Biplab K. Sikdar. Realizing reversible computing in QCA framework resulting in efficient design of testable ALU. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(3):30:1–30:??, December 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Sekanina:2007:EFR

[Sek07] Lukáš Sekanina. Evolutionary functional recovery in virtual reconfigurable circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(2):8:1–8:??, July 2007. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Salehi:2017:SSM

[SFD17] Soheil Salehi, Deliang Fan, and Ronald F. Demara. Survey of STT–MRAM cell design strategies: Taxonomy and sense amplifier tradeoffs for resiliency. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):48:1–48:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Sharad:2012:LPA

[SGR^{+}12] Mrigank Sharad, Sumeet K.

Gupta, Shriram Raghunathan, Pedro P. Irazoqui, and Kaushik Roy. Lowpower architecture for epileptic seizure detection based on reduced complexity DWT. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(2):10:1– 10:??, June 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Slijepcevic:2019:TRW

[SHAC19] Mladen Slijepcevic, Carles Hernandez, Jaume Abella, and Francisco J. Cazorla. Time-randomized wormhole NoCs for critical applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):3:1– 3:??, February 2019. CODEN ???? ISSN 1550-4832.

Song:2021:DRM

[SHB^{+}21] Shihao Song, Jui Hanamshet, Adarsha Balaji, Anup Das, Jeffrey L. Krichmar, Nikil D. Dutt, Nagarajan Kandasamy, and Francky Catthoor. Dynamic reliability management in neuromorphic computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):63:1– 63:27, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3462330.

Shukla:2009:GEI

[Shu09] Sandeep Shukla. Guest editorial: IEEE/ACM symposium on nanoscale architectures (NANOARCH07). ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(1):1:1–1:??, January 2009. CODEN ????

ISSN 1550-4832 (print), 1550-4840 (electronic).

Sakalis:2020:EPA

[SJKS20] Christos Sakalis, Alexandra Jimborean, Stefanos Kaxiras, and Magnus Själander. Evaluating the potential applications of quaternary logic for approximate computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):5:1– 5:25, February 2020. CO-DEN ???? ISSN 1550-4832.

URL https://dl.acm.org/
doi/abs/10.1145/3359620.

Sinanoglu:2016:GES

[SK16] Ozgur Sinanoglu and Ramesh Karri. Guest editorial special issue on secure and trustworthy computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1):1:1–1:??, December 2016. CODEN ???? ISSN 1550-4832.

Srinivasan:2013:NAF

[SKB13] S. Srinivasan, V. Kamakoti, and A. Bhattacharya. A novel algorithm for fast synthesis of DNA probes on microarrays. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(1): 1:1–1:??, February 2013. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Sun:2013:EVC

[SKRX13] Guangyu Sun, Eren Kursun, Jude A. Rivers, and Yuan Xie. Exploring the vulnerability of CMPs to soft errors with 3D stacked nonvolatile memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(3): 22:1–22:??, September 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Sartor:2017:EIH

[SLC^{+}17] Anderson L. Sartor, Arthur F. Lorenzon, Luigi Carro, Fernanda Kastensmidt, Stephan Wong, and Antonio C. S. Beck. Exploiting idle hardware to provide low overhead fault tolerance for VLIW processors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):13:1–13:??, March 2017. CODEN ???? ISSN 1550-4832.

Shi:2022:GEA

[SLCJ22] Yiyu Shi, Yongpan Liu, Jianxu Chen, and Steve Jiang. Guest editorial: ACM JETC special issue on hardwareaware learning for medical applications. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):24:1–24:3, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3503262.

Sun:2014:WAC

[SLS^{+}14] Jin Sun, Roman Lysecky, Karthik Shankar, Avinash Kodi, Ahmed Louri, and Janet Roveda. Workload assignment considering NBTI degradation in multicore systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):4:1–4:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Sheikh:2011:EEP

[SM11] Basit Riaz Sheikh and Rajit Manohar. Energy-Efficient pipeline templates for High-Performance asynchronous circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(4):19:1–19:??, December 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Shea:2019:HSD

[SM19] Colin Shea and Tinoosh Mohsenin. Heterogeneous scheduling of deep neural networks for low-power realtime designs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):36:1–36:??, December 2019. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3358699.

Sego:2012:IDC

[SMR^{+}12] Landon H. Sego, Andrés Márquez, Andrew Rawson, Tahir Cader, Kevin Fox, William I. Gustafson, Jr., and Christopher J. Mundy. Implementing the data center energy productivity metric. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(4):30:1– 30:??, October 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Sisejkovic:2021:CSL

[SMR^{+}21] Dominik Sisejkovic, Farhad Merchant, Lennart M. Reimann, Harshit Srivastava, Ahmed Hallawa, and Rainer Leupers. Challenging the security of logic locking schemes in the era of deep learning: a neuroevolutionary approach. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):30:1– 30:26, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3431389.

Samant:2023:NAL

[SMR23] Prerana Samant, Naveen Kumar Macha, and Mostafizur Rahman. A neoteric approach for logic with embedded memory leveraging crosstalk computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(1):2:1–2:??, January 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3569917.

Singhvi:2017:FGU

[SMT^{+}17] Ajay Singhvi, Matheus T.

Moreira, Ramy N. Tadros, Ney L. V. Calazans, and Peter A. Beerel. A fine-grain, uniform, energy-efficient delay element for 2-phase bundleddata circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):15:1–15:??, March 2017. CODEN ???? ISSN 1550-4832.

Sayed:2021:DBP

[SMT21] Nour Sayed, Longfei Mao, and Mehdi B. Tahoori. Dynamic behavior predictions for fast and efficient hybrid STT-MRAM caches. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(1):9:1–9:21, January 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3423135.

Saeed:2019:ISA

[SMZ^{+}19] Samah Mohamed Saeed, Nithin Mahendran, Alwin Zulehner, Robert Wille, and Ramesh Karri. Identification of synthesis approaches for IP/IC piracy of reversible circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(3):23:1–23:??, June 2019. CODEN ???? ISSN 1550-

                                        4832. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3289392.

Singh:2010:CPD

[SN10] Montek Singh and Steven M. Nowick. Call for papers: Deadline: March 15, 2011. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(4):15:1– 15:??, December 2010. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Singh:2011:ISI

[SN11] Montek Singh and Steven M. Nowick. Introduction to special issue: Asynchrony in system design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(4):14:1–14:??, December 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Srinivasan:2018:SBU

[SPR18] Gopalakrishnan Srinivasan, Priyadarshini Panda, and Kaushik Roy. STDP-based unsupervised feature learning using convolution-over-time in spiking neural networks for energy-efficient neuromorphic computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4):44:1–44:??, December 2018. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3266229.

Shafi:2024:RUD

[SPS^{+}24] Omais Shafi, Mohammad Khalid Pandit, Amarjeet Saini, Gayathri Ananthanarayanan, and Rijurekha Sen. Repercussions of using DNN compilers on edge GPUs for real time and safety critical systems: a quantitative audit. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(1):4:1– 4:??, January 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3611016.

Savage:2006:RAN

[SRD^{+}06] John E. Savage, Eric Rachlin, André DeHon, Charles M. Lieber, and Yue Wu. Radial addressing of nanowires. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(2):129–154, April 2006. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Shi:2015:ISI

[SS15] Yiyu Shi and Takashi Sato.

Introduction to: Special issue on cross-layer system design. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3): 20:1–20:??, September 2015.

CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Sitik:2015:FBL

[SSF^{+}15] Can Sitik, Emre Salman, Leo Filippini, Sung Jun Yoon, and Baris Taskin. FinFET-based low-swing clocking. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(2):13:1– 13:??, August 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Salam:2012:ICL

[SSN12] Muhammad Tariqus Salam, Mohamad Sawan, and Dang Khoa Nguyen. Implantable closedloop epilepsy prosthesis: Modeling, implementation and validation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(2):9:1–9:??, June 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Shafaei:2014:CSR

[SSP14] Alireza Shafaei, Mehdi Saeedi, and Massoud Pedram. Cofactor sharing for reversible logic synthesis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2):14:1–14:??, November 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Smith:2020:HDQ

[ST20] Kaitlin N. Smith and Mitchell A.

Thornton. Higher dimension quantum entanglement generators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):3:1–3:21, February 2020. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/doi/abs/10.
1145/3345501.

Sheikh:2012:EPA

[STA^{+}12] Hafiz Fahad Sheikh, Hengxing Tan, Ishfaq Ahmad, Sanjay Ranka, and Phanisekhar Bv. Energy- and performanceaware scheduling of tasks on parallel and distributed systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(4):32:1–32:??, October 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Sunny:2021:SSP

[STNP21] Febin P. Sunny, Ebadollah Taheri, Mahdi Nikdast, and Sudeep Pasricha. A survey on silicon photonics for deep learning. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):61:1–61:57, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3459009.

Senni:2017:NVP

[STSG17] Sophiane Senni, Lionel Torres, Gilles Sassatelli, and Abdoulaye Gamatie. Nonvolatile processor based on MRAM for ultra-low-power IoT devices. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):17:1–17:??, March 2017. CODEN ???? ISSN 1550-4832.

Sarwar:2018:EEN

[SVA^{+}18] Syed Shakib Sarwar, Swagath Venkataramani, Aayush Ankit, Anand Raghunathan, and Kaushik Roy. Energyefficient neural computing with approximate multipliers. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2): 16:1–16:??, July 2018. CO-DEN ???? ISSN 1550-4832.

Schulhof:2007:SRC

[SWJ07] Gabriel Schulhof, Konrad Walus, and Graham A. Jullien. Simulation of random cell displacements in QCA. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(1):2:1– 2:??, April 2007. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Soeken:2016:ELB

[SWK^{+}16] Mathias Soeken, Robert Wille, Oliver Keszocze, D. Michael Miller, and Rolf Drechsler. Embedding of large Boolean functions for reversible logic. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):41:1– 41:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Shang:2011:INC

[SX11] Li Shang and Qianfan Xu.

Introduction to nanophotonic communication technology integration. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(2):5:1–5:??, June 2011. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Srivastava:2012:CLV

[SXL^{+}12] Ashok Srivastava, Yao Xu, Yang Liu, Ashwani K. Sharma, and Clay Mayberry. CMOS LC voltage controlled oscillator design using multiwalled and single-walled carbon nanotube wire inductors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):15:1– 15:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Saeedi:2010:RCS

[SZSS10] Mehdi Saeedi, Morteza Saheb Zamani, Mehdi Sedighi, and Zahra Sasanian. Reversible circuit synthesis using a cyclebased approach. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(4):13:1–13:??, December 2010. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Tahoori:2006:AID

[Tah06] Mehdi B. Tahoori. Applicationindependent defect tolerance of reconfigurable nanoarchitectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(3):197–218, July 2006. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Tahoori:2009:LOD

[Tah09] Mehdi B. Tahoori. Lowoverhead defect tolerance in crossbar nanoarchitectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(2):11:1– 11:??, July 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Taskin:2009:SRB

[TCSV09] Baris Taskin, Andy Chiu, Jonathan Salkind, and Daniel Venutolo. A shift-register-based QCA memory architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(1): 4:1–4:??, January 2009. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Talukder:2023:NTD

[TFR23] B. M. S. Bahar Talukder, Farah Ferdaus, and Md Tauhidur Rahman. A noninvasive technique to detect authentic/counterfeit SRAM chips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(2):11:1– 11:??, April 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3597024.

Thakker:2021:CRK

[TFZ^{+}21] Urmish Thakker, Igor Fedorov, Chu Zhou, Dibakar Gope, Matthew Mattina, Ganesh Dasika, and Jesse Beu. Compressing RNNs to kilobyte budget for IoT devices using Kronecker products. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17 (4):46:1–46:18, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3440016.

Tyrrell:2007:ED

[TG07] Andy M. Tyrrell and Andrew J. Greensted. Evolving dependability. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(2):7:1–7:??, July 2007. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Tang:2016:DPM

[TGCJ16] Aoxiang Tang, Xun Gao, Lung-Yen Chen, and Niraj K. Jha. Delay/power modeling and optimization of Fin-FET circuit modules under PVT variations: Observing the trends between the 22nm and 14nm technology nodes. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):42:1– 42:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Tang:2013:DSE

[TJ13a] Aoxiang Tang and Niraj K. Jha. Design space exploration of FinFET cache. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(3):20:1– 20:??, September 2013. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Tang:2013:TCT

[TJ13b] Aoxiang Tang and Niraj K. Jha. Thermal characterization of test techniques for FinFET and 3D integrated circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(1):6:1–6:??, February 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Tolbert:2012:MDA

[TKBM12] Jeremy R. Tolbert, Pratik Kabali, Simeranjit Brar, and Saibal Mukhopadhyay. Modeling and designing for accuracy and energy efficiency in wireless electroencephalography systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(1):3:1–3:??, February 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Tsiokanos:2021:DPD

[TMG^{+}21] Ioannis Tsiokanos, Jack Miskelly, Chongyan Gu, Maire O’neill, and Georgios Karakonstantis. DTA-PUF: Dynamic timing-aware physical unclonable function for resourceconstrained devices. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):32:1– 32:24, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3434281.

Tempesti:2007:SRH

[TMM^{+}07] Gianluca Tempesti, Daniel Mange, Pierre-Andre Mudry, Joël Rossier, and Andre Stauffer. Self-replicating hardware for reliability: The Embryonics Project. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(2):9:1–9:??, July 2007. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Torres:2020:NZE

[TNWD20] Frank Sill Torres, Philipp Niemann, Robert Wille, and Rolf Drechsler. Near zero-energy computation using quantumdot cellular automata. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):11:1– 11:16, February 2020. CO-DEN ???? ISSN 1550-4832.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3365394.

Thapliyal:2010:DRS

[TR10] Himanshu Thapliyal and Nagarajan Ranganathan. Design of reversible sequential circuits optimizing quantum cost, delay, and garbage outputs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(4):14:1–14:??, December 2010. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Thapliyal:2013:DER

[TR13] Himanshu Thapliyal and Nagarajan Ranganathan. Design of efficient reversible logic-based binary and BCD adder circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(3):17:1–17:??, September 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Tirelli:2024:SBE

[TSÁ^{+}24] Cristian Tirelli, Juan Sapriza, Rubén Rodríguez Álvarez, Lorenzo Ferretti, Benoît Denkinger, Giovanni Ansaloni, José Miranda Calero, David Atienza, and Laura Pozzi. SAT-based exact modulo scheduling mapping for resource-constrained CGRAs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(3):8:1– 8:??, July 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3663675.

Todri-Sanial:2015:GES

[TSB15] Aida Todri-Sanial and Sanjukta Bhanja. Guest editorial: Special issue on advances in design of ultralow power circuits and systems in emerging technologies. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(2): 11:1–11:??, August 2015. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Todri-Sanial:2017:GES

[TSMCB17] Aida Todri-Sanial, Saraju P.

Mohanty, Mariane Comte, and Marc Belleville. Guest editorial: Special issue on nanoelectronic circuit and system design methods for the mobile computing era. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):12:1–12:??, March 2017. CODEN ???? ISSN 1550-4832.

Taha:2020:AMM

[TT20] Mohammad M. A. Taha and Christof Teuscher. Approximate memristive in-memory Hamming distance circuit. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2):18:1– 18:14, April 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3371391.

Tran:2021:CCC

[TT21] Dat Tran and Christof Teuscher. Computational capacity of complex memcapacitive networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):17:1–17:25, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3445795.

Thuries:2022:ISI

[TTS22] Sébastien Thuries and Aida Todri-Sanial. Introduction to the special issue on monolithic 3D: Technology, design and computing systems applications perspectives. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):19:1– 19:3, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3487869.

Topaloglu:2017:EJS

[TV17] Rasit O. Topaloglu and Naveen Verma. Editorial for JETC special issue on alternative computing systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):38:1– 38:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Thompson:2022:IPS

[TW22] David Thompson and Haibo Wang. Integrated power signature generation circuit for IoT abnormality detection. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):5:1– 5:13, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3460476.

Tang:2009:DET

[TWL09] Weiguo Tang, Lei Wang, and Fabrizio Lombardi. A defect/error-tolerant nanosystem architecture for DSP. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(4):18:1– 18:??, November 2009. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Tang:2019:MNT

[TWLL19] Yibin Tang, Ying Wang, Huawei Li, and Xiaowei Li. MV-Net: Toward realtime deep learning on mobile GPGPU systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):35:1– 35:??, December 2019. CO-DEN ???? ISSN 1550-

                                        4832. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358696.

Tann:2020:REE

[TZR20] Hokchhay Tann, Heng Zhao, and Sherief Reda. A resourceefficient embedded iris recognition system using fully convolutional networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(1):8:1–8:23, February 2020. CODEN ???? ISSN 1550-4832.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3357796.

Tida:2014:NTS

[TZS14] Umamaheswara Rao Tida, Cheng Zhuo, and Yiyu Shi. Novel through-silicon-via inductor-based on-chip DC– DC converter designs in 3D ICs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2): 16:1–16:??, November 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Uddin:2018:DCM

[UMB^{+}18] Mesbah Uddin, MD. Badruddoja Majumder, Karsten Beckmann, Harika Manem, Zahiruddin Alamgir, Nathaniel C. Cady, and Garrett S. Rose. Design considerations for memristive crossbar physical unclonable functions. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1):2:1–2:??, March 2018. CODEN ???? ISSN 1550-4832.

VanRynbach:2018:QCP

[VAK18] Andre Van Rynbach, Muhammad Ahsan, and Jungsang Kim. A quantum computing performance simulator based on circuit failure probability and fault path counting. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1): 13:1–13:??, March 2018. CO-DEN ???? ISSN 1550-4832.

Vatajelu:2016:SMB

[VDB^{+}16] Elena Ioana Vatajelu, Giorgio Di Natale, Mario Barbareschi, Lionel Torres, Marco Indaco, and Paolo Prinetto. STT–MRAM-based PUF architecture exploiting magnetic tunnel junction fabricationinduced variability. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(1):5:1–5:??, December 2016. CODEN ???? ISSN 1550-4832.

Beirendonck:2021:SCR

[VDK^{+}21] Michiel Van Beirendonck, Jan-Pieter D’anvers, Angshuman Karmakar, Josep Balasch, and Ingrid Verbauwhede. A side-channel-resistant implementation of SABER. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):10:1–10:26, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3429983.

Vacca:2011:ASN

[VGZ11] Marco Vacca, Mariagrazia Graziano, and Maurizio Zamboni. Asynchronous solutions for nanomagnetic logic circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(4):15:1–15:??, December 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Vanwinkle:2018:SSH

[VK18] Scott Vanwinkle and Avinash Karanth Kodi. SHARP: Shared heterogeneous architecture with reconfigurable photonic network-on-chip. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):25:1–25:??, July 2018. CODEN ???? ISSN 1550-4832.

VanMeter:2008:ADM

[VMNI08] Rodney Van Meter, W. J.

Munro, Kae Nemoto, and Kohei M. Itoh. Arithmetic on a distributed-memory quantum multicomputer. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(4):2:1–2:??, January 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Venken:2013:SBM

[VMV13] Lyn Venken, Kathleen Marchal, and Jos Vanderleyden. Synthetic biology and microdevices: a powerful combination. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(4):30:1–30:??, November 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

VanMeter:2006:AIQ

[VO06] Rodney Van Meter and Mark Oskin. Architectural implications of quantum computing technologies. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(1):31–63, January 2006. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Vu:2019:CAP

[VOB19] The H. Vu, Yuichi Okuyama, and Abderazek Ben Abdallah. Comprehensive analytic performance assessment and K-means based multicast routing algorithm and architecture for 3D-NoC of spiking neurons. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(4):34:1–34:??, December 2019. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3340963.

Beuningen:2016:PPR

[VRBS16] Anja Von Beuningen, Luca Ramini, Davide Bertozzi, and Ulf Schlichtmann. PRO-TON+: a placement and routing tool for 3D optical networks-on-chip with a single optical layer. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):44:1–44:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Ved:2019:PPA

[VSM19] Sneha N. Ved, Sarabjeet Singh, and Joycee Mekie. PANE: Pluggable asynchronous network-on-chip simulator. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):7:1– 7:??, February 2019. CODEN ???? ISSN 1550-4832.

Venkatesan:2015:EEA

[VSRR15]  Rangharajan  Venkatesan,

Mrigank Sharad, Kaushik Roy, and Anand Raghunathan. Energy-efficient allspin cache hierarchy using shift-based writes and multilevel storage. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(1):4:1–4:??, July 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Volkel:2022:DDI

[VTKT22] Kevin Volkel, Kyle J. Tomek, Albert J. Keung, and James M. Tuck. DINOS: Data INspired Oligo Synthesis for DNA data storage. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):53:1– 53:35, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3510853.

Wenck:2010:SST

[WCSA10] Justin Wenck, Jamie Collier, Jeff Siebert, and Rajeevan Amirtharajah. Scaling selftimed systems powered by mechanical vibration energy harvesting. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(2):5:1–5:??, June 2010. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Wille:2022:ISI

[WD22] Robert Wille and Rolf Drechsler. Introduction to the special issue on design automation for quantum computing.

ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):10:1– 10:2, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3485041.

Wang:2020:CDF

[WDG^{+}20] Lei Wang, Yu Deng, Rui Gong, Wei Shi, Li Luo, and Yongwen Wang. CSMO–DSE: Fast and precise applicationdriven DSE guided by criticality and sensitivity analysis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(2): 16:1–16:22, April 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3371406.

Wang:2009:UQD

[WDH^{+}09] Shuo Wang, Jianwei Dai, El-Sayed Hasaneen, Lei Wang, and Faquir Jain. Utilizing quantum dot transistors with programmable threshold voltages for low-power mobile computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(3):15:1–15:??, August 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Wille:2014:ISI

[WDT14] Robert Wille, Rolf Drechsler, and Mehdi B. Tahoori. Introduction to the Special Issue on Reversible Computation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(2): 8:1–8:??, November 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Wang:2013:HRD

[WDW13] Shuo Wang, Jianwei Dai, and Lei Wang. Hybrid redundancy for defect tolerance in molecular crossbar memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(1):7:1–7:??, February 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Wu:2009:SCD

[WFCX09] Xiaoxia Wu, Paul Falkenstern, Krishnendu Chakrabarty, and Yuan Xie. Scan-chain design and optimization for three-dimensional integrated circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(2):9:1–9:??, July 2009. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Wang:2021:QWY

[WGY21] Serena Wang, Maya Gupta, and Seungil You. Quit when you can: Efficient evaluation of ensembles by optimized ordering. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):55:1–55:20, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3451209.

Wang:2021:HTA

[WHB^{+}21] Xinmu Wang, Tamzidul Hoque, Abhishek Basak, Robert Karam, Wei Hu, Maoyuan Qin, Dejun Mu, and Swarup Bhunia. Hardware Trojan attack in embedded memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(1):6:1–6:28, January 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3422353.

Wu:2021:DRS

[WHL^{+}21] Chia-Cheng Wu, Yi-Hsiang Hu, Chia-Chun Lin, Yung-Chih Chen, Juinn-Dar Huang, and Chun-Yao Wang. Diagnosis for reconfigurable single-electron transistor arrays with a more generalized defect model. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):15:1–15:23, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3444751.

Witharana:2023:AGS

[WJWM23] Hasini Witharana, Aruna Jayasena, Andrew Whigham, and Prabhat Mishra. Automated generation of security assertions for RTL models. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(1):8:1– 8:??, January 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3565801.

Wang:2016:NPM

[WKL16] Qian Wang, Yongtae Kim, and Peng Li. Neuromorphic processors with memristive synapses: Synaptic interface and architectural exploration. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):35:1–35:??, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Wang:2019:TBN

[WL19] Weijia Wang and Bill Lin.

Trained biased number representation for ReRAM-based neural network accelerators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(2): 15:1–15:??, June 2019. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3304107.

Wang:2022:OUN

[WL22a] Weijia Wang and Bill Lin. Optimizing 3D U-Net-based brain tumor segmentation with integer-arithmetic deep learning accelerators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):25:1– 25:16, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3495210.

Wu:2022:ZHT

[WL22b] You Wu and Lin Li. Zallocator: a high throughput writeoptimized persistent allocator for non-volatile memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):80:1– 80:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3549528.

Wang:2021:GBA

[WLJC21] He Wang, Nicoleta Cucu Laurenciu, Yande Jiang, and Sorin Cotofana. Graphenebased artificial synapses with tunable plasticity. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):50:1– 50:21, June 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3447778.

Weerasena:2024:BCC

[WM24] Hansika Weerasena and Prabhat Mishra. Breaking on-chip communication anonymity using flow correlation attacks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 20(4):12:1– 12:??, October 2024. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3677034.

Wang:2010:DCS

[WOW^{+}10] W. S. Wang, T. O’Donnell, N. Wang, M. Hayes, B. O’Flynn, and C. O’Mathuna. Design considerations of sub-mW indoor light energy harvesting for wireless sensor systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(2):6:1– 6:??, June 2010. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Wang:2017:OSS

[WRWW17] Yao Wang, Liang Rong, Haibo Wang, and Guangjun Wen. One-step sneak-path free read scheme for resistive crossbar memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):25:1– 25:??, March 2017. CODEN ???? ISSN 1550-4832.

Wettin:2013:CNE

[WVGP13] Paul Wettin, Anuroop Vidapalapati, Amlan Gangul, and Partha Pratim Pande. Complex network-enabled robust wireless network-on-chip architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(3):24:1–24:??, September 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Winston:2023:VEF

[WWC23] Caleb Winston, Max Willsey, and Luis Ceze. Virtualizing existing fluidic programs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(3):27:1– 27:??, July 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3558550.

Walter:2019:PRT

[WWG^{+}19] Marcel Walter, Robert Wille, Daniel Große, Frank Sill Torres, and Rolf Drechsler. Placement and routing for tilebased field-coupled nanocomputing circuits is NP-complete (research note). ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(3):29:1–29:??, June 2019. CODEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3312661.

Wang:2009:TAR

[WWJ09] Shuo Wang, Lei Wang, and Faquir Jain. Towards achieving reliable and highperformance nanocomputing via dynamic redundancy allocation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(1):2:1–2:??, January 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Wan:2022:ARA

[WWZ^{+}22] Zhe Wan, Tianyi Wang, Yiming Zhou, Subramanian S. Iyer, and Vwani P. Roychowdhury. Accuracy and resiliency of analog computein-memory inference engines. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):33:1– 33:23, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3502721.

Wang:2015:WAS

[WX15] Jue Wang and Yuan Xie.

A write-aware STTRAMbased register file architecture for GPGPU. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(1):6:1–6:??, July 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Wang:2017:ACP

[WXW^{+}17] Xuan Wang, Jiang Xu, Zhe Wang, Haoran Li, Zhehui Wang, Peng Yang, Luan H. K.

Duong, Rafael K. V. Maeda, and Zhifei Wang. Alleviate chip pin constraint for multicore processor by on/off-chip power delivery system codesign. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2): 19:1–19:??, March 2017. CO-DEN ???? ISSN 1550-4832.

Wang:2020:HSS

[WYZ^{+}20] Xueyan Wang, Jienlei Yang, Yinglin Zhao, Xiaotao Jia, Gang Qu, and Weisheng Zhao. Hardware security in spin-based computing-in-memory: Analysis, exploits, and mitigation techniques. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(4):37:1– 37:18, October 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3397513.

Wu:2016:RCA

[WZL16] Chengwen Wu, Guangyan Zhang, and Keqin Li. Rethinking computer architectures and software systems for phase-change memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(4):33:1– 33:40, July 2016. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Wang:2009:ENP

[WZSC09] Z. F. Wang, Huaixiu Zheng, Q. W. Shi, and Jie Chen.

Emerging nanodevice paradigm: Graphene-based electronics for nanoscale computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(1):3:1– 3:??, January 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Xu:2008:IDR

[XC08] Tao Xu and Krishnendu Chakrabarty. Integrated droplet routing and defect tolerance in the synthesis of digital microfluidic biochips. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(3):11:1– 11:??, August 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Xie:2008:ESI

[XCF08] Yuan Xie, Jason Cong, and Paul Franzon. Editorial: Special issue on 3D integrated circuits and microarchitectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(4):15:1– 15:??, October 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Xu:2019:GGP

[XCS^{+}19] Zhen Xu, Xuhao Chen, Jie Shen, Yang Zhang, Cheng Chen, and Canqun Yang. GARDENIA: a graph processing benchmark suite for nextgeneration accelerators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):9:1–9:??, February 2019. CODEN ???? ISSN 1550-4832.

Xie:2014:TCP

[XDX14] Jing Xie, Yang Du, and Yuan Xie. Testable cross-power domain interface (CPDI) circuit design in monolithic 3D technology. ACM Journal on Emerging Technologies in Computing Systems (JETC), 11(1):5:1–5:??, September 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Xu:2007:ADP

[XHSC07] Tao Xu, William L. Hwang, Fei Su, and Krishnendu Chakrabarty. Automated design of pin-constrained digital microfluidic biochips under droplet-interference constraints. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(3):14:1–14:??, November 2007. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Xie:2006:DSE

[XLBB06] Yuan Xie, Gabriel H. Loh, Bryan Black, and Kerry Bernstein. Design space exploration for 3D architectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 2(2):65– 103, April 2006. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Xu:2018:IHS

[XLL^{+}18] Shi Xu, Zhang Luo, Mingche Lai, Zhengbin Pang, and Renfa Li. Integrated highspeed optical SerDes over 100GBd based on optical time division multiplexing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):27:1– 27:??, July 2018. CODEN ???? ISSN 1550-4832.

Xu:2018:EHI

[XLW^{+}18] Xiaowei Xu, Qing Lu, Tianchen Wang, Yu Hu, Chen Zhuo, Jinglan Liu, and Yiyu Shi. Efficient hardware implementation of cellular neural networks with incremental quantization and early exit. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4):48:1– 48:??, December 2018. CO-DEN ???? ISSN 1550-4832.

Xu:2018:SPC

[XNK18] Jiang Xu, Yuichi Nakamura, and Andrew Kahng. Silicon photonics for computing systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):20:1–20:??, July 2018. CODEN ???? ISSN 1550-4832.

Xu:2012:EPV

[XPD12] Hu Xu, Vasilis F. Pavlidis, and Giovanni De Micheli. Effect of process variations in 3D global clock distribution networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(3):20:1–20:??, August 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Xiang:2014:TDT

[XS14] Dong Xiang and Kele Shen. A thermal-driven test application scheme for pre-bond and post-bond scan testing of three-dimensional ICs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(2):18:1– 18:??, February 2014. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Xie:2021:AIB

[XYJ^{+}21] Wen Xie, Zeyang Yao, Erchao Ji, Hailong Qiu, Zewen Chen, Huiming Guo, Jian Zhuang, Qianjun Jia, and Meiping Huang. Artificial intelligencebased computed tomography processing framework for surgical telementoring of congenital heart disease. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):60:1– 60:24, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3457613.

Xu:2018:PVT

[XYM18] Yi Xu, Jun Yang, and Rami Melhem. A process-variation-tolerant method for nanophotonic on-chip network. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):22:1– 22:??, July 2018. CODEN ???? ISSN 1550-4832.

Xu:2021:MCC

[XZL^{+}21] Xiaowe Xu, Jiawei Zhang, Jinglan Liu, Yukun Ding, Tianchen Wang, Hailong Qiu, Haiyun Yuan, Jian Zhuang, Wen Xie, Yuhao Dong, Qianjun Jia, Meiping Huang, and Yiyu Shi. Multi-cycle-consistent adversarial networks for edge denoising of computed tomography images. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):58:1–58:16, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3462328.

Xu:2022:ASP

[XZR^{+}22] Kangwei Xu, Dongrong Zhang, Qiang Ren, Yuanqing Cheng, and Patrick Girard. Allspin PUF: an area-efficient and reliable PUF design with signature improvement for spin-transfer torque magnetic cell-based all-spin circuits. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):71:1– 71:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3517811.

Yu:2018:SOF

[YJ18] Ye Yu and Niraj K. Jha. Statistical optimization of Fin-FET processor architectures under PVT variations using dual device-type assignment. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1):3:1– 3:??, March 2018. CODEN ???? ISSN 1550-4832.

Yuan:2014:FEA

[YL14] Bo Yuan and Bin Li. A fast extraction algorithm for defect-free subcrossbar in nanoelectronic crossbar. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(3):25:1– 25:??, April 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Yogendra:2017:CST

[YLF^{+}17] Karthik Yogendra, Chamika Liyanagedera, Deliang Fan, Yong Shim, and Kaushik Roy. Coupled spin-torque nano-oscillator-based computation: a simulation study. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(4):56:1– 56:??, August 2017. CODEN ???? ISSN 1550-4832.

Yang:2023:HOE

[YLR^{+}23] Mingdai Yang, Qiuwen Lou, Ramin Rajaei, Mohammad Reza Jokar, Junyi Qiu, Yuming Liu, Aditi Udupa, Frederic T. Chong, John M. Dallesasse, Milton Feng, Lynford L. Goddard, X. Sharon Hu, and Yanjing Li. A hybrid opticalelectrical analog deep learning accelerator using incoherent optical signals. ACM Journal on Emerging Technologies in Computing Systems (JETC), 19(2):14:1–14:??, April 2023. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3584183.

Yang:2021:DRA

[YMWH21] Qing Yang, Jiachen Mao, Zuoguan Wang, and “Helen” Li Hai. Dynamic regularization on activation sparsity for neural network efficiency improvement. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(4):51:1– 51:16, October 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3447776.

Yuan:2017:VAR

[YP17] Bo Yuan and Keshab K.

Parhi. VLSI architectures for the Restricted Boltzmann Machine. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):35:1–35:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Yang:2013:MDC

[YW13] J. Joshua Yang and R. Stanley Williams. Memristive devices in computing system: Promises and challenges. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(2):11:1– 11:??, May 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Yang:2018:IAC

[YWF18] Xiaokun Yang, Wujie Wen, and Ming Fan. Improving AES core performance via an advanced ASBUS protocol. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(1): 6:1–6:??, March 2018. CO-DEN ???? ISSN 1550-4832.

Yang:2013:NAC

[YWH^{+}13] Shengqi Yang, Wenping Wang, Mark Hagan, Wei Zhang, Pallav Gupta, and Yu Cao. NBTI-aware circuit node criticality computation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(3): 23:1–23:??, September 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Yu:2017:RMA

[YXD^{+}17] Songping Yu, Nong Xiao, Mingzhu Deng, Fang Liu, and Wei Chen. Redesign the memory allocator for non-volatile main memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):49:1–49:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Ye:2012:TBH

[YXW^{+}12] Yaoyao Ye, Jiang Xu, Xiaowen Wu, Wei Zhang, Weichen Liu, and Mahdi Nikdast. A torus-based hierarchical optical-electronic Networkon-Chip for multiprocessor System-on-Chip. ACM Journal on Emerging Technologies in Computing Systems (JETC), 8(1):5:1–5:??, February 2012. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Yoon:2019:SLC

[YYBK19] Su-Kyung Yoon, Young-Sun Youn, Bernd Burgstaller, and Shin-Dug Kim. Self-learnable cluster-based prefetching method for DRAM-flash hybrid main memory architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):10:1– 10:??, February 2019. CO-DEN ???? ISSN 1550-4832.

Yuh:2007:PDT

[YYC07] Ping-Hung Yuh, Chia-Lin Yang, and Yao-Wen Chang. Placement of defect-tolerant digital microfluidic biochips using the T -tree formulation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(3):13:1– 13:??, November 2007. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Yoon:2017:MUM

[YYPK17] Su-Kyung Yoon, Young-Sun Youn, Kihyun Park, and Shin-Dug Kim. Mobile unified memory-storage structure based on hybrid non-volatile memories. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):40:1–40:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhang:2022:RRN

[ZBF^{+}22] Yongan Zhang, Anton Banta, Yonggan Fu, Mathews M. John, Allison Post, Mehdi Razavi, Joseph Cavallaro, Behnaam Aazhang, and Yingyan Lin. RT-RCG: Neural network and accelerator search towards effective and real-time ECG reconstruction from intracardiac electrograms. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):29:1–29:25, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3465372.

Zhao:2007:PTM

[ZC07] Wei Zhao and Yu Cao. Predictive technology model for nano-CMOS design exploration. ACM Journal on Emerging Technologies in Computing Systems (JETC), 3(1):1:1–1:??, April 2007. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhu:2022:DAT

[ZCB^{+}22] Lingjun Zhu, Arjun Chaudhuri, Sanmitra Banerjee, Gauthaman Murali, Pruek Vanna-Iampikul, Krishnendu Chakrabarty, and Sung Kyu Lim. Design automation and test solutions for monolithic 3D ICs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(1):21:1–21:49, January 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3473462.

Zand:2019:CPI

[ZCDD19] Ramtin Zand, Kerem Y. Camsari, Supriyo Datta, and Ronald F. Demara. Composable probabilistic inference networks using MRAM-based stochastic neurons. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(2):17:1– 17:??, June 2019. CO-DEN ???? ISSN 1550-

4832. URL https://dl.

                                        acm.org/ft_gateway.cfm?
                                        id=3304105.

Zhang:2021:FLI

[ZCSG21] Nathan Zhang, Kevin Canini, Sean Silva, and Maya Gupta. Fast linear interpolation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(2):20:1– 20:15, April 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3423184.

Zhang:2017:SSR

[ZCX^{+}17] Hang Zhang, Xuhao Chen, Nong Xiao, Lei Wang, Fang Liu, Wei Chen, and Zhiguang Chen. Shielding STT–RAM based register files on GPUs against read disturbance. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(2):27:1– 27:??, March 2017. CODEN ???? ISSN 1550-4832.

Zoni:2015:MDP

[ZF15] Davide Zoni and William Fornaciari. Modeling DVFS and power-gating actuators for cycle-accurate NoC-based simulators. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3):27:1–27:??, September 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhang:2013:DTU

[ZFT13] Xuehui Zhang, Andrew Ferraiuolo, and Mohammad Tehranipoor. Detection of Trojans using a combined ring oscillator network and offchip transient power analysis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(3): 25:1–25:??, September 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhuo:2015:CLA

[ZGSA15] Cheng Zhuo, Houle Gan, Wei-Kai Shih, and Alaeddin A. Aydiner. A cross-layer approach for early-stage power grid design and optimization. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3): 25:1–25:??, September 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhang:2020:MPR

[ZH20] Fan Zhang and Miao Hu. Mitigate parasitic resistance in resistive crossbar-based convolutional neural networks. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3): 25:1–25:20, July 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3371277.

Zhang:2011:FBP

[ZJ11] Meng Zhang and Niraj K. Jha. FinFET-based power management for improved DPA resistance with low overhead. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(3):10:1–10:??, August 2011. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zaman:2022:EDS

[ZJK22] Md Adnan Zaman, Rajeev Joshi, and Srinivas Katkoori. Early design space exploration framework for memristive crossbar arrays. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(2):42:1– 42:26, April 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3461644.

Zhang:2009:DSE

[ZJS09a] Wei Zhang, Niraj K. Jha, and Li Shang. Design space exploration and data memory architecture design for a hybrid nano/CMOS dynamically reconfigurable architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(4): 17:1–17:??, November 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhang:2009:HNCb

[ZJS09b] Wei Zhang, Niraj K. Jha, and Li Shang. A hybrid nano/ CMOS dynamically reconfigurable system — Part I: Architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(4):16:1–16:??, November 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhang:2009:HNCa

[ZJS09c] Wei Zhang, Niraj K. Jha, and Li Shang. A hybrid Nano/CMOS dynamically reconfigurable system — Part II: Design optimization flow. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(3):13:1– 13:??, August 2009. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhang:2010:LPN

[ZJS10] Wei Zhang, Niraj K. Jha, and Li Shang. Low-power 3D nano/CMOS hybrid dynamically reconfigurable architecture. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(3): 10:1–10:??, August 2010. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhao:2014:SRS

[ZJT^{+}14] Jing Zhao, Yuliang Jin, Kishor S. Trivedi, Rivalino Matias Jr., and Yanbin Wang. Software rejuvenation scheduling using accelerated life testing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 10(1):9:1–9:??, January 2014. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zyarah:2018:STM

[ZK18] Abdullah M. Zyarah and Dhireesha Kudithipudi. Semitrained memristive crossbar computing engine with in situ learning accelerator. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(4):43:1– 43:??, December 2018. CO-DEN ???? ISSN 1550-

4832. URL https://dl.
acm.org/ft_gateway.cfm?
id=3233987.

Zyarah:2019:NAH

[ZK19] Abdullah M. Zyarah and Dhireesha Kudithipudi. Neuromemrisitive architecture of HTM with on-device learning and neurogenesis. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(3):24:1–24:??, June 2019. CODEN ???? ISSN 1550-4832. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3300971.

Zahedi:2022:MTA

[ZLB^{+}22] Mahdi Zahedi, Muah Abu Lebdeh, Christopher Bengel, Dirk Wouters, Stephan Menzel, Manuel Le Gallo, Abu Sebastian, Stephan Wong, and Said Hamdioui. MNEMOSENE: Tile architecture and simulator for memristor-based computationin-memory. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(3):44:1–44:24, July 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3485824.

Zhou:2021:AMH

[ZLGL21] Jun Zhou, Mengquan Li, Pengxing Guo, and Weichen Liu. Attack mitigation of hardware Trojans for thermal sensing via micro-ring resonator in optical NoCs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 17(3):31:1– 31:23, July 2021. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

https://dl.acm.org/doi/
10.1145/3433676.

Zhang:2020:GEA

[ZLWB20] Wei Zhang, Helen Li, Wujie Wen, and Swarup Bhunia. Guest editorial: ACM JETC special issue on new trends in nanolectronic device, circuit, and architecture design:Part 2. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(4):35:1– 35:3, October 2020. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3412343.

Zaeemi:2022:HLM

[ZM22] Meysam Zaeemi and Siamak Mohammadi. High-level modeling and verification platform for elastic circuits with process variation considerations. ACM Journal on Emerging Technologies in Computing Systems (JETC), 18(4):77:1– 77:??, October 2022. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3534971.

Zhang:2015:DCP

[ZMC15] Tiansheng Zhang, Jie Meng, and Ayse K. Coskun. Dynamic cache pooling in 3D multicore processors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(2):14:1– 14:??, August 2015. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zamani:2013:IFV

[ZMT13] Masoud Zamani, Hanieh Mirzaei, and Mehdi B. Tahoori. ILP formulations for variation/defect-tolerant logic mapping on crossbar nanoarchitectures. ACM Journal on Emerging Technologies in Computing Systems (JETC), 9(3):21:1–21:??, September 2013. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhang:2020:LOO

[ZPL^{+}20] Xinyi Zhang, Clay Patterson, Yongpan Liu, Chengmo Yang, Chun Jason Xue, and Jingtong Hu. Low overhead online data flow tracking for intermittently powered non-volatile FPGAs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3): 26:1–26:20, July 2020. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3371392.

Zhan:2008:AMA

[ZS08] Yong Zhan and Sachin S. Sapatnekar. Automated module assignment in stacked-Vdd designs for high-efficiency power delivery. ACM Journal on Emerging Technologies in Computing Systems (JETC), 4(4):18:1–18:??, October 2008. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhao:2019:LST

[ZSPC19] Zhou Zhao, Ashok Srivastava, Lu Peng, and Qing Chen. Long short-term memory network design for analog computing. ACM Journal on Emerging Technologies in Computing Systems (JETC), 15(1):13:1–13:??, February 2019. CODEN ???? ISSN 1550-4832.

Zhang:2011:NPD

[ZSXY11] Xuefu Zhang, Delong Shang, Fei Xia, and Alex Yakovlev. A novel power delivery method for asynchronous loads in energy harvesting systems. ACM Journal on Emerging Technologies in Computing Systems (JETC), 7(4):16:1– 16:??, December 2011. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhao:2015:STD

[ZWL^{+}15] Chenyuan Zhao, Bryant T. Wysocki, Yifang Liu, Clare D. Thiem, Nathan R. McDonald, and Yang Yi. Spike-time-dependent encoding for neuromorphic processors. ACM Journal on Emerging Technologies in Computing Systems (JETC), 12(3):23:1– 23:??, September 2015. CO-DEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhao:2010:ICP

[ZXC10] Yang Zhao, Tao Xu, and Krishnendu Chakrabarty. Integrated control-path design and error recovery in the synthesis of digital microfluidic lab-on-chip. ACM Journal on Emerging Technologies in Computing Systems (JETC), 6(3):11:1–11:??, August 2010. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Zhang:2018:LBT

[ZY18] Zhe Zhang and Yaoyao Ye.

A learning-based thermalsensitive power optimization approach for optical NoCs. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(2):21:1– 21:??, July 2018. CODEN ???? ISSN 1550-4832.